A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1910–1919

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.03

Title word cross-reference

2^{tes} [7].

-und [536].

10 [305]. 11 [459]. 12 [604]. 13 [603]. 1770 [237]. 1789 [221]. 1790 [8]. 18 [1098]. 1800 [120]. 1832 [940]. 1850 [337]. 1906 [378, 95, 168, 421]. 1907 [276]. 1908 [274]. 1909 [110, 503, 140, 754]. 1910 [110, 292, 615, 631]. 1911 [379]. 1912 [375, 556, 411, 559]. 1913 [511, 562]. 1914 [638, 944, 857]. 1915 [756]. 1916 [1039, 968].

2 [937, 1015, 1087]. 20 [165].

30 [998].

4th [232].

9 [248, 740, 845, 937, 1015, 1087].

Abbott [150]. Abhilfe [423]. Ability [114]. Above [870]. Abridged [771]. Abroad [305, 428, 267]. Abstract [77]. Abteilung [7]. Accelerate [92]. Accident [1039]. Accidents [20]. According [437]. Account [521, 469, 92]. Accountants [1021]. Accounts [358, 1021, 704]. Accuracy [270]. Achille [657, 683]. Acre [56]. Act [379, 754, 378, 5]. Actual [777]. Actuarial [605, 886, 605]. Actuelle [964]. Actuelles [206]. Adam [873, 999, 752, 799, 1018, 1034]. Additions [129, 160, 19, 38, 53, 72, 90, 108, 128, 265, 173, 185, 195, 211, 226, 241, 286, 264, 394, 302, 315, 331, 347, 357, 372, 415, 393, 529, 550, 434, 452, 465, 477, 491, 508, 528, 669, 570, 587, 599, 611, 625, 635, 668, 764, 763, 691, 713, 736, 863, 862, 790, 812, 839, 953, 952, 900, 919, 932, 1028, 1027, 974, 992, 1010, 1110, 1045, 1063, 1083, 1109]. Address [398, 112]. Addresses [927]. Adeane [618]. Adelphi [110, 248, 375, 511, 638, 740, 845, 937, 1015, 1087]. Adhérents [138]. Adler [149]. Administration [45, 256, 1053, 915, 1021, 983, 782, 941]. Adolf [409]. Adolphe [829]. Adoption [757]. Adult [815]. Adultes [500]. Advance [109, 544]. Advancement [873, 398]. Advantages [828, 925]. Affected [955]. Africa [468, 304, 829]. Aftalion [558]. After [793, 1035, 922, 945, 957, 1034, 895, 832, 962, 946]. After-War [946]. Against [542, 324]. Age [777, 198, 817, 135, 166, 5, 139]. Agent [892]. Agents [256]. Ages [612, 162, 816]. Agglomerationen [485]. Agrarreform [631]. Agricole [796, 293]. Agricoles [592, 219]. Agricultural [651, 254, 560, 756, 805, 400, 378, 518, 428]. Agriculture [756, 833, 164, 378, 518, 805, 1001, 1020, 593, 118, 429, 832]. aide [500]. Al [516]. Albert [381, 558]. Albin [875]. Albrecht [322]. Alcoholic [419]. Alcoholism [114]. Alcool [969]. Alcooliques [834]. Aldo [117]. Alexander [909]. Alfred [565, 292, 946, 752, 804, 1089, 1103, 8, 10, 495]. Algébriques [202]. Allégations [969]. Allemagne [647]. Allemand [882]. Allemandes [807]. Allemands [190]. Allen [1071]. Alliés [947]. Almänna [278]. Alter [115]. Ambr [422]. America [884, 559, 783, 605, 231, 454, 806, 858]. American [1100, 621, 926]. Amerika [503]. Among [469]. Amount [133]. Amte [7, 81]. Analysis [927]. Analytical [573, 590, 613, 627, 640, 749, 984]. Anatole [500]. Andelsbevoegelsen [1020]. Anderson [149]. Andrew [1021, 139]. Andrews [1004]. Angel [103, 280]. Anglais [190]. Angleterre [483, 164, 746, 595, 12]. Anleihen [205]. Anna [337]. Annals [165, 998]. Anni [499]. Annie [167]. Annual [129, 265, 394, 529, 669, 764, 863, 953, 1028, 1110, 80, 535, 111, 110, 249, 248, 376, 375, 512, 511, 639, 638, 741, 740, 846, 845, 938, 937, 1016, 1015, 1087, 1088]. Annuities [166]. ans [948, 580]. Answers [64]. Anthropological [577, 914]. Antiquite [484]. Antonelli [659]. Anwendung [215]. Appareils [202]. Apparences [882]. Appendix [123, 960]. Application [847, 212]. Applications [403]. Applied [806]. Appliquée [11]. Appliquées [629]. Appointed [783]. Appreciation [766]. Après [947, 1019, 829]. Arbeit [648]. Arbeiten [277]. Arbeiter [337, 364, 7]. Arbeiter-Statistik [7]. Arbeitern [277]. Arbeitsbeschaffung [291]. Arbeitsblatte [7]. Arbeitslosenstatistik [684]. Arbeitslosenversicherung [291]. Arbeitsolhn [503]. Arbeitsverhältnisse [409]. Arbeitsvermittelung [291]. Arbeitszeit [503]. Arbitrage [153]. Arbitration [801]. Archibald [660]. Area [56]. Argentine [294]. Arguments [880]. Arias [1093]. Arithmétiques [202]. Armand [11]. Armenpfleger [685]. Armenpfleger-Konferenzen [685]. Armenwesen [685, 685]. Arminjon [235]. Arnaune [258, 681]. Arne [847]. Arranged [873]. Arthur [363, 1041, 59, 656, 878, 79]. Articles [18, 37, 52, 71, 89, 107, 127, 159, 172, 184, 194, 210, 225, 240, 263, 285, 301, 314, 330, 346, 356, 371, 392, 414, 433, 451, 464, 476, 490, 507, 527, 549, 569, 586, 598, 610, 624, 634, 667, 690, 712, 735, 762, 789, 811, 838, 861, 899, 918, 931, 951, 973, 991, 1009, 1026, 1044, 1062, 1082, 1108]. Artificial [725]. Arts [110, 375, 511, 638].

Arundel [831]. Arzt [460, 460]. Ascertained [777, 1017]. Ascertainment [869, 936]. Ascribed [288]. Ashby [745]. Ashley [881, 296, 66, 152]. Aspectos [887]. Aspects [256, 758, 934]. Aspetto [747]. Assessment [1002]. Assistance [123, 219]. Assisted [216]. Associated [641]. Association [521, 873, 1039, 398, 605, 348, 133]. Assurance [605, 408]. Attempt [213]. Attributes [348]. Audacious [746]. Auflösung [631]. Aufsätze [323]. Aug [681, 258]. Augier [280]. August [291]. Augusto [98]. Augustus [232]. Ausbildungs [409]. Ausbildungs- [409]. Auseinandersetzung [631]. Ausgang [337]. Auslande [85]. Australasia [801, 454]. Australia [1017, 960, 1051]. Australian [1075]. Austria [888]. Austria-Hungary [888]. Auswanderung [446]. Authentic [831]. Authorities [146]. Autriche [647]. Aux [500]. Available [694]. Avant [221, 1019, 829]. Avebury [496]. Avenir [879]. Average [270, 817, 92]. Averages [177]. Avril [292]. Avvocato [706].

B [289, 782, 119, 778, 780, 806, 895, 133, 203, 592, 151, 215, 580, 659, 708]. Bâties [351]. Baab [291]. Babson [858]. Bachelier [403, 645]. Bad [558]. Bagehot [742]. Baldwin [257]. Balkrishna [1079]. Ballod [536]. Banca [732]. Bancel [879]. Band [685, 83]. Banerjea [295]. Bank [1038, 29]. Banking [310]. Banks [124]. Bannington [782]. Banque [630]. Banques [630, 220]. Baranowsky [595]. Barbara [940, 327]. Barbour [682]. Baron [1041]. Barrington [742]. Barron [746]. Bart [176]. Bart. [871]. Barthélemy [501]. Barton [797, 725]. Bases [796, 994]. Basis [86]. Basler [277]. Bassett [541]. Baudin [825]. Bauer [277]. Bäuerlichen [631]. Bauernfrage [631]. Baumwollindustrie [461]. Bayerns [145]. be [288]. Beable [1097]. Bearbeitet [81, 7]. Beard [784]. Bearing [865, 73]. Beatrice [504, 65]. Beaulieu [290, 470]. Bedeutung [10]. Bedrijfsleven [819]. Before [305, 922, 327]. Begemann [444]. Beginn [361]. Behla [342]. Being [521, 104, 1018, 910, 873, 756]. Beitrag [446]. Beiträge [82]. Bekämpfung [445, 28]. Belgian [792]. Belgique [779, 592, 309, 62]. Belgium [146, 76, 174, 119]. Belief [914]. Bellet [290, 924, 882]. Bellom [13]. Below [133]. Benedict [1023]. Benefit [80, 420]. Bengal [890]. Benn [856]. Benoit [592]. Bento [887]. Beoefening [201, 537, 819, 980]. Bergfried [423]. Berlin [546]. Bernard [603, 947]. Berthold [306]. Bertillon [341]. Bertrand [581, 643]. Berucksichtigung [648, 145, 276, 631]. Beruf [83, 115]. Beruflicher [342]. Berufs [276]. Berufs- [276]. Beschränkung [445]. Besitzungen [168]. Besoins [502]. Besonderer [145, 648, 276, 631]. Besse [164]. Besso [144]. Best [533]. Bétail [138]. Beteiligung [364]. Betrieb [83]. Betriebe [648]. Betriebszählung [276]. Betterment [703]. Between [146, 335, 904, 348, 94]. Bevil [540]. Bevölkerungs [536]. Bevölkerungs- [536]. Bevolkerungsbewegung [485]. Bewegung [337]. Bewertung [445]. Beziehung [342]. Beziehungen [68]. Bi [763, 862, 952, 1027]. Bi-Monthly [763, 862, 952, 1027]. Biddulph [871]. Bigwood [1057]. Bilan [366]. Bilgram [654]. Bill [327]. Billiard [779]. Bills [153, 826]. Binney [386]. Biology [1031]. Birmingham [709]. Birth [848, 515, 701, 553, 641, 768, 935, 961]. Birth- [935]. Birth-Rate [848, 641, 961]. Birth-Rate.Part [701]. Birth-Rates [768]. Births [535, 396]. bis [337, 444, 81, 168]. Black [149]. Bland [753]. Blé [15]. Bloch [351]. Boag [405]. Board [189, 163, 76, 231, 556, 378, 95, 518]. Boards [1039, 754]. Boas [458]. Bodart [888].

Bodily [458]. Body [1023]. Bogart [884]. Bohn [980]. Boissons [834]. Bonds [926]. Bonger [907]. Bonynge [847]. Book [85, 99, 100, 146, 147, 165, 206, 207, 222, 235, 236, 272, 273, 275, 277, 308, 309, 325, 337, 343, 363, 380, 381, 405, 409, 410, 427, 442, 443, 444, 472, 517, 520, 561, 562, 565, 620, 621, 628, 759, 998, 1041, 1092, 1097, 1104, 13, 24, 27, 32, 45, 60, 62, 64, 80, 121, 122, 123, 124, 141, 142, 148, 199, 216, 217, 218, 256, 271, 281, 289, 290, 292, 296, 305, 339, 340, 379, 389, 401, 402, 406, 408, 420, 440, 470, 500, 501, 502, 503, 504, 521, 541, 544, 563, 576, 582, 603, 614, 617, 649, 651, 657, 658, 663, 706, 709, 729, 732, 747, 757, 776, 777, 782, 796, 848, 849, 851, 873, 885, 890, 891, 907, 927, 961, 1038, 1039, 1053]. Book [1056, 1074, 1079, 1093, 1095, 1099, 1100, 1105, 364, 15, 31, 44, 65, 119, 138, 155, 167, 189, 191, 233, 254, 255, 257, 294, 306, 407, 428, 430, 439, 447, 461, 560, 595, 653, 662, 721, 743, 753, 756, 758, 778, 780, 801, 803, 805, 806, 808, 824, 852, 895, 911, 912, 943, 946, 963, 987, 1001, 1004, 1020, 139, 163, 203, 253, 341, 499, 535, 593, 699, 1101, 49, 542, 629, 892, 939, 59, 61, 97, 120, 140, 144, 145, 214, 220, 278, 310, 311, 365, 425, 485, 592, 681, 797, 820, 1057, 1076, 14, 81, 115, 326, 342, 362, 382, 384, 385, 411, 445, 459, 460, 486, 524, 546, 577, 605, 644, 648, 697, 700, 701, 744, 745, 847, 853, 886, 888, 889, 896]. Book [942, 968, 981, 988, 1023, 1037, 1094, 33, 48, 76, 77, 149, 150, 151, 166, 200, 231, 232, 360, 677, 725, 742, 754, 755, 835, 967, 1075, 7, 116, 685, 748, 1077, 1102, 26, 114, 154, 215, 252, 403, 404, 29, 153, 205, 353, 388, 481, 482, 556, 607, 999, 63, 68, 96, 279, 321, 338, 361, 471, 522, 566, 594, 656, 676, 728, 730, 752, 804, 823, 830, 831, 876, 908, 915, 924, 970, 1006, 1021, 1035, 1073, 12, 30, 34, 46, 47, 66, 67, 86, 103, 104, 152, 204, 221, 234, 237, 258, 259, 280, 297, 298, 324, 351, 352, 366, 383, 386, 387, 426, 483, 484, 523, 543, 558, 564, 579, 580, 581, 606, 616, 618, 642, 655, 659, 660, 664, 678, 682]. Book [686, 687, 703, 704, 705, 707, 726, 749, 779, 784, 799, 800, 802, 807, 821, 832, 833, 850, 855, 878, 879, 881, 882, 883, 884, 940, 944, 945, 947, 962, 965, 966, 985, 986, 1000, 1003, 1018, 1019, 1033, 1040, 1051, 1055, 1058, 1059, 1068, 1069, 1070, 1078, 79, 118, 164, 178, 327, 368, 480, 539, 540, 604, 652, 654, 746, 84, 98, 102, 117, 180, 190, 201, 291, 367, 378, 441, 448, 498, 519, 538, 545, 557, 559, 583, 615, 630, 643, 661, 679, 680, 683, 702, 723, 724, 727, 750, 751, 773, 774, 775, 783, 785, 786, 795, 798, 822, 825, 854, 857, 874, 875, 877, 893, 909, 910, 914, 928, 941, 948, 964, 983, 984, 997, 1002, 1005, 1017, 1022, 1034, 1052, 1054]. Book [1071, 1089, 1090, 1091, 1096, 1103, 731, 913, 274, 578, 8, 9, 10, 11, 25, 28, 82, 83, 95, 101, 143, 168, 219, 276, 293, 295, 307, 322, 323, 422, 423, 429, 446, 518, 536, 537, 619, 631, 646, 647, 684, 708, 720, 722, 781, 819, 826, 827, 828, 829, 834, 856, 858, 880, 894, 925, 926, 969, 980, 982, 1036, 1072, 1098, 78, 179, 202, 336, 421, 424, 438, 458, 497, 516, 645, 650, 698, 960, 887]. Booth [1104, 872]. Borel [644, 26]. Borght [83]. Boris [943]. Borntraeger [445]. Borsa [732]. Bosanquet [677]. Bosco [98]. Boston [257]. Bosworth [257]. Bouché [592]. Bourgin [352]. Bowley [720, 743, 59, 744]. Boxes [1105]. Boxmakers [912]. Boyaval [367]. Brassey [703]. Brauindustrie [443]. Bread [985, 1003]. Breckinridge [150]. Brend [942]. Brief [908]. Brisco [729]. Britain [378, 1098, 161, 521, 1004, 983]. Britischen [364]. British [45, 141, 873, 133, 604, 913, 829, 358, 377, 637, 398, 902, 418, 591, 718, 510, 717, 561, 296, 603, 820, 893]. Broader [256]. Broke [894]. Brouihet [387]. Brouilhet [46]. Brown [753, 233, 728, 828, 925]. Bruneau [620]. Bruni [706]. Bruno [289]. Brunt [981]. Budget [825, 1091]. Budgétaire [206].

Budgets [603]. Buelens [647, 647]. Building [95]. Bullion [153]. Bundes [685]. Bureau [1017]. Burn [699, 29]. Burnett [720]. Burnett-Hurst [720]. Burnham [856]. Business [256, 541, 895, 558, 654, 970]. Businesses [601]. Butler [151, 894]. Byers [913].

C [272, 176, 123, 890, 852, 987, 1101, 49, 425, 820, 1076, 14, 326, 988, 149, 685, 772, 361, 67, 704, 986, 1068, 1069, 368, 746, 201, 559, 680, 724, 775, 795, 731, 537, 819, 894, 980]. Cadbury [709]. Cahill [518]. Cahn [798]. Calcul [403, 202, 403]. Calculated [1014, 958]. Calculating [817]. Calculation [162, 702]. Called [823, 1039]. Calmes [381]. Camille [9]. Can [1059]. Canada [294, 140, 835, 361, 1018, 255, 480]. Canal [454, 751]. Cancer [382, 853, 988]. Cane [731]. Cannan [447, 1054]. Cap [306]. Capital [767, 653, 1035, 739, 161, 1071, 1022, 680, 1089, 798, 1022]. Capitalism [911]. Carl [1094, 361, 536]. Carlile [353]. Carqueja [887]. Carr [339]. Carter [1098]. Cartwright [27]. Carvallo [403]. Case [122, 542, 324]. Casimir [362]. Castberg [606]. Cathcart [1023]. Catherine [14]. Causal [33]. Cause [654, 698]. Causes [142, 218, 848, 823, 558, 288, 5]. Cecil [946]. Censal [212]. Census [242, 320, 468, 251, 229, 377, 589, 439, 572, 816, 478, 674, 378, 1017, 574, 492, 869, 936, 437, 555, 438, 960, 614]. Census-Taker [439]. Cent [580]. Central [146, 756, 867, 881]. Centralised [749]. Centuries [985]. Century [165, 998, 436, 92, 6, 22, 42, 58, 146, 430, 780, 360, 1033]. Céréales [15]. Certain [573, 590, 613, 627, 640, 3, 33, 76, 231, 469, 556, 467]. Ch [619]. Chômage [218]. Chairman [133]. Chamber [756]. Chamberlain [324]. Chambre [880]. Chance [645]. Chancellor [910]. Change [681]. Changes [612, 1066, 400, 92, 458]. Chapman [703, 1035, 298]. Chapters [29, 909]. Character [823, 960]. Characters [553]. Charity [677]. Charles [872, 31, 577, 46, 280, 387, 618, 660, 784, 884, 829, 1104]. Charlotte [1105]. Charter [757]. Charters [757]. Chefs [62]. Chessa [545]. Chester [147]. Chevilliard [388]. Cheysson [424, 40]. Chief [1085]. Chih [786]. Child [163, 524, 203]. Child-Labour [524]. Children [1004, 479, 816, 78]. Chile [1018]. China [748, 786]. Ching [1102]. Chiozza [678]. Choisies [424]. Chomage [290]. Christ [783]. Christabel [881]. Christian [564]. Christians [320]. Christie [724]. Christopher [191, 803]. Chun [1102]. Churches [783]. Cigars [1105]. Cinquanta [499]. Cinquante [580]. Circle [1040]. Città [180]. City [757, 136, 257, 886, 968, 835]. Clair [558]. Clarence [662]. Class [493, 76, 231, 213, 556, 720]. Classe [502]. Classes [303, 1047, 844]. Clay [821]. Clement [908]. Clerks [770]. Cleveland [785]. Cleveland-Stevens [785]. Club [247]. Co [521, 805, 987, 749, 518, 551, 607, 1099, 1020, 593, 607, 429]. Co-operation [805, 518, 1099, 593, 607, 429, 1020]. Co-Operative [551, 521, 987]. Co-ordinated [749]. Co-Partnership [607]. Coût [61, 321]. Coats [140]. Coghlan [1051]. Cohen [830]. Cohn [311]. Coins [153]. Colajanni [582, 117]. Cole [1055, 800, 1055]. Collars [1105]. Colleges [513]. Colletti [499]. Colonial [104, 878, 161]. Coloniales [924]. Colonies [122]. Colson [1068, 1069]. Columbia [524, 1075]. Combination [981]. Commentaire [292]. Comments [541]. Commerce [363, 427, 221, 258, 580, 655, 882, 1018, 913, 829, 880, 925, 1041, 828, 808, 962]. Commercial [561, 1097, 296]. Commerciale [12, 779, 947, 1093]. Commission [32, 123, 199, 305, 848, 1095, 166, 967, 252, 480, 783, 1002]. Commissioner [80]. Commissions [1039]. Committee [628, 397, 514, 305, 617, 33, 835, 616, 133, 892]. Commodities [906, 979, 305, 602, 695, 794, 43, 188, 319, 455]. Common [86, 368]. Commons [178]. Commonwealth [1017, 960, 579, 1055, 913]. Communicated [411]. Communities [253]. Comolet [292]. Comolet-Tirman [292]. Comparative [139]. Comparées [875]. Comparison [379]. Comparisons [4]. Compete [1059]. Competition [454, 876, 751]. Competitive [361]. Complement [232]. Comyns [339]. con [732]. Concentration [12]. Concerning [334]. Conciliation [801]. Concise [310]. Condemnation [835]. Condition [33, 296]. Conditions [406, 1105, 833, 944, 720, 907]. Confectionery [1105]. Conference [134, 1039]. Conflit [46]. Conflitto [1096]. Confronting [1041]. Congress [411]. Connected [288]. Connecticut [1105]. Connection [793, 378, 1054]. Connections [626]. Conquest [857]. Conrad [646]. Conséquences [142, 218]. Considerable [20]. Considered [56]. Considérée [630]. Consistenza [706]. Consisting [133]. Consols [349, 318]. Consommation [581, 834]. Constance [660]. Constitution [29]. Constructing [198, 771]. Constructive [946, 385]. Consumption [419, 288]. Contemporaine [46]. Contemporaines [502]. Contento [117]. Continental [122]. Contract [721]. Contrat [647]. Contre [367]. Contribution [1098]. Control [728, 1002, 1098]. Conventions [343]. Convict [577]. Convictions [288]. Copeland [461, 970]. Corn [780]. Cornelissen [564]. Corporate [656]. Corporation [831]. Corporations [219]. Corps [9]. Corrado [24, 441]. Correlation [75, 212, 904, 775]. Correlations [575, 94].

Correspondence

[283, 474, 488, 157, 182, 261, 170]. Corso [851]. Cost [305, 767, 1066, 815, 715, 822, 844]. Costs [136, 925, 986]. Cotoniera [663]. Cotter [831]. Cotton [563, 461, 303, 144, 92, 6, 22, 42, 58, 1057]. Coulter [806]. Council [848, 783, 894, 110, 248, 246, 375, 511, 638, 740, 845, 937, 1015, 1087]. Countries [41, 161, 805, 498]. Courcy [908]. Cours [11, 259, 1068, 1069]. Course [1086, 267, 531]. Credit [293, 518, 681, 580, 799]. Credits [806]. Criminality [907]. Crise [164]. Crises [595, 558]. Criterion [533]. Critical [541, 1038, 385]. Critique [306]. Cromer [946]. Crookes [965]. Crop [41, 1067, 227, 696]. Crop-Cycles [1067]. Crowding [335]. Crown [244]. Cultivées [138]. Culture [775]. Cumulatively [5]. Cunningham [324, 911]. Currency [797, 519]. Current [1092, 17, 36, 51, 70, 88, 106, 126, 158, 171, 183, 193, 209, 224, 239, 262, 284, 300, 313, 329, 345, 355, 370, 391, 413, 432, 450, 463, 475, 489, 506, 526, 548, 568, 585, 597, 609, 623, 633, 666, 689, 711, 734, 761, 788, 810, 837, 860, 898, 917, 930, 950, 972, 990, 1008, 1025, 1043, 1061, 1081, 1107]. Curtis [913]. Curtler [118]. Curves [847]. Cycles [1067, 558, 698]. Cyclical [823]. Cyril [1103]. Czuber [215].

D

[363, 339, 1100, 1057, 14, 326, 149, 232, 676, 915, 800, 832, 855, 966, 1055, 604, 984, 1036]. D.C [1039]. D.C.L [175, 872]. D.Sc [175]. D’Abernon [1098]. d’Agriculture [138]. Daily [1003]. dal [24]. Dampier [946]. Dana [909]. Dangers [523]. Daniel [290, 15, 924, 882]. Danish [1020]. Danmark [1020, 311]. dans [502, 138, 46, 834]. D’Assurance [13]. Data [852, 843, 868, 905, 923, 533, 467, 577]. Daudé [879]. Daudé-Bancel [879]. d’Avenel [120]. Davenport [385, 1052]. David [981, 682, 1018, 78]. Davies [876]. Dawbarn [67]. Dawson [217, 946, 122, 379, 730]. Day [256, 798, 1053]. Death [75, 904, 769, 935, 5]. Death-Rates [75, 904, 769, 935]. Deaths [535, 5]. Debito [706]. Debt [1071]. Debts [29]. Decay [253]. December [1039, 110, 248, 375, 417, 511, 638, 740, 845, 937, 1015, 1087]. Décennale [351]. Decline [1079]. Declining [848]. d’Économie [387, 581, 659, 1068, 1069, 501]. Deel [819, 980]. Defective [78]. Déficit [825]. Dehn [563]. Deite [484]. del [1096]. Dell’ [499]. Della [24, 121, 851, 98, 84]. Delusions [615]. Demain [779]. Demand [386]. Demi [948]. d’Émission [220]. Democracy [784]. Demografia [117]. Demografica [117]. Demografici [1096]. Démographie [219]. d’Emploi [647, 647]. Denmark [420]. Dennis [823]. Denrées [120]. Density [612]. Dépôts [630]. Department [1105, 361, 910]. Departmental [628, 480]. Departments [541, 541]. Dépopulation [341]. Depressions [654]. Depuis [120, 484, 580, 31]. Deren [277]. Dernier [297]. Descendants [458]. Deschanel [883]. Descriptive [912]. Descriptives [190]. Designed [852]. Dess [278]. Detention [122]. Deterioration [888]. Determination [553, 986]. Deutsch [153]. Deutsche [99, 460, 578, 422, 446, 81]. Deutschen [100, 277, 409, 444, 520, 485, 7, 205, 82, 83, 101, 423, 446, 684]. Deutschland [275, 445, 236, 101]. Deutschlands [28, 410, 276]. Deux [948]. Deuxième [351, 1069]. Devant [874]. Development [363, 1065, 749, 833, 785, 854, 984, 1002, 1075, 798]. Developments [1013]. Dewing [656]. d’Expansion [807]. d’Exportation [829]. Diagramme [422]. Dibblee [386].

Dictionary [232, 310]. Diets [1032]. Different [41, 844]. Différents [834]. Difficulties [4]. Digest [305]. Dilke [176]. Dimensions [3]. d’Industrie [62]. d’Industries [190]. Dinner [247]. Directors [605]. Direkte [207]. Disappearance [79]. Discount [349, 318]. Discussion [256, 757, 77, 288]. Disease [139]. Distress [123, 32, 199]. Distribution [721, 133, 384, 385, 553, 471, 615, 93, 5]. Distributions [5]. District [890]. Districts [137, 1]. Dittman [485]. Diverses [321]. Divorce [9]. Doctrines [31, 46, 581]. Documents [880]. Dollar [822]. Domestic [33, 104, 925, 894]. Domestica [657]. Dominions [296]. Dottrine [747]. Douane [258]. Douanière [280]. Douglas [389, 542]. Downie [234]. Dr [133, 101]. Dr. [350]. Drage [649, 204]. Drawn [213]. Dresden [83]. Drew [1097]. Drink [1098]. Drinks [288]. Drysdale [705]. Dublin [217]. Dudley [752]. Dufour [522]. Duplice [747]. Durata [121]. Duration [673]. During [146, 45, 716, 1098, 22, 42, 58, 873, 288, 6]. Durning [14].

Earl [946, 664, 703]. Early [1018, 833, 1018]. Earners [1105, 303]. Earning [167]. Earnings [1105, 189, 335, 95, 95]. East [829]. Eclipse [895]. Ecole [15]. Economia [1093]. Economic [147, 562, 18, 37, 52, 71, 89, 107, 127, 159, 172, 184, 194, 210, 225, 240, 263, 285, 301, 314, 330, 346, 356, 371, 392, 414, 433, 451, 464, 476, 490, 507, 527, 549, 569, 586, 598, 610, 624, 634, 667, 690, 712, 735, 762, 789, 811, 838, 861, 899, 918, 931, 951, 973, 991, 1009, 1026, 1044, 1062, 1082, 1108, 124, 271, 873, 890, 907, 927, 447, 753, 287, 637, 398, 1065, 902, 469, 174, 86, 426, 707, 833, 884, 1018, 1033, 1070, 683, 798, 1103, 720, 722, 828, 1013, 879, 165]. Economic [998, 651, 1004, 47, 784, 944, 945, 1034, 698]. Economicamente [121]. Economico [582, 747]. Economicos [887]. Economics [621, 256, 852, 1030, 49, 471, 752, 722, 254, 407, 542, 365, 353, 606, 448, 295, 336, 729, 821]. Economie [750, 629, 46, 619]. Economique [483, 235, 120, 543, 879, 883, 1019, 102, 948, 308, 807]. Économiques [31, 964, 219, 1068]. Economisch [819, 980]. Economist [850]. Economy [86, 941, 66, 298, 660]. Eden [683]. Edgar [461]. Edith [150]. Edition [232]. Edmond [308, 219]. Edmund [338]. Edouard [523]. Education [411]. Edward [1023, 915, 1021, 234, 833, 962, 1000, 654]. Edwards [579]. Edwin [273, 1038, 428, 447, 542, 279, 566, 618, 1054]. Effect [743, 1012]. Effects [123, 1004, 835, 1018, 738, 848, 1074]. Effets [655]. Efficiency [1023, 1098, 729]. Efter [97]. Egbert [444]. Égypte [235]. Égyptien [235]. Eheliche [145]. Eighteenth [780, 1033, 436]. Eighth [376, 375]. Eightieth [639, 638]. Eighty [741, 740, 846, 845, 938, 937, 1016, 1015, 1087, 1088]. Eighty-Fifth [1087, 1088]. Eighty-First [741, 740]. Eighty-Fourth [1016, 1015]. Eighty-Second [846, 845]. Eighty-Third [938, 937]. Einar [311]. Einaudi [851]. Einfluss [115]. Eingeleitet [277]. Eingeleiteten [631]. Einiger [648]. Einkommen [307]. Einkommen-Entwicklung [307]. Elbourne [1021]. Elderton [114, 25, 701, 25]. Election [73]. Elementaire [259]. Elementary [59, 407, 722]. Elements [402, 827, 26, 581]. Eliz [151]. Ellis [687, 749]. Ely [722, 721]. Emanuel [100, 215]. Emanuele [84]. Emigration [559, 694]. Emigrazione [499]. Emil [443]. Emile [250, 26, 644, 40, 424]. Emmet [943]. Empire [377, 589, 895, 55, 104, 962, 510, 717, 913, 961, 255, 30, 664]. Empires [637]. Employed [112, 92]. Employees [886]. Employés [647]. Employing [303]. Employment [123, 374, 359, 228]. Encyclopœdia [310]. Ended [110, 248, 375, 511, 638, 1087, 740, 845, 937, 1015]. Ending [110, 248, 375, 511, 638, 1087, 740, 845, 937, 1015]. Enever [324]. Engineers [1021]. England [146, 163, 535, 701, 29, 327, 137, 841, 612, 287, 39, 436, 994, 674, 694, 56, 1, 23, 768, 769, 770, 935, 551, 911, 1102, 854, 363, 253]. English [651, 254, 780, 577, 701, 118, 652, 1032, 109, 437, 504, 782, 65, 753, 802, 539, 538, 785]. Enigma [340]. Enquêtes [44]. Enquiries [793, 378]. Enquiry [617, 892, 823, 616, 518, 76, 556, 894]. Enrico [706]. Enter [479]. Enterprises [915]. Enthaltend [536]. Entwickelung [99]. Entwicklung [236, 361, 101, 276, 307]. Environment [78, 179]. Épargne [219]. Epidemic [968]. Epidemics [889]. Époques [321]. Ereditaria [545]. Erfolge [82]. Erhebung [7]. Ernährung [277]. Ernest [884, 856]. Errata [137]. Error [816, 177, 494]. Errors [673]. Erven [980]. Escher [827]. Espagne [103]. Essai [306, 497, 500]. Essais [9]. Essay [538, 775, 336]. Essays [104, 678, 48, 566]. Essentials [757]. Esslen [423]. Establishment [1105, 754]. Estates [513, 93]. Estimates [230]. Estimation [212]. Eta [121]. État [442]. Ethel [701, 114, 25]. Etienne [483]. Étranger [13]. Etrangères [924]. Etrangers [352]. Etude [501, 102, 62]. Etudes [482, 219]. Eugen [879]. Eugenics [793, 1030, 411, 326, 48]. Europa [503]. Europaische [759]. Europe [206, 806, 217, 881]. European [1018]. Europeo [1096]. Evans [752]. Eversley [178]. Every [626]. Evidence [305]. Evolution [780, 749, 164]. Examination [3, 154, 481, 615, 774, 877, 1071]. Excess [835]. Exchange [29, 388, 826, 827]. Exchanges [675]. Exchequer [430, 269, 910]. Exclusion [352]. Exemption [133, 835]. Exeter [946].

Exhibition [702]. Existe [142]. Existe-t-il [142]. Existence [321]. Existing [886]. Expenditure [33]. Expenditures [257]. Experience [420, 886, 830]. Experiment [374]. Experimental [494]. Experiments [966]. Explained [699]. Export [680]. Exportation [880]. Exports [781]. Expressly [852]. Extent [469]. Exterieur [258]. External [743]. Extremity [186].

F [275, 277, 27, 401, 1074, 257, 758, 133, 365, 524, 361, 471, 1019, 983, 1022, 1091, 826, 980, 650]. F.R.S [74, 872]. F.R.S. [175, 496]. Faber [1020]. Fabian [910]. Fabric [726]. Fabrik [381]. Fabrik- [381]. Factories [20]. Factors [334, 641, 960]. Factory [1037, 915, 1021]. Facts [64, 406, 47, 686, 892]. Faculty [524, 1075]. Faggot [560]. Failure [469]. Fallacies [686, 1103, 47]. Fallacy [574]. Falling [1054]. Familien [7]. Families [33, 335]. Familles [421]. Family [14, 553, 626]. Farm [855, 704]. Farmers [1099]. Farming [986, 539, 604, 705]. Fatality [186]. Father [479]. Fatigue [1037]. Fattori [1096]. Federal [783]. Federico [657, 545]. Fehlerausgleichung [215]. Félice [142]. Felix [546]. Ferdinand [28]. Fertility [572, 673, 574]. Field [1037, 802]. Fifth [1087, 1088]. Fifty [45, 756]. Figures [574]. Final [628, 438]. Finance [146, 541, 873, 1102, 104, 749, 1018, 814, 723, 1092, 824, 999, 519, 727, 982, 748, 1000]. Finances [259, 746, 875]. Financial [561, 110, 248, 375, 511, 638, 740, 845, 937, 1015, 1087, 333, 541, 287, 55, 726, 679, 983, 909]. Financiere [259, 483, 235, 883]. Financières [219]. Finanz [520, 536]. Finanz- [520]. Finanza [851]. Finanze [747]. Finanzgeschichte [100]. Finanziarie [747]. Finanzreformversuche [444]. Fingland [408]. First [741, 740, 411, 114, 583, 438]. Fiscal [47]. Fishberg [179]. Fisher [233, 847, 407, 822].

Fisheres [518]. Fisheries [378, 274]. Fishery [112]. Fishing [921]. Fit [533]. Fitting [533]. Five [1105, 830]. Fleischpreise [361]. Fleischproduktion [99]. Fleischteuerung [423]. Fleischversorgung [423]. Florence [1037]. Flour [985]. Fluctuation [823]. Fluctuations [558, 228, 552, 1012, 960]. Fondamentali [747]. Fondements [362]. Food [803, 815, 399, 957, 844]. Footpaths [178]. Force [749]. Forces [92]. Forcher [557]. Fordeling [97]. Foreign [675, 418, 591, 718, 161, 826, 827, 925, 926, 826]. Forest [200]. Forestry [893]. Forests [178]. form [458]. Formation [482]. Forms [757, 1021]. Formule [706]. Forsörgelsesbyrde [97]. Forsyth [772]. Fortune [338]. Fortunes [116]. Fothergill [340, 521]. Foundation [895]. Foundations [1070, 1103]. Fourteenth [532]. Fourth [1016, 1015, 535]. Foville [495, 11]. Foxwell [1092]. Fractures [186]. Frage [291]. Francais [804, 190]. Francaise [880, 259]. France [146, 142, 44, 321, 924, 221, 237, 655, 1019, 630, 874, 948, 922, 93, 427, 620, 341, 888, 116, 338, 152, 280, 366, 878, 879, 964, 293]. Francesco [499]. Franches [880]. Francis [175, 697, 1023, 48, 643]. Franco [707, 842]. Franco-German [707]. Franco-Prussian [842]. Frank [64, 1036]. Frankel [217]. Franklin [827]. Franz [440, 458, 57]. Frau [648]. Frederick [256, 853]. Free [104, 324]. Freien [337]. Freiherr [85]. Freiwillige [685]. French [671]. Frequency [139, 847, 5]. Friedrich [889, 881, 276]. Friend [439]. Friendly [521, 467, 716]. Fruchtbarkeit [145]. Funds [80, 886]. Furniss [928]. Further [673]. Future [1001, 678, 717, 913, 858, 996].

G [222, 472, 121, 216, 281, 401, 420, 576, 782, 233, 593, 120, 214, 486, 1023, 149, 215, 388, 297, 366, 558, 678, 800, 1055, 1059, 448, 724, 775, 795, 960]. G.C.B. [766]. Gabrielsson [834]. Gait [614]. Galton [175, 48, 697]. Garnett [339]. Garton [895]. Garvin [1070]. Gas [228]. Gaston [888, 259, 655, 746]. Gebied [819]. Geburtenrückgang [445]. Geburtenzahl [445]. Gegenwart [444]. Gegenwärtiger [99]. Gehe [83]. Gehe-Stiftung [83]. Gelphart [1074]. Gemeindesteurn [207]. General [111, 110, 249, 248, 376, 375, 512, 511, 639, 638, 741, 740, 846, 845, 938, 937, 1016, 1015, 1087, 852, 535, 577, 886, 1088, 821, 73, 120, 13]. Generale [117, 321, 1068, 11]. Geo [203]. Geoffrey [649, 204]. Geometry [573, 590, 613, 627, 640]. George [405, 60, 709, 1057, 386, 884, 1000, 1003, 1033, 1103, 722, 339]. Georges [62, 220, 237, 352, 102, 964]. Gerard [21]. Gérault [878]. Gerhard [322]. Gerlich [99]. Gerloff [520]. German [563, 707, 637, 1032, 55]. Germany [379, 420, 152, 333, 39, 112, 730, 518, 562]. Gesamtwirtschaft [101]. Geschlecht [115]. Geschlechter [24]. Geschlechtsverhältnis [404]. Gesellschaftliche [83]. Gesellschaftslehre [60]. Gesetz [631]. Gesetzliche [685]. Gestion [442]. Gesundheitliche [648]. Gewerbestatistik [646]. Gewerkschaften [82]. Ghose [1101]. Giacomo [732]. Gibbon [216, 420]. Gide [31, 660]. Giffen [74, 576]. Gilbert [662]. Gilds [521]. Gini [24, 441]. Gino [1093]. Ginsberg [775]. Giorgio [180]. Giuseppe [352]. Given [123]. Gleanings [478]. Gliederung [83]. Globe [24]. God [914]. Gold [682, 481, 1072]. Goldstein [28]. Gomme [777]. Gonner [368]. Good [558]. Goodness [533]. Gordon [987, 1057, 1036, 1099]. Gorham [884]. Goring [577]. Goulichambaroff [24]. Governing [717]. Government [504, 541, 163, 730, 327, 877, 222]. Grandes [630]. Grandi [180]. Grant [139]. Gras [780]. Gray [895, 802]. Great [521, 1004, 378, 983, 1098, 161]. Greater [705]. Grèce [883]. Green [988]. Gregory [696]. Grey [664]. Grice [146]. Grignon [15]. Grisewood [687]. Groat [884]. Grossbanken [101]. Grossstadt [143]. Grosstadte [485, 685]. Grosstadtischen [485]. Growing [147, 294]. Growth [253, 1017]. Grundlagen [517, 215]. Grundriss [536]. Grundung [485]. Grundzüge [661]. Grunzel [879]. Gschwendtner [361]. Guerre [804, 746, 875, 964, 829, 947, 1019, 874, 948]. Guilford [47]. Gunnison [828, 925]. Günther [409]. Gürtler [8]. Guyot [442, 594, 969].

H [277, 1092, 27, 339, 541, 777, 907, 233, 560, 753, 756, 895, 912, 946, 1020, 133, 140, 385, 754, 403, 772, 471, 86, 426, 484, 606, 660, 800, 855, 883, 944, 1055, 79, 118, 615, 783, 914, 928, 1022, 1052, 731, 720, 829, 698, 960]. Habitants [219]. Haig [835]. Halbwachs [502, 497]. Haldane [946]. Hall [966, 832, 604]. Hammond [940, 327, 940, 327]. Handbook [702]. Handel [461]. Handels [536]. Handels-Statistik [536]. Handelspolitik [759]. Handicap [829]. Handley [707]. Haney [426]. Hans [207]. Harald [1020, 939, 888, 889]. Harbutt [122, 379, 730]. Haristoy [96]. Harley [1002]. Harman [687]. Harold [593]. Harris [222]. Harry [828, 925]. Hartley [824, 963, 726]. Hasard [644, 645]. Hauser [807]. Haushaltungsstatiski. [322]. Hawtrey [558]. Health [782, 456, 942]. Hearings [305]. Heilbehandlung [459]. Heinrich [99, 759, 143]. Held [111, 110, 249, 248, 375, 512, 639, 638, 1016, 1087, 411, 1088, 376, 511, 741, 740, 846, 845, 938, 937, 1015]. Helen [677]. Helfferich [562]. Henri [807, 643]. Henry [1097, 124, 148, 576, 1001, 153, 830, 821, 679, 941, 293, 429, 1098, 336, 218]. Hepburn [797, 725]. Her [912, 287, 1013]. Herausgegeben [277, 685]. Herbert [541, 835]. Herbst [684]. Herman [798]. Hermann [254, 68, 323]. Heron [78]. Hertel [1020]. Herzog [1040]. Hesse [646]. Heutigen [275]. Heyking [1041]. Heyman [205]. Hicks [361]. Hidden [896]. Hiérarchie [502]. Higgs [576, 679, 941]. High [162, 822]. Highlands [1053]. Highway [504]. Hilton [727]. Hirst [401]. Histoire [427, 31, 120, 221]. Historical [363, 1038, 749, 984]. Historie [483, 484]. History [147, 379, 408, 756, 797, 360, 677, 626, 471, 752, 831, 908, 152, 884, 1033, 118, 559, 753, 1006, 722, 650, 426]. Hoag [662]. Hoare [777]. Hobbs [1004]. Hobhouse [775, 216]. Hobson [481, 642, 680, 1090]. Hodges [944]. Hoffman [256, 853]. Holdings [254]. Holme [823]. Home [267, 78, 428]. Homemorton [1021]. Homeworker [912]. Homme [497]. Homograde [847]. Hon [176, 496, 872, 766]. Hooker [829]. Horace [155, 707, 997]. Horowicz [404]. Horsburgh [702]. Horton [907]. Hospital [186, 213]. Hostudbytte [311]. Hostudsigter [311]. Hough [708]. Hours [189, 95]. Household [1006]. Households [493, 720]. Houses [122]. Housing [76, 231, 967, 976, 1005]. Howard [802]. Hsu [786]. Huart [875]. Huber [364]. Hugh [745, 579]. Hugo [205, 654, 557]. Huishouding [819, 980]. Human [767, 1023, 86, 642]. Humber [701]. Hungary [888]. Hurst [720]. Hutchins [778]. Hutchinson [751]. Hutton [1005]. Hythe [664].

Ida [325]. Ideas [910]. Ihr [99]. Ihre [99, 409, 101, 215, 361, 28]. Ihrem [275]. ihren [68]. II [305, 617, 886, 685, 98, 117, 8, 9, 83]. III [305, 703, 95, 1000]. Illegitimate [641]. Illusions [523]. Illustration [554].

Illustrierte [422]. im [85, 381, 444, 81, 7, 10, 83, 101, 168]. Immigrants [458]. Immortality [914]. Impôt [63]. Impôts [219]. Imperial [204]. Importée [804]. Imports [781]. Impot [484]. Improve [921]. Improvements [835]. Inavalidenversicherung [459]. Incidence [139]. Inclosure [368]. Including [86]. Income [27, 133, 279, 773, 1085, 384, 683, 498, 774]. Incomes [257, 820, 471]. Increase [515, 615, 20, 844]. Increased [56]. Indemnity [707]. Index [575]. India [544, 614, 320, 229, 141, 271, 891, 1079, 45]. Indian [45, 1101, 104, 295, 1095, 519]. Indicating [253]. Indices [94]. Indien [24]. Indigène [804]. Indigène-Viande [804]. Individual [626, 161]. Indtoegt [97]. Industria [663]. Industrial [363, 187, 995, 296, 406, 1039, 1079, 1095, 1105, 895, 1037, 1049, 76, 231, 967, 556, 823, 915, 1006, 928, 1103, 894, 1097]. Industrie [68]. Industrielle [309, 755, 12, 779]. Industriellen [648]. Industrielles [502, 595, 501]. Industriels [62, 594]. Industries [1105, 601, 693, 1036, 921]. Industry [461, 303, 144, 1076, 149, 754, 830, 799, 1018, 1051, 228, 731, 92, 6, 22, 42, 58, 363, 563, 778, 808, 150, 999]. Infant [163, 745]. Infantile [397, 903, 396, 968, 534]. Infants [198]. Inflation [934]. Influence [114, 386, 682, 78, 650, 960]. Information [2, 91, 130, 243, 266, 373, 395, 509, 530, 636, 670, 737, 765, 840, 864, 933, 954, 1011, 1029, 1084]. Inheritance [14]. Injury [515]. Inleiding [201, 537, 819, 980]. Innes [363]. Inquiries [873]. Inquiry [469, 558, 95, 189, 231]. Institut [138]. Institute [805, 268, 532]. Institutions [716, 775]. Instruments [702]. Insurance [187, 80, 148, 216, 217, 256, 379, 927, 1039, 1074, 551, 216, 339, 658, 927, 256, 776, 166]. Intégrateurs [202]. Intelligence [78]. Inter [281]. Inter-Racial [281]. Interest [418, 457, 591, 718, 317, 662]. Interested [256]. International [1039, 138, 805, 411, 112, 751, 4, 134, 268, 532, 824, 828]. Internationale [796, 543]. Internationalen [342]. Internazionali [657]. Interni [657]. Interpretation [574, 94]. Intoxicating [288]. Intoxication [288]. Introduction [408, 214, 326, 360, 821, 884, 1055, 997, 1094]. Investigation [253, 886, 1037, 112, 783, 605, 305]. Investment [926]. Investments [29, 418, 457, 591, 718, 161]. Investors [621]. Ireland [987, 1033, 287, 492, 555]. Irene [1004]. Ireson [64]. Iron [1040]. Irregularity [335]. Irving [233, 407, 822]. Isbell [773]. Islands [1053]. Issued [252]. Italiana [499]. Italiani [180]. Italiano [706]. Italy [995, 144]. IV [305].

J [146, 123, 292, 296, 339, 777, 890, 778, 801, 808, 852, 1020, 139, 820, 385, 445, 1094, 149, 116, 404, 481, 1035, 1073, 30, 66, 104, 298, 579, 642, 703, 881, 940, 966, 1058, 1070, 1078, 327, 652, 519, 1005, 1071, 1090, 518, 646, 647, 781, 856, 982]. Jack [408, 890]. Jackman [854]. Jacob [202]. Jacquart [9]. Jacques [341, 750]. Jahre [459]. Jahrhunderts [337, 361]. Jahrzehnt [81]. James [961, 806, 234, 914, 141]. Jan [309]. Janssen [343]. January [1048, 73]. Japan [783, 786, 197, 113, 1013, 1056]. Japanese [783]. Jeffersonian [784]. Jenks [1073]. Jenny [631]. Jethro [728]. Jeu [645]. Jews [179]. Jeze [746, 259, 746]. Johannes [236, 364, 834]. John [1053, 439, 806, 1021, 66, 985, 110, 375, 511, 638]. Johnson [79, 559]. Joint [33]. Jones [583, 915, 850]. Jörgensen [661]. Josef [879, 8]. Joseph [595, 699, 29, 876, 324, 383, 423]. Jourdan [1019]. Jours [31, 484, 164]. Juglar [908]. Julian [445]. Julin [11]. Julius [503, 445, 830]. July [411]. June [111, 110, 245, 249, 248, 376, 375, 512, 511, 639, 638, 741, 740, 846, 845, 938, 937, 1016, 1015, 1087, 1088]. Juraschek [57]. Juridique [102]. Jusqu’a [484, 31]. Jusqu’en [120, 321]. Just [96]. Juvenile [359].

K.C.B [74]. Kaiserlichen [81, 7]. Kantone [685]. Kapital [364]. Kapital- [364]. Kapitalanlage [85]. Kapitalistischen [68]. Karl [562, 697, 114]. Kartelle [68]. Kasseneinrichtungen [459]. Käte [648]. Kaufmann [630, 516]. Kaufpreise [168]. Kazimierz [404]. Kellogg [888, 151]. Kelynack [154]. Kemmerer [1038]. Kendall [592]. Kennan [498]. Kennedy [538]. Kent [498]. Keynes [519]. Kinds [573, 590, 613, 627, 640]. King [504, 402, 771, 773, 696]. Kingdom [556, 559, 95, 438, 514, 189, 133, 135, 1067, 738, 93, 419, 671, 288, 6, 22, 42, 58, 1013, 777, 743, 29, 679]. Kinley [1018]. Kirchhoff [24]. Kirkaldy [752, 799, 873, 999]. Kirkland [985]. Klara [307]. Klotz [351]. Knibbs [960]. Knoop [542, 389]. Knott [795]. Knox [1057, 1036]. Knudsen [631]. Kobatsch [543]. Kohlhammer [277]. Kokichi [1056]. Kollektivmasslehre [215]. Kommission [685]. Kommunalen [236]. Konferenzen [685]. Konigreich [168]. Konzentration [101]. Kossuth [498]. Kosten [277]. Kowatsch [422, 422]. Krankheit [486]. Krankheits [115]. Krankheits- [115]. Krebs [342]. Kritik [307]. Krömmelbein [277]. Kuczynski [503]. Kulturausgaben [410]. Kulturproblem [445].

L [405, 216, 218, 256, 430, 743, 778, 806, 133, 59, 144, 744, 853, 607, 351, 678, 705, 855, 940, 1018, 1040, 1070, 327, 724, 775, 913, 720, 722, 926, 202, 698]. Labor [80, 361]. Labour [122, 1105, 119, 189, 524, 884, 95, 653, 873, 1035, 800, 1051, 1055]. Labourer [592, 33, 940, 327]. Labourers [33]. Labours [697]. Lage [486].

L’Alimentation [924]. L’Allemagne [620]. Lamb [439]. Land [617, 895, 892, 579, 616, 618, 686, 705, 368, 652, 540, 119, 191, 687]. Landanteile [631]. Landauer [461]. Landliche [168]. Landowner [79]. Landry [829]. Landwirtschaft [578]. Lanfranco [1096]. L’Appauvrissement [306]. Large [335, 467, 254]. Lasker [289]. Last [288]. Latifi [406]. Latin [454]. Latine [102]. Laurence [777, 987]. Lavergne [218, 947]. Law [123, 340, 65, 310, 818, 494, 554, 698, 696, 471, 86, 830]. Lawrence [983, 1022]. Laws [32, 123, 199, 177, 876, 386, 683, 336]. Lawson [561, 255]. Lawyers [256]. Layton [653, 360]. Lead [177]. Lebens [409]. Lebens- [409]. Lebensdauer [10]. Lebensversicherung [215, 215, 661]. Lectures [104]. Lee [217]. Leedham [724]. Leener [62]. Lefeuvre [883]. Lefeuvre-Méaulle [883]. L’Effort [948]. Legal [801]. Legale [755]. Leggi [24]. Legislation [259, 501, 647, 524]. Legislative [1102]. Legislazione [706]. Lehfeldt [1072]. Leipzig [115]. Leist [1002]. L’Emploi [750]. Lennard [651]. Lenoir [482]. Léon [12]. Leroy [470, 290, 34]. Leroy-Beaulieu [290]. less [502]. Lessons [119]. L’État [472]. l’Étranger [655]. Letters [439, 697]. Leuba [914]. Leur [647]. L’Europe [308]. Levasseur [427, 250, 61, 221]. Leven [819, 980]. Levi [802]. L’Évolution [309, 580]. Levy [1022, 1071, 254, 68, 654, 220]. Lewinski [309]. Lewis [426]. Lezioni [98]. Liberty [67]. Library [129, 160, 19, 38, 53, 72, 90, 108, 128, 265, 173, 185, 195, 211, 226, 241, 286, 264, 394, 302, 315, 331, 347, 357, 372, 415, 393, 529, 550, 434, 452, 465, 477, 491, 508, 528, 669, 570, 587, 599, 611, 625, 635, 668, 764, 763, 691, 713, 736, 863, 862, 790, 812, 839, 953, 952, 900, 919, 932, 1028, 1027, 974, 992, 1010, 1110, 1045, 1063, 1083, 1109]. Lichamen [819, 980]. Lichtenfelt [277]. Liebmann [403]. Liegenschaften [143].

Life

[408, 890, 162, 903, 1031, 155, 439, 198, 771, 605, 697, 888, 742, 626, 959, 730, 904, 1078]. Life-Table [1031]. Life-Tables [198, 959]. Limit [133]. L’Impôt [96]. Lincoln [751]. Lindheim [10]. Lindsay [983]. L’Industria [657]. L’Industrie [594]. Lines [533, 109]. Lionel [1099, 987]. Liquors [419]. List [129, 160, 19, 38, 53, 72, 90, 108, 128, 265, 173, 185, 195, 211, 226, 241, 286, 264, 394, 302, 315, 331, 347, 357, 372, 415, 393, 529, 550, 434, 452, 465, 477, 491, 508, 528, 669, 570, 587, 599, 611, 625, 635, 668, 764, 763, 691, 713, 736, 863, 862, 790, 812, 839, 953, 952, 900, 919, 932, 1028, 1027, 974, 992, 1010, 1110, 1045, 1063, 1083, 1109]. Literarhistorische [322]. Lithuania [977]. Live [551]. Live-Stock [551]. Livelihood [720]. Liverpool [33]. Lives [33, 592]. Living [305, 1056, 1066, 257, 822]. Livre [1068, 1069]. LL.D. [74]. Lloyd [339]. Loafers [122]. Lobar [186]. Local [146, 222, 628, 27, 504, 163, 994, 29]. Lohne [337, 443]. Lohnsystem [443]. Loi [292]. Lois [13]. Lombard [726]. Löndlichen [337]. London [248, 375, 511, 638, 740, 845, 937, 1015, 1087, 411, 677, 110, 531]. Londres [388, 676, 676]. Lord [496, 178, 856, 1098, 766]. Lordier [619]. L’Organisation [62]. Loria [683, 657]. Loro [732, 747]. Losses [888]. Lottin [383]. Louis [620, 217, 403, 63, 654, 645]. Louise [257]. Lowell [983]. Lower [186]. Ludlow [884]. Ludwell [336]. Ludwig [410, 700]. Luigi [851]. L’Union [947]. Lutte [367]. Lutz [1002]. Luzzatti [732]. Lyman [888].

M

[325, 621, 217, 292, 340, 521, 544, 563, 658, 927, 306, 461, 486, 701, 149, 114, 522, 351, 660, 881, 1040, 1058, 519, 683, 702, 775, 11, 25]. M. [250]. M.P. [176]. Macarthur [149]. Machine [1023]. Maciejewski [362]. MacLean [167]. Made [378]. Madsen [877]. Magnitude [694]. Mai [631]. Mainly [601]. Maitland [643]. Mallet [603]. Mallock [1022, 615]. Mallon [778]. Man [895]. Management [136, 513]. Managers [256, 1021]. Manes [565]. Manfred [884]. Mangold [203]. Manual [59, 1037, 405]. Manuale [117]. Manufacturer [877]. Manufactures [1006]. Manufacturing [461]. Mappin [1059]. Marcel [482]. Marchant [961]. Marconcini [657]. Marcuse [445]. Margaretta [1004]. Maria [8]. Marion [167, 257]. Market [780, 749]. Marketing [855]. Markoff [403]. Marks [579]. Maroi [1096]. Marriage [572, 673]. Marriages [535, 574]. Marriott [652]. Marseille [880]. Marshall [946]. Martin [871, 483]. Marvaud [103, 280]. Mary [801, 896]. Massachusetts [166]. Massé [804]. Massenverbrauch [277]. Masses [306]. Material [552, 775, 307]. Mathematical [843, 868, 905, 923, 978, 847, 1094, 960]. Mathematiques [750, 629]. Mathematische [215]. Mathusianism [849]. Matthews [757, 756]. Maughfling [1021]. Maurice [13, 502, 885, 179, 497, 351]. Maurice-Bloch [351]. Max [380]. Maxime [34]. May [592, 149, 288]. Mayr [60]. McCune [983]. McVickar [566]. Mean [1014, 958, 869, 936]. Means [123]. Measurement [270, 374, 1066, 744]. Measures [122]. Measuring [348]. Measurments [577]. Méaulle [883]. Mécanique [202]. Mechanism [925]. Median [719]. Medical [163, 605, 456, 420, 154]. Medico [605]. Medico-Actuarial [605]. Medizin [700]. Meet [1071]. Meeting [111, 110, 249, 248, 376, 375, 512, 511, 639, 638, 741, 740, 846, 845, 938, 937, 1016, 1015, 1087, 1088]. Melrose [852]. Melvin [970]. Members [553]. Memoir [1104]. Memorandum [770, 45, 895]. Memorial [795]. Men [1105]. Mengarini [98]. Menschlichen [10].

Mercantile [1105]. Merchants [256, 770]. Meredith [881]. Metabolic [1023]. Method [198, 817, 939, 1050, 771, 533, 212, 402, 643]. Methode [557, 819, 980]. Methoden [684, 516, 307]. Méthodes [807]. Methodologische [322]. Methods [256, 389, 541, 847, 41, 112, 498, 702, 1017, 348, 997]. Metodologia [98]. Michel [595, 219]. Midas [1103]. Migrations [253]. Miles [217]. Military [888]. Mill [66]. Millis [783]. Milnes [1103]. Milton [1006]. Minderbemittelter [7]. Minimum [501, 754]. Ministère [321]. Minnesota [833]. Minority [305]. Mirs [631]. Misère [500]. Missiaen [306]. Mitchell [558]. Mittel [423]. MM [746]. Models [757]. Modern [778, 888, 854, 152, 702]. Modernen [10, 143]. Molesworth [47]. Molinari [297]. Molyneux [1105]. Mon [297]. Monarchs [650]. Mönchaltorf [685]. Mönckmeier [446]. Moneta [732]. Monétaire [102]. Monétaires [343]. Monetary [252, 353]. Money [963, 749, 678, 233, 1054]. Mongton [786]. Monnaie [681]. Monograph [1101, 968]. Monographie [631]. Monographiques [44]. Monographs [805]. Monopole [68]. Monopolies [728, 877]. Monopoly [361]. Montagu [222]. Monthly [19, 38, 53, 72, 90, 108, 128, 264, 393, 434, 452, 465, 477, 491, 508, 528, 570, 587, 599, 611, 625, 635, 668, 763, 862, 952, 1027]. Moore [336, 698]. Moral [469]. Morale [9, 497]. Morals [848]. Morbidity [514]. Moret [750]. Morgan [1078]. Morimoto [1056]. Morison [271]. Morman [806]. Morse [927]. Mortalità [121]. Mortalities [77]. Mortality [397, 514, 903, 253, 396, 515, 605, 853, 886, 186, 534, 334, 467, 163, 745]. Mortara [121, 180]. Mosse [486]. Mossop [272]. Mouvement [482, 237]. Movement [521, 113]. Movements [823]. Moyen [497]. Mr. [771, 772]. Mulhall [232]. Multiple [212]. Multiplier [739]. Munchener [361]. Municipal [389, 886, 757, 730]. Municipalités [442]. Murray [835, 747]. Muscular [1023]. Musgrave [829]. Mutabilità [441]. Mutualité [219].

N [472, 1097, 729, 780, 149, 154, 117, 448, 661]. Nach [275, 631]. Naissances [142]. Napier [702, 795]. Napoleone [582]. Nathan [757]. Nation [27, 946, 1065, 14, 399, 401]. National [562, 27, 64, 379, 848, 191, 252, 29, 842, 727, 1091, 1098, 146, 339, 941]. Nationalisation [542, 273]. Nationalization [786]. Natural [14]. Nature [744, 583]. Naumann [881]. Nearing [305, 774]. Nebst [307, 631]. Nécessiteux [500]. Needed [793]. nei [657, 732]. nel [747]. nella [663]. Net [351]. Neumann [337]. Nicholson [801, 30, 982]. Nicolaas [21]. Nine [886]. Nineteenth [146, 165, 998, 360, 92, 6, 22, 42, 58]. Ninth [512, 511]. niveaux [502]. Nogaro [580, 581]. Noms [190]. Norges [97]. Normal [818, 494, 227]. Norman [913]. North [701, 559]. Northampton [720]. nos [31, 484, 164]. Note [870, 162, 673, 816, 1032, 772, 574, 717, 719, 437, 935, 696]. Notes [17, 36, 51, 70, 88, 106, 126, 158, 171, 183, 193, 209, 224, 239, 262, 284, 300, 313, 329, 345, 355, 370, 391, 413, 468, 432, 450, 463, 475, 489, 506, 526, 548, 568, 585, 597, 609, 623, 633, 666, 689, 711, 734, 761, 788, 810, 837, 860, 898, 917, 930, 950, 972, 990, 1008, 1025, 1043, 1061, 1081, 1107, 651, 251, 377, 1031, 771, 55, 674, 93, 4, 555]. Notices [190]. Notre [829]. Nouveaux [362]. November [631, 132, 672, 866, 956]. Noyes [909]. Number [20].

O [907, 306, 460, 887]. O’Brien [1033]. Observations [981]. Occupation [77]. Occupations [612, 76, 231, 674, 556, 23, 770]. Occupying [540]. Occurrence [908]. Ocean [708]. Œuvres [424]. O’Farrell [707]. Offences [288]. Offered [910]. Office [867]. Officer [163, 1077]. Officers [246]. Official [541, 867, 480]. Officials [256]. Officiella [278]. Officielle [97]. Offspring [114]. Ogburn [524]. Ogle [350]. Old [269, 135, 166, 777]. Olney [708]. One [198]. op [819, 980]. Operating [272, 92]. Operation [379, 886, 805, 518, 1099, 593, 607, 429, 1020]. Operations [29]. Operative [551, 521, 987]. Operazioni [706]. Opinion [538]. Opportunities [902]. Options [153]. Order [553, 830]. ordinated [749]. Organica [747]. Organisation [428, 1049, 677, 68, 865, 987, 278]. Organisierte [685]. Organization [204, 723, 1077]. Organized [884]. Orient [454]. Origines [755]. Origins [784]. Orlo [439]. Ormerod [1036]. Ortskrankenkasse [115]. Orwin [704, 986]. Osgood [1004]. Other [16, 35, 50, 69, 87, 105, 125, 156, 169, 181, 192, 208, 223, 238, 260, 282, 299, 312, 328, 344, 354, 369, 390, 412, 431, 449, 462, 473, 487, 505, 525, 547, 567, 584, 596, 608, 622, 632, 665, 688, 710, 733, 760, 787, 809, 836, 859, 897, 916, 929, 949, 971, 989, 1007, 1024, 1042, 1060, 1080, 1106, 256, 716, 678, 133]. Oualid [580]. Our [325, 895, 1001, 921, 726, 803, 963, 1003]. Out-Patients [213]. Outbreak [675, 844]. Outline [27, 941]. Outlines [542, 298]. Outlook [912, 447, 928]. Output [378]. Ouvrage [297]. Ouvrier [352]. Ouvrière [13, 502]. Ouvrieres [292, 1069]. Ouvriers [592, 676]. Over-Production [386]. Overseas [738]. Owned [774]. Ownership [686, 540]. Oxford [513].

P [272, 141, 544, 891, 1100, 294, 753, 939, 1023, 150, 151, 606, 686, 98, 893, 856]. Paillard [102]. Paish [1103]. Palin [25].

Panama [751, 454]. Panics [908]. Panjab [406]. Pankraz [485]. Paper [1105]. Papers [927, 411, 1092, 281]. Paralysis [968]. Parental [114]. Pareto [554]. Parliament [1052]. Parmelee [885]. Part [24, 886, 703, 117, 92, 6, 22, 42, 58]. Particular [253]. Partie [221]. Partnership [607]. Parts [768, 769, 935]. Pasquet [676]. Past [45, 539, 731]. Patients [213]. Paul [218, 290, 470, 883, 367, 683]. Pauperism [131]. Pawn [104]. Pay [910]. Payen [523]. Payment [1055]. Pays [138, 834]. Peace [895, 945, 1034, 1070]. Pearson [697, 114]. Peddie [781]. Pendant [948, 829]. Pension [886]. Pensions [777, 135, 166, 1074]. People [64, 773, 1047, 770, 124]. Peoples [775]. Percival [1053]. Percy [152]. Perils [803]. Periodical [54, 196, 316, 332, 392, 435, 453, 466, 571, 588, 600, 692, 714, 791, 813, 901, 920, 975, 993, 1046, 1064, 908]. Periodicals [18, 37, 52, 71, 89, 107, 127, 159, 172, 184, 194, 210, 225, 240, 263, 285, 301, 314, 330, 346, 356, 371, 414, 433, 451, 464, 476, 490, 507, 527, 549, 569, 586, 598, 610, 624, 634, 667, 690, 712, 735, 762, 789, 811, 838, 861, 899, 918, 931, 951, 973, 991, 1009, 1026, 1044, 1062, 1082, 1108]. Périodiques [558]. Peris [307]. Peter [207, 759, 143]. Pethick [1022]. Petite [44]. Pfarrer [685]. Pfitzner [236]. Phenomena [744]. Phénomenès [1068]. Philip [1037, 86]. Philippovich [879]. Physiocrates [31]. Physiocratique [237]. Physique [114, 78]. Picard [292]. Pierre [235, 292, 164, 825]. Pierson [448, 472, 21]. Pigou [425]. Pilgrimage [604]. Pinot [292]. Place [388]. Placement [388]. Planning [709]. Plunkett [155]. Pneumonia [186]. Poincaré [403, 643]. Point [895, 630]. Points [149]. Poland [977]. Policy [878, 538, 65]. Poliomyelitis [968]. Political [524, 1075, 66, 86, 298, 660]. Politicians [687]. Politico [747].

Politico-Economico [747]. Politics

[756, 946]. Politique [206, 46, 280, 543, 581, 1068, 1069, 746, 619, 629, 387, 750]. Poole [430]. Poor [32, 123, 199, 340, 65, 649]. Popolazioni [180]. Population [24, 891, 841, 136, 612, 436, 967, 230, 694, 717, 113, 869, 936, 960, 470, 849]. Populations [212]. Porter [401]. Portion [717]. Portugez [887]. Position [995, 287]. Positive [82]. Possibilities [406]. Possible [574]. Post [212]. Post-Censal [212]. Postage [984]. Postal [1038, 1038]. Potts [927]. Poverty [335, 720, 885]. Povo [887]. Powell [593, 749]. Power [233, 857]. Powers [1085]. Pox [141]. Practical [256]. Practice [428, 29, 310, 915, 826]. Pramathanath [295]. Pratique [292]. Prato [352]. Pratt [273, 428, 542, 857]. Pre [977]. Pre-War [977]. Précis [387, 11]. Preference [104]. Preis [143]. Preisbewegung [277]. Preliminary [242, 1018]. Premier [1068]. Première [221]. Preparation [252]. Present [256, 296, 112, 539, 731]. Present-Day [256]. Presented [110, 248, 375, 511, 638, 1087, 740, 845, 937, 1015]. President [246]. Preussen [168, 307]. Preventable [382]. Prevention [728]. Prévoyance [219]. Prévoyange [321]. Prezzi [663, 732]. Price [349, 318, 696, 607, 361, 722]. Prices [906, 979, 305, 140, 76, 231, 360, 267, 481, 556, 361, 682, 985, 602, 695, 794, 43, 188, 319, 455, 1072, 1054]. Primer [827, 25]. Principal [76, 231, 177]. Principalement [630]. Principes [659]. Principi [747]. Principii [1093]. Principle [583]. Principles [407, 428, 806, 722, 389, 365, 66, 448, 925]. Pringle [123]. Prinsen [731]. Prinzing [889]. Private [1017]. Prix [120, 482]. Probabilités [403, 26, 403]. Probabilities [162, 416, 847]. Probable [835]. Problem [122, 891, 155, 988, 783, 1091, 913, 1073, 965]. Problème [500]. Problemes [746].

Problems [1041, 256, 191, 895, 1049, 131, 945, 73, 786, 1034, 467, 976, 1103, 281, 946, 203, 222, 411, 664]. Proceedings [132, 111, 245, 249, 376, 417, 512, 639, 741, 672, 846, 938, 866, 1016, 956, 1048, 1088, 1039]. Prochaine [522]. Production [138, 478, 386, 378, 438, 696, 606]. Products [855]. Produktion [461]. Produktivgenossenschaften [364]. Produttiva [121]. Produzione [24]. Professioni [545]. Professions [190]. Professors [133]. Profit [943]. Profit-Sharing [943]. Profits [705, 1012, 682]. Program [148]. Progress [562, 64, 401, 716, 911, 253, 174, 818, 671, 92, 885, 67, 524]. Progressif [63]. Progresso [582]. Project [30]. Projetée [523]. Promotions [656]. Property [774, 1002, 719, 820, 721]. Proportion [534]. Proposals [686]. Proposed [1071]. Propriété [44, 120]. Propriétés [351]. Prospects [957]. Protection [755]. Protectionism [879]. Protectionnisme [352, 366, 879]. Prothero [539]. Provident [716]. Provisional [229]. Prussia [146]. Prussian [842]. Psychical [1050]. Psychological [914]. Pubblico [706]. Public [256, 782, 848, 941, 471, 1053]. Public-Law [471]. Publications [16, 35, 50, 69, 87, 105, 125, 156, 169, 181, 192, 208, 223, 238, 260, 282, 299, 312, 328, 344, 354, 369, 390, 412, 431, 449, 462, 473, 487, 505, 525, 547, 567, 584, 596, 608, 622, 632, 665, 688, 710, 733, 760, 787, 809, 836, 859, 897, 916, 929, 949, 971, 989, 1007, 1024, 1042, 1060, 1080, 1106, 252]. Publics [220, 34]. Publiek [819, 980]. Publiek-Rechtelijke [819, 980]. Publique [338]. Pulsford [962]. Punto [24]. Pupin [874]. Pura [747]. Purchasing [233]. Pure [659]. Purpose [744].

Quantity [481]. Question [470, 103, 969, 758]. Questions [1069].

Quetelet [497, 383].

R [561, 401, 503, 563, 255, 560, 753, 912, 133, 140, 460, 149, 754, 748, 279, 566, 558, 944, 118, 652, 661, 724, 913, 83, 518, 720, 1072]. Race [320, 888, 179]. Racial [281]. Radford [1003]. Raffalovich [1041]. Raffety [27]. Rail [857]. Railroad [621]. Railroads [723, 723]. Railway [405, 1101, 49, 542, 1077, 1102, 272, 1077, 724, 786]. Railways [561, 358, 542, 785, 786, 273]. Rankin [801]. Ranks [303]. Raoul [142]. Raphael [220, 200]. Raphael-Georges [220]. Rapid [817]. Rappard [755]. Rapporti [657, 732]. Ratan [1076]. Rate [349, 848, 418, 457, 591, 718, 317, 641, 318, 961]. Rate.Part [701]. Rates [186, 76, 231, 754, 556, 75, 723, 724, 984, 904, 768, 769, 935, 1101]. Rationalisierung [445]. Ratios [94]. Rawlley [1076]. Ray [722]. Raymond [926]. Raynaud [501]. Re [1077]. Re-Organization [1077]. Reader [852, 821]. Reading [493, 720]. Readings [884]. Real [531]. Réalisée [523]. Réalités [882, 969]. Realities [615]. Reception [246]. Recherches [502]. Rechtelijke [819, 980]. Reconstruction [987, 1103]. Record [124, 987]. Recording [626]. Recruiting [303]. Reduction [56]. Reference [709, 1074, 1023, 360, 902, 915, 1021, 842, 386]. Reform [327, 615, 324, 148]. Refugees [792]. Regard [829]. Reginald [651, 430]. Register [792]. Registrar [535]. Registrar-General [535]. Registration [865]. Réglementation [523]. Règlements [655]. Regression [533]. Regular [469]. Regulation [1102, 723]. Reiche [444, 7, 83]. Reiches [423, 485, 100, 520]. Reichs [7, 485]. Reichs-Arbeitsblatte [7]. Reichversicherungsordnung [460]. Related [615]. Relating [416, 686, 574]. Relation [256, 136, 903, 1030, 396, 335, 785, 913, 56, 781, 1013]. Relations [146, 653, 721, 287, 637, 783]. Relative [305]. Relief [32, 123, 199]. Relieving [123]. Remede [290]. Remédes [142, 218]. Remedies [803]. Renard [964]. Rendered [774]. René [874]. Rents [76, 231, 556]. Reorganisation [867]. Reorganisations [656]. Repartition [24, 116, 581]. Répercussions [964]. Répertoire [190]. Replacement [1105]. Reply [892, 978]. Report [628, 110, 248, 375, 397, 511, 514, 638, 740, 845, 937, 1015, 1087, 80, 123, 305, 617, 535, 892, 478, 33, 166, 835, 616, 378, 1017, 438, 133, 614, 1095, 1105, 189, 701, 886, 76, 231, 967, 556, 480, 518]. Reported [20]. Reporting [41]. Reports [227]. Represent [573, 590, 613, 627, 640]. Representation [843, 868, 905, 923, 978, 3, 73]. Research [1050, 33, 910, 382, 1037]. Researches [334]. Resources [200, 902]. Responsibility [14]. Ressources [924]. Restraint [122]. Résultats [351]. Resulting [889]. Results [242, 45, 256, 379, 777, 873, 3, 73, 498, 1002, 707]. Retail [305, 76, 231, 556]. Retard [92]. Retort [618]. Retraites [292]. Returns [54, 196, 316, 332, 435, 453, 466, 571, 588, 600, 692, 714, 791, 813, 901, 920, 975, 993, 1046, 1064, 774, 869, 936]. Réunis [880]. Revendications [647]. Revenu [351, 96]. Revenus [472]. Review [85, 99, 100, 146, 147, 165, 206, 207, 222, 235, 236, 272, 273, 275, 277, 308, 309, 325, 337, 343, 363, 380, 381, 405, 409, 410, 427, 442, 443, 444, 472, 517, 520, 561, 562, 565, 620, 621, 628, 759, 998, 1041, 1092, 1097, 1104, 13, 24, 27, 32, 45, 60, 62, 64, 80, 121, 122, 123, 124, 141, 142, 148, 199, 216, 217, 218, 256, 271, 281, 289, 290, 292, 296, 305, 339, 340, 379, 389, 401, 402, 406, 408, 420, 440, 470, 500, 501, 502, 503, 504, 521, 541, 544, 563, 576, 582, 603, 614, 617, 649, 651, 657, 658, 663, 706, 709, 729, 732, 747, 757, 776, 777, 782, 796, 848, 849, 851, 873, 885, 890, 891, 907, 927, 961, 1038, 1039, 1053]. Review [1056, 1074, 1079, 1093, 1095, 1099, 1100, 1105, 364, 15, 31, 44, 65, 119, 138, 155, 167, 189, 191, 233, 254, 255, 257, 294, 306, 407, 428, 430, 439, 447, 461, 560, 595, 653, 662, 721, 743, 753, 756, 758, 778, 780, 801, 803, 805, 806, 808, 824, 852, 895, 911, 912, 943, 946, 963, 987, 1001, 1004, 1020, 139, 163, 203, 253, 341, 499, 535, 593, 699, 1101, 49, 542, 629, 892, 939, 59, 61, 97, 120, 140, 144, 145, 214, 220, 278, 310, 311, 365, 425, 485, 592, 681, 797, 820, 1057, 1076, 14, 81, 115, 326, 342, 362, 382, 384, 385, 411, 445, 459, 460, 486, 524, 546, 577, 605, 644, 648, 697, 700, 701, 744, 745, 847, 853, 886, 888, 889, 896]. Review [942, 968, 981, 988, 1023, 1037, 1094, 33, 48, 76, 77, 149, 150, 151, 166, 200, 231, 232, 360, 677, 725, 742, 754, 755, 835, 967, 1075, 7, 116, 685, 748, 1077, 1102, 26, 114, 154, 215, 252, 403, 404, 29, 153, 205, 353, 388, 481, 482, 556, 607, 999, 63, 68, 96, 279, 321, 338, 361, 471, 522, 566, 594, 656, 676, 728, 730, 752, 804, 823, 830, 831, 876, 908, 915, 924, 970, 1006, 1021, 1035, 1073, 12, 30, 34, 46, 47, 66, 67, 86, 103, 104, 152, 204, 221, 234, 237, 258, 259, 280, 297, 298, 324, 351, 352, 366, 383, 386, 387, 426, 483, 484, 523, 543, 558, 564, 579, 580, 581, 606, 616, 618, 642, 655, 659, 660, 664, 678, 682]. Review [686, 687, 703, 704, 705, 707, 726, 749, 779, 784, 799, 800, 802, 807, 821, 832, 833, 850, 855, 878, 879, 881, 882, 883, 884, 940, 944, 945, 947, 962, 965, 966, 985, 986, 1000, 1003, 1018, 1019, 1033, 1040, 1051, 1055, 1058, 1059, 1068, 1069, 1070, 1078, 79, 118, 164, 178, 327, 368, 480, 539, 540, 604, 652, 654, 746, 84, 98, 102, 117, 180, 190, 201, 291, 367, 378, 441, 448, 498, 519, 538, 545, 557, 559, 583, 615, 630, 643, 661, 679, 680, 683, 702, 723, 724, 727, 750, 751, 773, 774, 775, 783, 785, 786, 795, 798, 822, 825, 854, 857, 874, 875, 877, 893, 909, 910, 914, 928, 941, 948, 964, 983, 984, 997, 1002, 1005, 1017, 1022, 1034, 1052, 1054]. Review [1071, 1089, 1090, 1091, 1096, 1103, 731, 913, 274, 578, 8, 9, 10, 11, 25, 28, 82, 83, 95, 101, 143, 168, 219, 276, 293, 295, 307, 322, 323, 422, 423, 429, 446, 518, 536, 537, 619, 631, 646, 647, 684, 708, 720, 722, 781, 819, 826, 827, 828, 829, 834, 856, 858, 880, 894, 925, 926, 969, 980, 982, 1036, 1072, 1098, 288, 78, 179, 202, 336, 421, 424, 438, 458, 497, 516, 645, 650, 698, 960, 887]. Révision [351]. Revolution [755, 671, 522]. Rew [560]. Rhondda [1078]. Ricchezza [84]. Ricci [796]. Richard [871, 721, 684, 722]. Richesse [874]. Rickman [439]. Riesser [101]. Right [176, 496, 872, 766]. Ripley [723, 723]. Rise [428, 749, 857]. Rising [1054]. Rist [31]. Ritter [57]. Robert [24, 74, 576, 927, 342, 835, 779, 799, 945, 583, 1034]. Roberto [747]. Robertson [823, 1058]. Robins [687]. Robinson [1103, 340, 521, 833]. Roger [858]. Rogers [148]. Rolla [1006]. Rollo [382]. Rooms [110, 248, 375, 511, 638, 740, 845, 937, 1015, 1087]. Rossignol [234]. Rothamsted [966]. Rothkegel [168]. Roulleau [655]. Rowe [1018]. Rowland [539]. Rowntree [289, 592, 119]. Royal [967, 110, 248, 375, 511, 638, 740, 845, 938, 937, 1016, 1015, 1087, 1088, 32, 123, 199]. Rubinow [658]. Rudolf [543]. Rufus [927]. Rule [45]. Rumpf [460]. Rural [612, 155, 806, 967, 437, 987, 616]. Rurale [44]. Rusch [380]. Russell [808, 742, 1020, 382, 966, 643]. Russia [1041, 1097, 902, 1041, 1097]. Russland [631]. Rutter [294].

S [1092, 24, 141, 339, 849, 780, 1101, 49, 144, 835, 656, 1035, 298, 704, 986, 1018, 1040, 630]. Sa [142, 24]. Sacker [630]. Saggio [747]. Sagnier [293]. Sakolski [621]. Salaire [501]. Salaires [120, 321, 219]. Salariata [657]. Saluti [10]. Sample [1014, 958]. Samples [553]. sämtlicher [685]. Samuel [895, 983]. Sanderson [928]. Sargant [1037]. Sartorius [85]. Sauvaire [1019]. Sauvaire-Jourdan [1019]. Savill [618]. Savings [1038]. Schädigungen [648].

Schapiro [595]. Scharlieb [896]. Schedule [577]. Schelle [366]. Schemes [216, 709]. Schiffahrtsstatisik [517]. Schmid [685]. Schmidt [759]. Scholars [154]. School [78]. Schools [154]. Schott [485]. Schulden [236]. Schumacher [323]. Schweiz [277, 759, 685, 207]. Schweizerischen [685]. Science [256, 873, 398, 524, 1075, 259, 650, 398, 856, 643]. Scienza [747, 851]. Scope [939]. Scotland [967, 287, 572, 673, 123, 1053]. Scott [305, 774, 1034, 133, 799, 945]. Scottish [400]. Scourge [896]. Sea [112]. Seager [148]. Séailles [116]. Seasonal [228]. Second [846, 845, 850]. Secondo [121]. Secret [705]. Secretary [133]. Secrist [997]. Sect [320]. Section [873, 398, 92]. Seebohm [289, 119, 592]. seine [445]. seit [485, 361, 307]. Selbständige [557]. Selection [888]. Self [717]. Self-Governing [717]. Seligman [566, 279]. Sella [84]. Sénéchal [12]. Senectutis [10]. Sense [86]. Séparation [9]. Serbian [1065]. Series [541]. Service [886, 894]. Services [34, 774]. Ses [142, 523, 403]. Sessi [24]. Session [532]. Sessional [110, 248, 375, 511, 638, 740, 845, 937, 1015, 1087]. Sesso [24]. Seven [149]. Seventh [249, 248]. Seventy [111, 110, 249, 248, 376, 375, 512, 511, 535]. Seventy-Eighth [376, 375]. Seventy-Fourth [535]. Seventy-Ninth [512, 511]. Seventy-Seventh [249, 248]. Seventy-Sixth [111, 110]. Sevin [410]. Sex [139]. Sexes [24, 534]. Sexuallebens [445]. Shackleton [149]. Shares [153]. Sharing [943]. Sharp [27]. Shield [801, 30, 982]. Shirts [1105]. Short [198, 771, 118, 722]. Shortt [1018]. Show [213]. Shrinivas [748]. Shrinking [822]. Sibships [553]. Sickness [467]. Sidney [146, 504, 65, 583, 910]. Sigmund [485]. Silk [1076]. Similar [5]. Simpler [775]. Since [123, 675, 317, 844]. Sir [176, 175, 871, 74]. Sistemazione [747]. Sisyphusarbeit [82]. Situation [235, 142, 895, 647]. Sixth [111, 110]. Size [553]. Sizes [601]. Skatternes [97]. Sketch [722]. Slav [1065]. Small [254, 79, 141]. Small-Pox [141]. Smaller [137, 1]. Smart [660, 850, 165, 998, 384]. Smelser [1100]. Smith [987, 149, 1099, 808, 984, 1034]. Smith-Gordon [987, 1099]. So-Called [823]. Sociaal [819, 980]. Sociaal-Economisch [819, 980]. Sociaes [887]. Social [1097, 124, 148, 289, 885, 927, 1039, 14, 744, 1075, 703, 775, 148, 379, 658, 677, 615]. Sociale [500, 103, 647, 306, 321]. Sociales [219]. Socialism [234]. Socialists [64]. Socialstatistik [97]. Societies [521, 467, 502, 716]. Society [110, 375, 511, 638, 416, 605, 844, 110, 248, 375, 511, 638, 740, 845, 938, 937, 1016, 1015, 1087, 892, 411, 1088, 677]. Sociologue [383]. Some [45, 521, 359, 835, 674, 334, 641, 686, 552, 73, 197, 467, 4, 64, 92, 758]. Sommaires [190]. son [290]. Sonderheft [7]. Sophonisba [150]. Soudan [235]. Source [574]. South [468, 858, 304]. Southern [1065]. Soward [1089]. Sowie [276]. Sozial [10]. Sozial-Statistische [10]. Soziale [486, 700]. Soziologie [440]. Spalding [826]. Special [397, 514, 709, 1074, 886, 1023, 360, 902, 915, 842, 386, 829]. Specimen [1021]. Spencer [257]. Spiller [281]. Spirit [521]. Springer [503]. St [309]. Staate [10]. Staats [207]. Staats- [207]. Stamp [820]. Stand [99, 275]. Standard [1056, 227]. Ständigen [685]. Standpunkte [342]. Stanley [559, 720]. Staple [1036]. Staples [987]. State [649, 1105, 112, 877, 1002, 256, 963, 942, 785, 1090, 234]. Statement [673, 816]. States [305, 294, 231, 831, 1006, 152, 783, 274, 252, 1067, 147, 1038, 155, 461, 943, 797, 835, 471, 908, 884, 773, 774, 80]. Statistica [98, 117, 98]. Statistical [110, 18, 37, 52, 71, 89, 107, 127, 159, 248, 247, 172, 184, 194, 210, 225, 240, 263, 285, 375, 301, 314, 330, 346, 356, 371, 392, 414, 511, 433, 451, 464, 476, 490, 507, 527, 549, 638, 569, 586, 598, 610, 624, 634, 667, 740, 690, 712, 735, 762, 845, 789, 811, 838, 861, 938, 937, 899, 918, 931, 951, 1016, 1015, 973, 991, 1009, 1026, 1087, 1044, 1062, 1082, 1108, 402, 398, 793, 1065, 867, 843, 868, 905, 923, 1050, 382, 577, 847, 988, 131]. Statistical [112, 1088, 934, 842, 268, 914, 997, 532, 467, 4, 288, 336]. Statistician [1017, 1017]. Statisticien [383]. Statistico [24]. Statistics [358, 514, 873, 253, 699, 1030, 867, 977, 416, 573, 590, 613, 627, 640, 978, 39, 304, 1037, 359, 232, 112, 865, 955, 480, 197, 109, 996, 92, 6, 22, 42, 58, 272, 405, 939, 59, 214, 847, 1094, 970, 25, 576]. Statistiek [819, 980, 201, 537]. Statistik [275, 381, 97, 278, 215, 422, 536, 440, 342, 536, 516, 380, 60, 459, 7]. Statistique [796, 138, 9, 11, 497, 362, 619, 321, 421]. Statistiques [219]. Statistische [557, 10, 819, 980]. Statistischen [81, 7]. Stebbing [893]. Steel [831]. Stephan [277]. Steps [650]. Sterbetafeln [81]. Sterblichkeitsverhältnisse [115]. Sterile [546]. Stevens [785]. Stewart [234]. Stiftung [83]. Still [396]. Still-Births [396]. Stock [388, 551, 29]. Stocks [153]. Storia [499]. Storico [98]. Story [504]. Strand [248, 740, 845, 937, 1015, 1087]. Street [110, 375, 511, 638, 726]. Streightoff [471]. Stuart [339, 66, 980, 201, 537, 819]. Students [256, 621]. Studie [546]. Studien [323]. Studies [471, 1018]. Study [147, 1097, 64, 122, 216, 420, 849, 1038, 612, 575, 254, 257, 912, 946, 1065, 572, 14, 1023, 186, 360, 114, 823, 1006, 86, 606, 749, 884, 327, 552, 683, 798, 822, 1002, 295, 720, 828, 925, 179, 289, 890, 385, 577, 988, 914, 984]. Stufenkarten [422]. Stuk [819, 980]. Subsequent [903]. Substitution [1074]. Success [124]. Successful [705]. Successive [904]. Sugar [731]. Suggested [806, 198, 467, 109]. Suggestions [626]. Suisse [102, 755]. Summary [112]. Superficies [138]. Supplément [13].

Supplies [957]. Supply [803, 386, 682, 399]. Suret [63]. Surproduction [558]. Survey [363, 406, 1065, 131, 1103, 561]. Sveriges [278]. Svolgimento [98]. Swan [835]. Sweating [367]. Sweating-System [367]. Switzerland [144]. Sydney [703]. Syndicale [62]. Syndicalisme [522]. Syndicats [62, 219, 34]. Synthesis [683]. System [85, 1038, 55, 679, 727, 983, 422, 652, 367]. Systems [333, 802].

T [216, 776, 461, 653, 721, 745, 360, 154, 970, 1021, 579, 749, 1051, 775, 854, 722, 142]. Table [1031]. Tables [198, 771, 959, 1021]. Tailoresses [912]. Tailoring [754]. Taken [553]. Taker [439]. Tally [269]. Tamburini [98]. Tarbell [325]. Tariff [147, 325, 758, 152, 324, 878]. Tariffism [1058]. Tarifs [258]. Tarifvertrage [443]. Tätigkeit [82]. Taussig [758, 365]. Tawney [753, 912, 754, 720]. Tax [133, 1002, 279]. Taxation [835, 994, 686, 498, 538, 1089, 1047, 27, 566, 583, 1002, 628, 1090]. Taxation-Methods [498]. Taxpayer [1052]. Taylor [339, 781]. Teachers [256]. Techniker [409]. Technologique [190]. Teilbau [631]. Teilbaugrossbetriebs [631]. Teleky [700]. Ten [716]. Tendances [206]. Tendencies [253]. Tendency [479]. Tennant [149]. Tenure [686]. Teoria [117]. Teorica [747]. Tercentenary [702, 795]. Terrace [248, 740, 845, 937, 1015, 1087]. Terrestre [24]. Test [494]. Teubner [215, 215]. Text [621]. Text-Book [621]. Textile [601, 693]. Th [460]. Theilhaber [546]. Their [358, 296, 777, 479, 386, 513, 785, 698, 721, 908, 1013]. Them [177, 533, 960]. Theodore [271]. Theoreon [77]. Theorie [362, 661, 26, 1068, 497, 516, 63, 564]. Theorique [292]. Théoriques [796]. Theory [662, 416, 214, 847, 29, 826, 960, 481].

Theresa [801]. Theresias [8]. Thereto [686]. Théry [308, 338]. these [288]. Third [938, 937, 80]. Thom [908]. Thomas [253, 686, 850]. Thompson [884, 849]. Thomson [310]. Thought [426]. Thoughts [850]. Three [3, 985]. Throughout [853]. Time [1032, 800]. Times [325]. Tirman [292]. To-day [798]. Tobacco [1105, 877]. Todd [324]. Toepasing [980]. Toepassing [819]. Together [76, 231, 556, 1017, 722]. Tollemache [540]. Toshiharn [786]. tot [201, 537, 819, 980]. Totals [229]. Tougan [595]. Tougan-Baranowsky [595]. tous [120]. Town [940, 709]. Towns [76, 231, 556]. Trade [1041, 521, 1100, 189, 743, 76, 231, 754, 454, 556, 922, 558, 1000, 1055, 552, 510, 738, 751, 1012, 828, 829, 1098, 1013, 95, 1076, 823, 104, 204, 324, 708, 925, 856]. Trader [877]. Trades [479, 95, 1036, 151]. Trading [389]. Traffic [708, 724]. Traite [292]. Tramps [122]. Transition [271]. Transport [49, 752]. Transportation [854, 925]. Transports [580]. Trasmissione [545]. Travail [755, 321, 523, 1069]. Travailleurs [352]. Travis [724]. Treatment [456]. Trend [841]. Trésors [220]. trois [630]. Troupeau [804]. Trust [1073, 876]. Trusts [68]. Tryon [1006]. Tuberkulose [342]. Tuckwell [149]. Tuesday [741, 846, 938, 1016, 1088]. Tugendreich [486]. Turner [895]. Turnor [191, 803]. Turrentine [1040]. Twelfth [430, 780]. Twenty [253, 288, 80]. Twenty-Third [80]. Two [3, 348]. Types [469].

U [151]. Überseeische [446].

Übervölkerung [28]. Udny [576, 214]. Ueber [700, 404]. Ukas [631]. Ultima [297]. Umberto [796]. Umgegend [115]. unbebauter [143]. Uneheliche [145]. Unemployables [122]. Unemployed [123]. Unemployment [359, 134, 386, 216, 289, 1100]. Unfair [876].

Unfavourable [78]. Uniformity [524]. Uninsured [456]. Union [102, 468]. Unionism [1055]. Unions [521, 1100]. United [147, 80, 305, 777, 1038, 155, 189, 294, 461, 743, 943, 797, 231, 835, 252, 29, 556, 471, 831, 908, 1006, 152, 884, 559, 679, 773, 774, 783, 274, 95, 438, 514, 133, 135, 1067, 738, 93, 419, 671, 288, 6, 22, 42, 58, 1013]. University [411, 524, 1075]. Unofficial [892]. Unserer [445]. Untersuchung [423, 10, 322]. Untersuchungen [115]. Unwin [1000]. Upon [541, 1004, 771, 772, 1018, 1012, 1074]. Urban [137, 39, 976, 1, 437, 617, 967]. Ursachen [423, 361]. Usages [388]. Use [416, 573, 590, 613, 627, 640, 227, 1037].

V [912, 97, 460, 894, 92]. Vaccination [141]. Vagrancy [122]. Valeur [564]. Valeurs [388]. Valides [500]. Valori [732]. Valuation [642]. Value [1014, 958, 385]. Vandyke [833]. Variabilità [441]. Variations [437, 288]. Various [805, 498, 1017, 1047, 768, 769, 935]. Végétale [138]. Verba [297]. Vergleichenden [342]. Verhältnis [24]. Vernon [888]. Verrijn [201, 537, 819, 980]. Vers [501]. Versicherungsanstalten [459]. Versicherungswesen [565]. Verwaltungs [364]. Verwaltungs-Beteiligung [364]. Vesselitsky [912]. Viande [804]. Vicomte [120]. Victoria [801]. vie [502, 61]. View [895, 149]. Views [305]. VIII [97]. Village [327]. Viscount [946, 664, 1078]. Vista [24]. Vita [121, 84]. Vital [39, 865, 955, 842, 109, 699]. Vitalité [1019]. Vogel [517]. Vol [305, 614, 617, 616, 960, 123]. Volksinkomen [980]. Volksrijksdom [980]. Volkswirtschaftliche [85, 277]. Volkszählung [276]. Volkszählungen [8]. Volkwirtschaftliche [546]. Volume [2, 91, 130, 243, 266, 373, 395, 509, 530, 636, 670, 737, 765, 840, 864, 933, 954, 1011, 1029, 1084, 847, 795]. Voluntary [213]. vom [337, 342, 631]. Vorarlberg [144]. Voraussichtliche [99]. Vorlesungen [700]. Vorträge [323, 83]. vue [630]. Vyrnwy [1078].

W [277, 561, 176, 27, 124, 296, 339, 401, 873, 907, 1038, 255, 294, 653, 758, 911, 946, 1001, 133, 365, 14, 326, 755, 353, 999, 728, 752, 908, 1073, 66, 324, 580, 799, 881, 118, 746, 615, 854, 877, 983, 1022, 1089, 1091, 25, 429, 829, 858]. W.C [937, 1015, 1087]. W.C. [248, 375, 511, 638, 740, 845]. Wage [1105, 257, 801, 303, 92, 167]. Wage-Earners [1105, 303]. Wage-Earning [167]. Wagel [748]. Wages [305, 1086, 133, 76, 231, 481, 556, 703, 92, 6, 22, 42, 58, 531, 336, 1055, 651, 1022]. Wagner [484]. Wahrscheinlichkeitsrechnung [403, 215]. Wahrscheinlichkeitstheorie [215]. Wales [137, 841, 612, 163, 674, 694, 1, 23, 768, 769, 770, 935, 551, 535]. Wall [726]. Walter [939, 360, 742, 726, 168]. Waltershausen [85]. Walther [517]. Wanderungsgeschichte [446]. Wang [1102]. War [675, 873, 1074, 743, 895, 946, 1004, 977, 1032, 842, 707, 726, 800, 1018, 814, 857, 910, 1017, 1022, 1071, 1098, 844, 767, 693, 715, 793, 955, 922, 842, 957, 738, 671, 895, 1035, 799, 832, 945, 746, 909, 1034, 946, 726, 982]. War-Time [1032]. Warenshandelsbetrieb [381]. Warren [849]. Warrington [720]. Wars [888, 889]. Washington [1039]. Watarai [786]. Waterways [725]. Watson [146]. Wattal [891]. Wealth [562, 64, 739, 642, 615, 773, 1017, 1085, 721, 425]. Webb [504, 65, 146, 504, 544, 65, 232, 583, 910]. Weber [500]. Weekly [1105]. Weiler [207]. Welby [766]. Weld [855]. Welfare [256, 191, 425, 1005]. Welton [253]. Weltwirtschaftliche [323]. Wert [143]. Wertschätzung [82]. Wesley [558]. West [1094]. Westel [983]. Westergaard [939, 888, 889]. Weulersse [237]. Wheat [965, 985, 56, 294]. Wheeler [775]. Whetham [326, 946, 14, 326]. which [177, 288, 960]. Whitney [147]. Whittaker [686]. Wholesale [906, 979, 305, 361, 140]. Wicker [722]. Wicksteed [86]. Wieth [631]. Wieth-Knudsen [631]. Wild [685]. Wilhelm [520, 446]. Willan [1089]. Willcox [939]. Willford [402, 773]. William [165, 998, 1097, 350, 122, 379, 1074, 946, 310, 384, 524, 847, 942, 730, 234, 660, 726, 799, 850, 945, 965, 538, 723, 983, 1034, 926]. Williams [439, 49]. Willoughby [983, 894, 1091]. Wilson [104]. Winkelmann [648]. Wirkungen [423]. Wirtschaftliche [276]. Wirtschafts [536]. Wirtschafts-Finanz [536]. Wirtschaftsrechnungen [7]. Wissenschaft [557]. Withers [726, 824, 963]. within [4]. Witwatersrand [1072]. Wl [629]. Wm [826]. Wolf [445]. Wolff [443, 124, 1001, 429]. Woman [815, 149]. Women [1105, 257, 1004, 894, 1086, 167, 778, 150, 151]. Woodbury [927]. Woods [650]. Woodworking [95]. Wool [147, 1036]. Work [1023, 677, 678, 642, 703]. Workers [257, 469]. Working [777, 493, 76, 231, 556, 720]. Working-Class [493, 76, 556, 720]. Workingmen [217]. Workmen [80]. Workpeople [189]. Works [742]. World [957, 731, 913, 853, 200, 725]. Wotzel [448]. Wright [147]. Wurm [100].

XIX [92]. xixa [123]. XV [6, 22]. XVII [42]. XVIII [58].

Year [110, 248, 375, 511, 638, 740, 845, 937, 1015, 1087, 198, 830]. Years [45, 873, 716, 756, 253, 904, 288]. Yield [56]. York [886, 968, 835, 1105]. Young [776, 816, 727]. Yule [576, 214]. Yves [442, 594, 969]. Yves-Guyot [969].

Z [723]. Zahn [275, 276]. Zawadzki [629]. Zealand [801, 234]. Zeit [631, 445]. Zivilrechtspflege [380]. Zizek [440]. Zolla [15]. Zollpolitik [520]. Zon [200]. Zones [880]. Zugelassene [459]. zum [7]. zur [444, 68, 82, 423, 446, 291]. Zusammenhange [101]. Zwillingsgeburten [404].

References

Welton:1904:SUD

[1] Thomas A. Welton. On the smaller urban districts of England and Wales. Journal of the Royal Statistical Society, 67(1):1–51, March 1904. ISSN 0952-

                                              8385. URL http://www.jstor.org/

stable/2339334. See errata [137].

Anonymous:1910:VIa

[2] Anonymous. Volume information. Journal of the Royal Statistical Society, 73(1):i–viii, January 1910. ISSN

                                              0952-8385. URL http://www.jstor.
                                              org/stable/2340008.

Garnett:1910:RCE

[3] W. Garnett. On the representation of certain examination results in two and in three dimensions. Journal of the Royal Statistical Society, 73 (1):1–9, January 1910. ISSN 0952-

                                              8385. URL http://www.jstor.org/
                                              stable/2340009.

Webb:1910:NSD

[4] Augustus D. Webb. Notes on some difficulties met within international statistical comparisons. Journal of the Royal Statistical Society, 73(1):

10–25, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340010.

Yule:1910:DDA

[5] G. Udny Yule. On the distribution of deaths with age when the causes of death act cumulatively, and similar frequency distributions. Journal of the Royal Statistical Society, 73(1): 26–38, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340011.

Wood:1910:SWUa

[6] George Henry Wood. The statistics of wages in the United Kingdom during the Nineteenth Century. (Part XV). The cotton industry. Journal of the Royal Statistical Society, 73(1): 39–58, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340012.

J:1910:BRBa

[7] H. S. J. Book review: Erhebung von Wirtschaftsrechnungen Minderbemittelter Familien im Deutschen Reiche. 2^{tes} Sonderheft zum Reichs-Arbeitsblatte. Bearbeitet im Kaiserlichen Statistischen Amte; Abteilung für Arbeiter-Statistik. Journal of the Royal Statistical Society, 73(1): 59–65, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340013.

W:1910:BRBa

[8] A. D. W. Book review: Die Volkszählungen Maria Theresias und Josef II, 1753–1790, by Alfred Gürtler. Journal of the Royal Statistical Society, 73(1):65–66, January 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340014.

W:1910:BRBb

[9] A. D. W. Book review: Essais de Statistique Morale. II. Le Divorce et la Séparation de Corps, by Camille Jacquart. Journal of the Royal Statistical Society, 73(1):66–67, January 1910. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340015.

W:1910:BRBc

[10] A. D. W. Book review: Saluti Senectutis. Die Bedeutung der Menschlichen Lebensdauer im Modernen Staate. Eine Sozial-Statistische Untersuchung, by Alfred von Lindheim. Journal of the Royal Statistical Society, 73(1): 67–68, January 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340016.

W:1910:BRBd

[11] A. D. W. Book review: Précis du Cours de Statistique Générale et Appliquée, by Armand Julin; M. A. de Foville. Journal of the Royal Statistical Society, 73(1):68–69, January 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340017.

P:1910:BRBa

[12] L. L. P. Book review: La Concentration Industrielle et Commerciale en Angleterre, by Léon Sénéchal. Journal of the Royal Statistical Society, 73 (1):69–71, January 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340018.

B:1910:BRBa

[13] E. B. Book review: Les Lois D’Assurance Ouvrière à L’Étranger.

Supplément Général, by Maurice Bellom. Journal of the Royal Statistical Society, 73(1):71–72, January 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340019.

G:1910:BRBa

[14] M. G. Book review: The Family and the Nation: A Study in Natural Inheritance and Social Responsibility, by W. C. D.; Catherine Durning Whetham. Journal of the Royal Statistical Society, 73(1):72–73, January 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340020.

C:1910:BRBa

[15] P. G. C. Book review: Le Blé et les Céréales, by Daniel Zolla; L’Ecole de Grignon. Journal of the Royal Statistical Society, 73(1):74–76, January 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340021.

Anonymous:1910:ONPa

[16] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(1):77–79, January 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340022.

Anonymous:1910:CNa

[17] Anonymous. Current notes. Journal of the Royal Statistical Society, 73 (1):80–86, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340023.

Anonymous:1910:SEAa

[18] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73 (1):87–89, January 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340024.

Anonymous:1910:MLAa

[19] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(1):90–93, January 1910. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340025.

Verney:1910:RCI

[20] H. Verney. On the recent considerable increase in the number of reported accidents in factories. Journal of the Royal Statistical Society, 73(2): 95–125, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340152.

Stuart:1910:NGP

[21] C. A. Verrijn Stuart. Nicolaas Gerard Pierson. Journal of the Royal Statistical Society, 73(2):126–127, February 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340153.

Wood:1910:SWUb

[22] George Henry Wood. The statistics of wages in the United Kingdom during the Nineteenth Century. (Part XV). The cotton industry. Journal of the Royal Statistical Society, 73(2): 128–163, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340154.

Welton:1910:OEW

[23] Thomas A. Welton. Occupations in England and Wales, 1881 and 1901. Journal of the Royal Statistical Society, 73(2):164–166, February 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340155.

B:1910:BRBb

[24] J. A. B. Book review: Il Sesso dal Punto di Vista Statistico. Part I. Le Leggi Della Produzione dei Sessi, by Corrado Gini; Population du Globe Terrestre et sa Repartition par Sexes by S. Goulichambaroff; Über das Verhältnis der Geschlechter in Indien by Robert Kirchhoff. Journal of the Royal Statistical Society, 73(2):167– 170, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340156.

W:1910:BRBe

[25] A. D. W. Book review: Primer of Statistics, by W. Palin Elderton; Ethel M. Elderton. Journal of the Royal Statistical Society, 73(2):170–171, February 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340157.

K:1910:BRBa

[26] J. M. K. Book review: Éléments de la Théorie des Probabilités, by Emile Borel. Journal of the Royal Statistical Society, 73(2):171–172, February 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340158.

B:1910:BRBc

[27] A. L. B. Book review: The Nation’s Income. An Outline of National and Local Taxation, by F. W. Raffety; W. H. Cartwright Sharp. Journal of the Royal Statistical Society, 73(2):172, February 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340159.

W:1910:BRBf

[28] A. D. W. Book review: Die Übervölkerung Deutschlands und ihre Bekämpfung, by Ferdinand Goldstein. Journal of the Royal Statistical Society, 73(2):173, February 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340160.

L:1910:BRBa

[29] W. R. L. Book review: Stock Exchange Investments in Theory and Practice, with Chapters on the Constitution and Operations of the Bank of England and the National and Local Debts of the United Kingdom, by Joseph Burn. Journal of the Royal Statistical Society, 73(2):174–175, February 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340161.

P:1910:BRBb

[30] L. L. P. Book review: A Project of Empire, by J. Shield Nicholson. Journal of the Royal Statistical Society, 73(2): 175–177, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340162.

C:1910:BRBb

[31] E. C. Book review: Histoire des Doctrines Économiques depuis les Physiocrates Jusqu’à nos Jours, by Charles Gide; Charles Rist. Journal of the Royal Statistical Society, 73(2):177– 179, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340163.

B:1910:BRBd

[32] E. B. Book review: Royal Commission on the Poor Laws and Relief of Distress. Journal of the Royal Statistical Society, 73(2):179–182, February 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340164.

H:1910:BRBa

[33] B. L. H. Book review: How the Causal Labourer Lives. Report of the Liverpool Joint Research Committee on the Domestic Condition and Expenditure of the Families of Certain Liverpool Labourers. Journal of the Royal Statistical Society, 73(2):182–183, February 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340165.

P:1910:BRBc

[34] L. L. P. Book review: Syndicats et Services Publics, by Maxime Leroy. Journal of the Royal Statistical Society, 73(2):184–185, February 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340166.

Anonymous:1910:ONPb

[35] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(2):186–187, February 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340167.

Anonymous:1910:CNb

[36] Anonymous. Current notes. Journal of the Royal Statistical Society, 73(2): 188–199, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340168.

Anonymous:1910:SEAb

[37] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73(2): 200–203, February 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340169.

Anonymous:1910:MLAb

[38] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(2):204–206, February 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340170.

Flux:1910:UVS

[39] A. W. Flux. Urban vital statistics in England and Germany. Journal of the Royal Statistical Society, 73(3): 207–253, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339848.

deFoville:1910:EC

[40] A. de Foville. Émile Cheysson. Journal of the Royal Statistical Society, 73 (3):254–256, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339849.

Godfrey:1910:MCR

[41] Ernest H. Godfrey. Methods of crop reporting in different countries. Journal of the Royal Statistical Society, 73 (3):257–283, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339850.

Wood:1910:SWUc

[42] George Henry Wood. The statistics of wages in the United Kingdom during the Nineteenth Century. (Part XVII). The cotton industry. Journal of the Royal Statistical Society, 73 (3):283–315, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339851.

Sauerbeck:1910:PC

[43] A. Sauerbeck. Prices of commodities in 1909. Journal of the Royal Statistical Society, 73(3):315–329, March 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339852.

C:1910:BRBc

[44] P. G. C. Book review: La Petite Propriété Rurale en France. Enquêtes Monographiques (1908–09). Journal of the Royal Statistical Society, 73(3): 330–333, March 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339853.

B:1910:BRBe

[45] J. A. B. Book review: Memorandum on Some of the Results of Indian Administration During the Past Fifty Years of British Rule in India. Journal of the Royal Statistical Society, 73 (3):334–335, March 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339854.

P:1910:BRBd

[46] L. L. P. Book review: Le Conflit des Doctrines dans l’Économie Politique Contemporaine, by Charles Brouilhet. Journal of the Royal Statistical Society, 73(3):335–337, March 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339855.

P:1910:BRBe

[47] L. L. P. Book review: Economic and Fiscal Facts and Fallacies, by Guilford Molesworth. Journal of the Royal Statistical Society, 73(3):337–339, March 1910. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339856.

H:1910:BRBb

[48] D. H. Book review: Essays in Eugenics, by Francis Galton. Journal of the Royal Statistical Society, 73(3):339, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339857.

E:1910:BRB

[49] C. L. E. Book review: The Economics of Railway Transport, by S. C. Williams. Journal of the Royal Statistical Society, 73(3):340–341, March 1910. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2339858.

Anonymous:1910:ONPc

[50] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(3):341–343, March 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339859.

Anonymous:1910:CNc

[51] Anonymous. Current notes. Journal of the Royal Statistical Society, 73 (3):344–355, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339860.

Anonymous:1910:SEAc

[52] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73 (3):356–358, March 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339861.

Anonymous:1910:MLAc

[53] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(3):359–361, March 1910. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2339862.

Anonymous:1910:PR

[54] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 73 (3):362–379, March 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339863.

Hooper:1910:NFS

[55] Wynnard Hooper. Notes on the financial system of the German empire. Journal of the Royal Statistical Society, 73(4):381–395, April 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340412.

Vigor:1910:IYP

[56] H. D. Vigor. The increased yield per acre of wheat in England considered in relation to the reduction of the area. Journal of the Royal Statistical Society, 73(4):396–408, April 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340413.

Craigie:1910:FRJ

[57] P. G. Craigie. Franz Ritter von Juraschek. Journal of the Royal Statistical Society, 73(4):409–410, April 1910. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340414.

Wood:1910:SWUd

[58] George Henry Wood. The statistics of wages in the United Kingdom during the Nineteenth Century. (Part XVIII). The cotton industry. Journal of the Royal Statistical Society, 73 (4):411–434, April 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340415.

F:1910:BRBa

[59] A. W. F. Book review: An Elementary Manual of Statistics, by Arthur L. Bowley. Journal of the Royal Statistical Society, 73(4):435–437, April 1910.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340416.

B:1910:BRBf

[60] A. L. B. Book review: Statistik und Gesellschaftslehre, by George von Mayr. Journal of the Royal Statistical Society, 73(4):437–439, April 1910. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340417.

F:1910:BRBb

[61] A. W. F. Book review: Le Coût de la Vie, by E. Levasseur. Journal of the Royal Statistical Society, 73 (4):439–440, April 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340418.

B:1910:BRBg

[62] E. B. Book review: L’Organisation Syndicale des Chefs d’Industrie. Étude sur les Syndicats Industriels en Belgique, by Georges de Leener. Journal of the Royal Statistical Society, 73 (4):440–441, April 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340419.

M:1910:BRBa

[63] B. M. Book review: Théorie de l’Impôt Progressif, by Louis Suret. Journal of the Royal Statistical Society, 73 (4):441–442, April 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340420.

B:1910:BRBh

[64] A. L. B. Book review: The People’s Progress. A Study of the Facts of National Wealth, with some Answers to Socialists, by Frank Ireson. Journal of the Royal Statistical Society, 73 (4):442–445, April 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340421.

C:1910:BRBd

[65] W. C. Book review: English Poor Law Policy, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 73(4):445–447, April 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340422.

P:1910:BRBf

[66] L. L. P. Book review: Principles of Political Economy, by John Stuart Mill; W. J. Ashley. Journal of the Royal Statistical Society, 73(4):447–449, April 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340423.

P:1910:BRBg

[67] L. L. P. Book review: Liberty and Progress, by C. Y. C. Dawbarn. Journal of the Royal Statistical Society, 73 (4):449–451, April 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340424.

M:1910:BRBb

[68] H. W. M. Book review: Monopole, Kartelle, und Trusts in ihren Beziehungen zur Organisation der Kapitalistischen Industrie, by Hermann Levy. Journal of the Royal Statistical Society, 73(4):451–453, April 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340425.

Anonymous:1910:ONPd

[69] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(4):454–456, April 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340426.

Anonymous:1910:CNd

[70] Anonymous. Current notes. Journal of the Royal Statistical Society, 73 (4):457–463, April 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340427.

Anonymous:1910:SEAd

[71] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73 (4):464–467, April 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340428.

Anonymous:1910:MLAd

[72] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(4):468–472, April 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340429.

Rosenbaum:1910:GEJ

[73] S. Rosenbaum. The general election of January, 1910, and the bearing of the results on some problems of representation. Journal of the Royal Statistical Society, 73(5):473–528, May 1910. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340171.

B:1910:SRG

[74] A. E. B. Sir Robert Giffen, K.C.B., F.R.S., LL.D. Journal of the Royal Statistical Society, 73(5):529–533, May 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340172.

Pearson:1910:CDR

[75] Karl Pearson, Alice Lee, and Ethel M. Elderton. On the correlation of deathrates. Journal of the Royal Statistical Society, 73(5):534–539, May 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340173.

H:1910:BRBc

[76] E. J. H. Book review: Report of an Enquiry, by the Board of Trade into Working-Class Rents, Housing and Retail Prices, Together with the Rates of Wages in Certain Occupations in the Principal Industrial Towns of Belgium. Journal of the Royal Statistical Society, 73(5):540–544, May 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340174.

H:1910:BRBd

[77] N. A. H. Book review: Occupation Mortalities, with an Abstract of Discussion Theoreon. Journal of the Royal Statistical Society, 73(5):544–547, May 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340175.

Y:1910:BRB

[78] G. U. Y. Book review: The Influence of Defective Physique and Unfavourable Home Environment on the Intelligence of School Children, by David Heron. Journal of the Royal Statistical Society, 73(5):547–551, May 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340176.

R:1910:BRBa

[79] R. H. R. Book review: The Disappearance of the Small Landowner, by Arthur H. Johnson. Journal of the Royal Statistical Society, 73(5): 552–555, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340177.

B:1910:BRBi

[80] E. B. Book review: Twenty-Third Annual Report of the Commissioner of Labor, 1908. Workmen’s Insurance and Benefit Funds in the United States. Journal of the Royal Statistical Society, 73(5):555–557, May 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340178.

G:1910:BRBb

[81] M. G. Book review: Deutsche Sterbetafeln für das Jahrzehnt 1891 bis 1900. Bearbeitet im Kaiserlichen Statistischen Amte. Journal of the Royal Statistical Society, 73(5): 557–558, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340179.

W:1910:BRBg

[82] A. D. W. Book review: Sisyphusarbeit oder Positive Erfolge? Beiträge zur Wertschätzung der Tätigkeit der Deutschen Gewerkschaften. Journal of the Royal Statistical Society, 73 (5):558–559, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340180.

W:1910:BRBh

[83] A. D. W. Book review: Vorträge der Gehe-Stiftung zu Dresden. Band II. 1910. Beruf, Gesellschaftliche Gliederung und Betrieb im Deutschen Reiche, by R. van der Borght. Journal of the Royal Statistical Society, 73(5):559, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340181.

S:1910:BRBa

[84] S. S. Book review: La Vita della Ricchezza, by Emanuele Sella. Journal of the Royal Statistical Society, 73 (5):559–560, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340182.

A:1910:BRBa

[85] P. A. Book review: Das Volkswirtschaftliche System der Kapitalanlage im Auslande, by A. Sartorius Freiherr von Waltershausen. Journal of the Royal Statistical Society, 73 (5):560–562, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340183.

P:1910:BRBh

[86] L. L. P. Book review: The Common Sense of Political Economy, Including a Study of the Human Basis of Economic Law, by Philip H. Wicksteed. Journal of the Royal Statistical Society, 73(5):562–565, May 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340184.

Anonymous:1910:ONPe

[87] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(5):565–568, May 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340185.

Anonymous:1910:CNe

[88] Anonymous. Current notes. Journal of the Royal Statistical Society, 73 (5):569–577, May 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340186.

Anonymous:1910:SEAe

[89] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73 (5):578–581, May 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340187.

Anonymous:1910:MLAe

[90] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(5):582–584, May 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340188.

Anonymous:1910:VIb

[91] Anonymous. Volume information. Journal of the Royal Statistical Society, 73(6/7):833–848, June/July 1910. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339903.

Wood:1910:SWN

[92] George Henry Wood. The statistics of wages in the Nineteenth Century. Part XIX. — The cotton industry. Section V. Changes in the average wage of all employed, with some account of the forces operating to accelerate or retard the progress of the industry. Journal of the Royal Statistical Society, 73(6/7): 585–633, June/July 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339904.

Strutt:1910:NDE

[93] H. C. Strutt. Notes on the distribution of estates in France and the United Kingdom. Journal of the Royal Statistical Society, 73(6/7):634–644, June/ July 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339905.

Yule:1910:ICB

[94] G. Udny Yule. On the interpretation of correlations between indices or ratios. Journal of the Royal Statistical Society, 73(6/7):644–647, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339906.

W:1910:BRBi

[95] G. H. W. Book review: Earnings and Hours Inquiry. Report…, by the Board of Trade. … Earnings and Hours of Labour… in the United Kingdom. III. Building and Woodworking Trades in 1906. Journal of the Royal Statistical Society, 73(6/7):648–650, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339907.

M:1910:BRBc

[96] B. M. Book review: L’Impôt sur le revenu, by Just Haristoy. Journal of the Royal Statistical Society, 73(6/7): 651–652, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339908.

F:1910:BRBc

[97] A. W. F. Book review: Skatternes Fordeling Efter Indtoegt og Forsörgelsesbyrde. Norges Officielle Statistik V. 104. Socialstatistik VIII. Journal of the Royal Statistical Society, 73(6/7):652–653, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339909.

S:1910:BRBb

[98] S. S. Book review: Lezioni di Statistica. I. Metodologia Statistica. II. Svolgimento Storico Della Statistica, by Augusto Bosco; P. Mengarini; A. Tamburini. Journal of the Royal Statistical Society, 73(6/7):653–654, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339910.

A:1910:BRBb

[99] P. A. Book review: Die Deutsche Fleischproduktion, Ihr Gegenwärtiger Stand und Ihre Voraussichtliche Entwickelung, by Heinrich Gerlich. Journal of the Royal Statistical Society, 73 (6/7):654–656, June/July 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339911.

A:1910:BRBc

[100] P. A. Book review: Die Finanzgeschichte des Deutschen Reiches, by Emanuel Wurm. Journal of the Royal Statistical Society, 73(6/7):656– 657, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339912.

W:1910:BRBj

[101] A. D. W. Book review: Die Deutschen Grossbanken und Ihre Konzentration im Zusammenhange mit der Entwicklung der Gesamtwirtschaft in Deutschland, by Dr. Riesser. Journal of the Royal Statistical Society, 73(6/7):657– 659, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339913.

S:1910:BRBc

[102] J. S. Book review: La Suisse et l’Union Monétaire Latine. Etude Économique et Juridique, by Georges Paillard. Journal of the Royal Statistical Society, 73 (6/7):659, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339914.

P:1910:BRBi

[103] L. L. P. Book review: La Question Sociale en Espagne, by Angel Marvaud. Journal of the Royal Statistical Society, 73(6/7):659–661, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339915.

P:1910:BRBj

[104] L. L. P. Book review: An Empire in Pawn: Being Lectures and Essays on Indian, Colonial, and Domestic Finance, Preference, Free Trade, by A. J. Wilson. Journal of the Royal Statistical Society, 73(6/7):661–663, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339916.

Anonymous:1910:ONPf

[105] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(6/7):663–665, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339917.

Anonymous:1910:CNf

[106] Anonymous. Current notes. Journal of the Royal Statistical Society, 73(6/7): 666–676, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339918.

Anonymous:1910:SEAf

[107] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73(6/7): 677–680, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339919.

Anonymous:1910:MLAf

[108] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(6/7):681–684, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339920.

Stevenson:1910:SLA

[109] T. H. C. Stevenson. Suggested lines of advance in English vital statistics. Journal of the Royal Statistical Society, 73(6/7):685–713, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339921.

Anonymous:1910:RCF

[110] Anonymous. Report of the Council for the financial year ended December 31, 1909, and for the sessional year ending June 21, 1910, presented at the seventy-sixth Annual General Meeting of the Royal Statistical Society, held at the rooms of the Royal Society of Arts, John Street, Adelphi, London, on June 21, 1910. Journal of the Royal Statistical Society, 73(6/7):714–725, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339922.

Anonymous:1910:PSS

[111] Anonymous. Proceedings of the seventy-sixth Annual General Meeting, held on June 21, 1910. Journal of the Royal Statistical Society, 73(6/7): 726–727, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339923.

Henking:1910:SFS

[112] Dr. Henking. The sea fishery statistics of Germany: A summary of an address on the present state of the sea fishery statistics of Germany and of the statistical methods employed in international sea investigation. Journal of the Royal Statistical Society, 73(6/7):

728–738, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339924.

Takano:1910:RMP

[113] Iwasaburo Takano. The recent movement of population in Japan. Journal of the Royal Statistical Society, 73 (6/7):738–768, June/July 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339925.

K:1910:BRBb

[114] J. M. K. Book review: A First Study of the Influence of Parental Alcoholism on the Physique and Ability of the Offspring, by Ethel M. Elderton; Karl Pearson. Journal of the Royal Statistical Society, 73(6/7):769–773, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339926.

G:1910:BRBc

[115] M. G. Book review: Krankheits- und Sterblichkeitsverhältnisse in der Ortskrankenkasse für Leipzig und Umgegend. Untersuchungen über den Einfluss von Geschlecht, Alter und Beruf. Journal of the Royal Statistical Society, 73(6/7):774–776, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339927.

J:1910:BRBb

[116] H. S. J. Book review: La Répartition des Fortunes en France, by J. Séailles. Journal of the Royal Statistical Society, 73(6/7):776–777, June/July 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339928.

S:1910:BRBd

[117] S. S. Book review: Teoria Statistica, Generale e Demografica, by Aldo Contento; Manuale di Demografia (Part II of Statistica e Demografia), by N. Colajanni. Journal of the Royal Statistical Society, 73(6/7):778–779, June/ July 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339929.

R:1910:BRBb

[118] R. H. R. Book review: A Short History of English Agriculture, by W. H. R. Curtler. Journal of the Royal Statistical Society, 73(6/7):779–781, June/ July 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339930.

C:1910:BRBe

[119] P. G. C. Book review: Land and Labour: Lessons from Belgium, by B. Seebohm Rowntree. Journal of the Royal Statistical Society, 73(6/7):781– 785, June/July 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339931.

F:1910:BRBd

[120] A. W. F. Book review: Histoire Économique de la Propriété, des Salaires, des Denrées et de tous les Prix en Général Depuis l’an 1200 Jusqu’en l’an 1800, by Le Vicomte G. d’Avenel. Journal of the Royal Statistical Society, 73(6/7):785, June/July 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339932.

B:1910:BRBj

[121] J. A. B. Book review: La Mortalità Secondo L’Eta, e la Durata Della Vita Economicamente Produttiva, by G. Mortara. Journal of the Royal Statistical Society, 73(6/7):785–787, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339933.

B:1910:BRBk

[122] E. B. Book review: The Vagrancy Problem: The Case for Measures of Restraint for Tramps, Loafers, and Unemployables; with a Study of Continental Detention Colonies and Labour Houses, by William Harbutt Dawson. Journal of the Royal Statistical Society, 73 (6/7):787, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339934.

B:1910:BRBl

[123] E. B. Book review: Royal Commission on the Poor Laws and Relief of Distress. Appendix, vol. xixa. Report on the Effects of Employment or Assistance Given to the “Unemployed” Since 1886 as a Means of Relieving Distress Outside the Poor Law in Scotland, by J. C. Pringle. Journal of the Royal Statistical Society, 73(6/7):787–788, June/ July 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339935.

B:1910:BRBm

[124] E. B. Book review: People’s Banks; A Record of Social and Economic Success, by Henry W. Wolff. Journal of the Royal Statistical Society, 73(6/7): 788–789, June/July 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339936.

Anonymous:1910:ONPg

[125] Anonymous. Other new publications. Journal of the Royal Statistical Society, 73(6/7):790–792, June/July 1910.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339937.

Anonymous:1910:CNg

[126] Anonymous. Current notes. Journal of the Royal Statistical Society, 73(6/7): 793–800, June/July 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339938.

Anonymous:1910:SEAg

[127] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 73(6/7): 801–806, June/July 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339939.

Anonymous:1910:MLAg

[128] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 73(6/7):807–810, June/ July 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339940.

Anonymous:1910:ALA

[129] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 73(6/7):811–831, June/ July 1910. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339941.

Anonymous:1910:VIc

[130] Anonymous. Volume information. Journal of the Royal Statistical Society, 74(1):i–viii, December 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339977.

Hamilton:1910:SSP

[131] George Hamilton. A statistical survey of the problems of pauperism. Journal of the Royal Statistical Society, 74 (1):1–34, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339978.

Anonymous:1910:PN

[132] Anonymous. Proceedings on November 15, 1910. Journal of the Royal Statistical Society, 74(1):34–36, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339979.

Cannan:1910:ADI

[133] E. Cannan, A. L. Bowley, F. Y. Edgeworth, H. B. Lees Smith, and W. R. Scott. The amount and distribution of income (other than wages) below the income tax exemption limit in the United Kingdom. Report of the Committee of the British Association, consisting of Professors E. Cannan (Chairman), A. L. Bowley (Secretary), F. Y. Edgeworth, and H. B. Lees Smith, and Dr. W. R. Scott. Journal of the Royal Statistical Society, 74(1): 37–66, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339980.

Anonymous:1910:ICU

[134] Anonymous. The International Conference on Unemployment, 1910. Journal of the Royal Statistical Society, 74 (1):67–70, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339981.

H:1910:OAP

[135] N. A. H. Old age pensions in the United Kingdom. Journal of the Royal Statistical Society, 74(1):71–73, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339982.

Baker:1910:PCR

[136] C. Ashmore Baker. Population and costs in relation to city management. Journal of the Royal Statistical Society, 74(1):73–79, December 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339983.

Anonymous:1910:ESU

[137] Anonymous. Errata: On the smaller urban districts of England and Wales. Journal of the Royal Statistical Society, 74(1):79, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339984. See [1].

C:1910:BRBf

[138] P. G. C. Book review: Institut International d’Agriculture: Statistique des Superficies Cultivées, de la Production Végétale et du Bétail dans les Pays Adhérents. Journal of the Royal Statistical Society, 74(1):80–83, December 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339985.

D:1910:BRB

[139] R. D. Book review: Age Incidence, Sex, and Comparative Frequency in Disease, by J. Grant Andrew. Journal of the Royal Statistical Society, 74 (1):83–84, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339986.

F:1910:BRBe

[140] A. W. F. Book review: Wholesale Prices in Canada, 1890–1909, by R. H. Coats. Journal of the Royal Statistical Society, 74(1):84–85, December 1910. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339987.

B:1910:BRBn

[141] J. A. B. Book review: Small-Pox and Vaccination in British India, by S. P. James. Journal of the Royal Statistical Society, 74(1):85–87, December 1910. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339988.

B:1910:BRBo

[142] J. A. B. Book review: Les Naissances en France. Sa Situation. Ses Conséquences. Ses Causes. Existe-t-il des Remédes?, by Raoul de Félice. Journal of the Royal Statistical Society, 74(1):88–89, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339989.

W:1910:BRBk

[143] A. D. W. Book review: Wert und Preis unbebauter Liegenschaften in der Modernen Grossstadt, by Heinrich Peter. Journal of the Royal Statistical Society, 74(1):89–90, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339990.

F:1910:BRBf

[144] A. W. F. Book review: The Cotton Industry in Switzerland, Vorarlberg, and Italy, by S. L. Besso. Journal of the Royal Statistical Society, 74 (1):90–92, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339991.

F:1910:BRBg

[145] A. W. F. Book review: Die Eheliche und Uneheliche Fruchtbarkeit mit Besonderer Berücksichtigung Bayerns. Journal of the Royal Statistical Society, 74(1):92–93, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339992.

A:1910:BRBd

[146] P. A. Book review: National and Local Finance: A Review of the Relations Between the Central and Local Authorities in England, France, Belgium, and Prussia During the Nineteenth Century, by J. Watson Grice; Sidney Webb. Journal of the Royal Statistical Society, 74(1):93–95, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339993.

A:1910:BRBe

[147] P. A. Book review: Wool Growing and the Tariff: A Study in the Economic History of the United States, by Chester Whitney Wright. Journal of the Royal Statistical Society, 74(1): 95–97, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339994.

B:1910:BRBp

[148] E. B. Book review: Social Insurance: A Program of Social Reform, by Henry Rogers Seager. Journal of the Royal Statistical Society, 74(1):98, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339995.

H:1910:BRBe

[149] B. L. H. Book review: Woman in Industry, from Seven Points of View, by G. M. Tuckwell; C. Smith; M. R. Macarthur; May Tennant; N. Adler; A. M. Anderson; C. Black; D. J. Shackleton. Journal of the Royal Statistical Society, 74(1):99, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339996.

H:1910:BRBf

[150] B. L. H. Book review: Women in Industry, by Edith Abbott; Sophonisba P. Breckinridge. Journal of the Royal Statistical Society, 74(1):99–100, December 1910. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339997.

H:1910:BRBg

[151] B. L. H. Book review: Women and the Trades, by Eliz. B. Butler; P. U. Kellogg. Journal of the Royal Statistical Society, 74(1):100–102, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339998.

P:1910:BRBk

[152] L. L. P. Book review: Modern Tariff History. Germany, United States, France, by Percy Ashley. Journal of the Royal Statistical Society, 74(1): 102–103, December 1910. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339999.

L:1910:BRBb

[153] W. R. L. Book review: Arbitrage in Bullion, Coins, Bills, Stocks, Shares, and Options, by Henry Deutsch. Journal of the Royal Statistical Society, 74 (1):103–104, December 1910. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340000.

K:1910:BRBc

[154] J. K. Book review: Medical Examination of Schools and Scholars, by T. N. Kelynack. Journal of the Royal Statistical Society, 74(1):104–105, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340001.

C:1910:BRBg

[155] P. G. C. Book review: The Rural Life Problem of the United States, by Horace Plunkett. Journal of the Royal Statistical Society, 74(1):105– 106, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340002.

Anonymous:1910:ONPh

[156] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(1):106–113, December 1910. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340003.

Keynes:1910:C

[157] J. M. Keynes. Correspondence. Journal of the Royal Statistical Society, 74(1): 114–121, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340004.

Anonymous:1910:CNh

[158] Anonymous. Current notes. Journal of the Royal Statistical Society, 74(1): 122–142, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340005.

Anonymous:1910:SEAh

[159] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74(1): 143–153, December 1910. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340006.

Anonymous:1910:LAL

[160] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(1):154–166, December 1910. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340007.

Paish:1911:GBC

[161] George Paish. Great Britain’s capital investments in individual colonial and foreign countries. Journal of the Royal Statistical Society, 74(2): 167–200, January 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339835.

Brownlee:1911:NCP

[162] John Brownlee and R. M. Morison. Note on the calculation of the probabilities of life at high ages. Journal of the Royal Statistical Society, 74(2): 201–206, January 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339836.

D:1911:BRBa

[163] R. D. Book review: Infant and Child Mortality, by Medical Officer of the Local Government Board of England and Wales. Journal of the Royal Statistical Society, 74(2):207–208, January 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339837.

R:1911:BRBa

[164] R. H. R. Book review: La Crise et L’Evolution de l’Agriculture en Angleterre de 1875 a nos Jours, by Pierre Besse. Journal of the Royal Statistical Society, 74(2):208–210, January 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339838.

A:1911:BRBa

[165] P. A. Book review: Economic Annals of the Nineteenth Century, 1801– 20, by William Smart. Journal of the Royal Statistical Society, 74(2): 210–212, January 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339839.

H:1911:BRBa

[166] N. A. H. Book review: Massachusetts– Report of the Commission on Old Age Pensions, Annuities and Insurance. Journal of the Royal Statistical Society, 74(2):212–216, January 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339840.

C:1911:BRBa

[167] C. E. C. Book review: Wage-Earning Women, by Annie Marion MacLean. Journal of the Royal Statistical Society, 74(2):216, January 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339841.

W:1911:BRBa

[168] A. D. W. Book review: Die Kaufpreise fur Landliche Besitzungen im Konigreich Preussen von 1895 bis 1906, by Walter Rothkegel. Journal of the Royal Statistical Society, 74(2): 216–217, January 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339842.

Anonymous:1911:ONPa

[169] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(2):217–220, January 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339843.

Pearson:1911:C

[170] Karl Pearson. Correspondence. Journal of the Royal Statistical Society, 74 (2):221–229, January 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339844.

Anonymous:1911:CNa

[171] Anonymous. Current notes. Journal of the Royal Statistical Society, 74(2): 230–238, January 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339845.

Anonymous:1911:SEAa

[172] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74(2): 239–244, January 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339846.

Anonymous:1911:LALa

[173] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(2):245–249, January 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339847.

Julin:1911:EPB

[174] Armand Julin. The economic progress of Belgium from 1880 to 1908. Journal of the Royal Statistical Society, 74(3): 251–313, February 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340441.

Anonymous:1911:SFG

[175] Anonymous. Sir Francis Galton, D.C.L., D.Sc., F.R.S. Journal of the Royal Statistical Society, 74(3): 314–320, February 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340442.

Anonymous:1911:RHS

[176] Anonymous. The Right Hon. Sir C. W. Dilke, Bart., M.P. Journal of the Royal Statistical Society, 74(3): 320–321, February 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340443.

Keynes:1911:PAL

[177] J. M. Keynes. The principal averages and the laws of error which lead to them. Journal of the Royal Statistical Society, 74(3):322–331, February 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340444.

R:1911:BRBb

[178] R. H. R. Book review: Commons, Forests, and Footpaths, by Lord Eversley. Journal of the Royal Statistical Society, 74(3):332–333, February 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340445.

Y:1911:BRBa

[179] G. U. Y. Book review: The Jews: A Study of Race and Environment, by Maurice Fishberg. Journal of the Royal Statistical Society, 74(3): 333–334, February 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340446.

S:1911:BRBa

[180] S. S. Book review: Le Popolazioni delle Grandi Città Italiani, by Giorgio Mortara. Journal of the Royal Statistical Society, 74(3):335, February 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340447.

Anonymous:1911:ONPb

[181] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(3):335–338, February 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340448.

Keynes:1911:C

[182] J. M. Keynes. Correspondence. Journal of the Royal Statistical Society, 74(3): 339–345, February 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340449.

Anonymous:1911:CNb

[183] Anonymous. Current notes. Journal of the Royal Statistical Society, 74(3): 346–357, February 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340450.

Anonymous:1911:SEAb

[184] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74(3): 358–361, February 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340451.

Anonymous:1911:LALb

[185] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(3):362–364, February 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340452.

Greenwood:1911:FFL

[186] M. Greenwood and R. H. Candy. The fatality of fractures of the lower extremity and of lobar pneumonia. a study of hospital mortality rates, 1751–1901. Journal of the Royal Statistical Society, 74(4):365–405, March 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340401.

Anonymous:1911:II

[187] Anonymous. Industrial insurance. Journal of the Royal Statistical Society, 74(4):406–407, March 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340402.

Sauerbeck:1911:PC

[188] A. Sauerbeck. Prices of commodities in 1910. Journal of the Royal Statistical Society, 74(4):408–420, March 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340403.

C:1911:BRBb

[189] P. G. C. Book review: Report of an Inquiry, by the Board of Trade into the Earnings and Hours of Labour of Workpeople in the United Kingdom. Journal of the Royal Statistical Society, 74(4):421–422, March 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340404.

S:1911:BRBb

[190] T. H. C. S. Book review: Répertoire Technologique des Noms d’Industries et de Professions — Français, Anglais, Allemands — Avec Notices Descriptives Sommaires. Journal of the Royal Statistical Society, 74(4):423– 424, March 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340405.

C:1911:BRBc

[191] P. G. C. Book review: Land Problems and National Welfare, by Christopher Turnor. Journal of the Royal Statistical Society, 74(4):424–427, March 1911.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340406.

Anonymous:1911:ONPc

[192] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(4):427–429, March 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340407.

Anonymous:1911:CNc

[193] Anonymous. Current notes. Journal of the Royal Statistical Society, 74 (4):430–439, March 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340408.

Anonymous:1911:SEAc

[194] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74 (4):440–443, March 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340409.

Anonymous:1911:LALc

[195] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(4):444–447, March 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340410.

Anonymous:1911:PR

[196] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 74 (4):448–465, March 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340411.

Sale:1911:SSJ

[197] Charles V. Sale. Some statistics of Japan. Journal of the Royal Statistical Society, 74(5):467–539, April 1911.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2339878.

Dudfield:1911:SSM

[198] Reginald Dudfield. A suggested short method of constructing life-tables for infants under one year of age. Journal of the Royal Statistical Society, 74 (5):540–543, April 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339879.

B:1911:BRBa

[199] J. A. B. Book review: Royal Commission on the Poor Laws and Relief of Distress. Journal of the Royal Statistical Society, 74(5):544–546, April 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339880.

H:1911:BRBb

[200] R. H. H. Book review: Forest Resources of the World, by Raphael Zon. Journal of the Royal Statistical Society, 74 (5):546–547, April 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339881.

S:1911:BRBc

[201] S. S. Book review: Inleiding tot de Beoefening der Statistiek, by C. A. Verrijn Stuart. Journal of the Royal Statistical Society, 74(5):547–548, April 1911. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339882.

Y:1911:BRBb

[202] G. U. Y. Book review: Le Calcul Mécanique: Appareils Arithmétiques et Algébriques Intégrateurs, by L. Jacob. Journal of the Royal Statistical Society, 74(5):548–549, April 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339883.

D:1911:BRBb

[203] R. D. Book review: Child Problems, by Geo. B. Mangold. Journal of the Royal Statistical Society, 74(5):549–550, April 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339884.

P:1911:BRBa

[204] L. L. P. Book review: The Imperial Organization of Trade, by Geoffrey Drage. Journal of the Royal Statistical Society, 74(5):550–552, April 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339885.

L:1911:BRB

[205] W. R. L. Book review: Die Deutschen Anleihen, by Hugo Heyman. Journal of the Royal Statistical Society, 74 (5):552–553, April 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339886.

A:1911:BRBb

[206] P. A. Book review: La Politique Budgétaire en Europe: Les Tendances Actuelles. Journal of the Royal Statistical Society, 74(5):553–555, April 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339887.

A:1911:BRBc

[207] P. A. Book review: Direkte Staats- und Gemeindesteurn in der Schweiz, by Peter Hans Weiler. Journal of the Royal Statistical Society, 74(5):555, April 1911. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339888.

Anonymous:1911:ONPd

[208] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(5):556–558, April 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339889.

Anonymous:1911:CNd

[209] Anonymous. Current notes. Journal of the Royal Statistical Society, 74 (5):559–566, April 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339890.

Anonymous:1911:SEAd

[210] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74 (5):567–570, April 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339891.

Anonymous:1911:LALd

[211] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(5):571–574, April 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339892.

Snow:1911:AMM

[212] E. C. Snow. The application of the method of multiple correlation to the estimation of post-censal populations. Journal of the Royal Statistical Society, 74(6):575–629, May 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340076.

Johnson:1911:ASW

[213] Stewart Johnson. An attempt to show from what class the out-patients of a voluntary hospital are drawn. Journal of the Royal Statistical Society, 74 (6):630–640, May 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340077.

F:1911:BRBa

[214] A. W. F. Book review: An Introduction to the Theory of Statistics, by G. Udny Yule. Journal of the Royal Statistical Society, 74(6):641–643, May 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340078.

K:1911:BRBa

[215] J. M. K. Book review: Wahrscheinlichkeitsrechnung und ihre Anwendung auf Fehlerausgleichung, Statistik, und Lebensversicherung, by Emanuel Czuber; Wahrscheinlichkeitstheorie, Fehlerausgleichung, Kollektivmasslehre by B. G. Teubner; Mathematische Statistik, Mathematische Grundlagen der Lebensversicherung by B. G. Teubner. Journal of the Royal Statistical Society, 74(6):643–647, May 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340079.

B:1911:BRBb

[216] E. B. Book review: Unemployment Insurance: A Study of Schemes of Assisted Insurance, by I. G. Gibbon; L. T. Hobhouse. Journal of the Royal Statistical Society, 74(6):647, May 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340080.

B:1911:BRBc

[217] E. B. Book review: Workingmen’s Insurance in Europe, by Lee K. Frankel; Miles M. Dawson; Louis I. Dublin. Journal of the Royal Statistical Society, 74(6):648–649, May 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340081.

B:1911:BRBd

[218] E. B. Book review: Le Chômage: Causes, Conséquences, Remèdes, by A. de Lavergne; L. Paul Henry. Journal of the Royal Statistical Society, 74(6):649, May 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340082.

W:1911:BRBb

[219] A. D. W. Book review: Études Statistiques, Économiques, Sociales, Financières et Agricoles. Les Habitants. Démographie, Salaires, Corporations et Syndicats, Assistance, Mutualité, Épargne, Prévoyance, Impôts, by Edmond Michel. Journal of the Royal Statistical Society, 74(6):650–651, May 1911. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340083.

F:1911:BRBb

[220] H. S. F. Book review: Banques d’Émission et Trésors Publics, by Raphael-Georges Lévy. Journal of the Royal Statistical Society, 74(6): 652–653, May 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340084.

P:1911:BRBb

[221] L. L. P. Book review: Histoire du Commerce de la France. Première Partie: Avant 1789, by E. Levasseur. Journal of the Royal Statistical Society, 74 (6):653–656, May 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340085.

A:1911:BRBd

[222] P. A. Book review: Problems of Local Government, by G. Montagu Harris. Journal of the Royal Statistical Society, 74(6):656–657, May 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340086.

Anonymous:1911:ONPe

[223] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(6):657–660, May 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340087.

Anonymous:1911:CNe

[224] Anonymous. Current notes. Journal of the Royal Statistical Society, 74 (6):661–670, May 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340088.

Anonymous:1911:SEAe

[225] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74 (6):671–673, May 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340089.

Anonymous:1911:LALe

[226] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(6):674–678, May 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340090.

Vigor:1911:UNC

[227] H. D. Vigor. On the use of the “normal crop” as a standard in crop reports. Journal of the Royal Statistical Society, 74(7):679–692, June 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339741.

Popplewell:1911:SFE

[228] Frank Popplewell. Seasonal fluctuations in employment in the gas industry. Journal of the Royal Statistical Society, 74(7):693–734, June 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339742.

Baines:1911:PTC

[229] J. A. Baines. The provisional totals of the census of India. Journal of the Royal Statistical Society, 74 (7):735–737, June 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339743.

Snow:1911:EP

[230] E. C. Snow. Estimates of population. Journal of the Royal Statistical Society, 74(7):738–743, June 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339744.

H:1911:BRBc

[231] E. J. H. Book review: Report of an Inquiry, by the Board of Trade into Working Class Rents, Housing and Retail Prices, Together with the Rates of Wages in Certain Occupations in the Principal Industrial Towns of the United States of America. Journal of the Royal Statistical Society, 74 (7):744–749, June 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339745.

H:1911:BRBd

[232] R. H. H. Book review: The New Dictionary of Statistics. — A Complement to the 4th Edition of Mulhall’s “Dictionary of Statistics.”, by Augustus D. Webb. Journal of the Royal Statistical Society, 74(7):750–752, June 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339746.

C:1911:BRBd

[233] S. J. C. Book review: The Purchasing Power of Money, by Irving Fisher; H. G. Brown. Journal of the Royal Statistical Society, 74(7):752–754, June 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339747.

P:1911:BRBc

[234] L. L. P. Book review: State Socialism in New Zealand, by James Edward Le Rossignol; William Downie Stewart. Journal of the Royal Statistical Society, 74(7):754–756, June 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339748.

A:1911:BRBe

[235] P. A. Book review: Le Situation Économique et Financière de l’Égypte. Le Soudan Égyptien, by Pierre Arminjon. Journal of the Royal Statistical Society, 74(7):756–758, June 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339749.

A:1911:BRBf

[236] P. A. Book review: Die Entwicklung der Kommunalen Schulden in Deutschland, by Johannes Pfitzner. Journal of the Royal Statistical Society, 74(7):758, June 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339750.

P:1911:BRBd

[237] L. L. P. Book review: Le Mouvement Physiocratique en France de 1756 á 1770, by Georges Weulersse. Journal of the Royal Statistical Society, 74 (7):759–761, June 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339751.

Anonymous:1911:ONPf

[238] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(7):761–764, June 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339752.

Anonymous:1911:CNf

[239] Anonymous. Current notes. Journal of the Royal Statistical Society, 74 (7):765–770, June 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339753.

Anonymous:1911:SEAf

[240] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74 (7):771–776, June 1911. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339754.

Anonymous:1911:LALf

[241] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 74(7):777–780, June 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339755.

Anonymous:1911:PCR

[242] Anonymous. The preliminary census results. Journal of the Royal Statistical Society, 74(7):781–786, June 1911.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2339756.

Anonymous:1911:VIa

[243] Anonymous. Volume information. Journal of the Royal Statistical Society, 74(8):911–926, July 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340196.

Baines:1911:UC

[244] J. Athelstane Baines. Under the crown. Journal of the Royal Statistical Society, 74(8):787–796, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340197.

Anonymous:1911:PJ

[245] Anonymous. Proceedings on June 13. Journal of the Royal Statistical Society, 74(8):796–798, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340198.

Anonymous:1911:RPC

[246] Anonymous. Reception by the President, Council, and Officers. Journal of the Royal Statistical Society, 74 (8):799–805, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340199.

Anonymous:1911:SDC

[247] Anonymous. Statistical dinner club. Journal of the Royal Statistical Society, 74(8):806–817, July 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340200.

Anonymous:1911:RCF

[248] Anonymous. Report of the Council for the financial year ended December 31, 1910, and for the sessional year ending June 15, 1911, presented at the seventy-seventh Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C., on June 15, 1911. Journal of the Royal Statistical Society, 74(8):818–828, July 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340201.

Anonymous:1911:PSS

[249] Anonymous. Proceedings of the seventy-seventh Annual General Meeting, held on June 15, 1911. Journal of the Royal Statistical Society, 74 (8):829–831, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340202.

Anonymous:1911:MEL

[250] Anonymous. M. Emile Levasseur. Journal of the Royal Statistical Society, 74(8):832–834, July 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340203.

Baines:1911:CN

[251] J. A. Baines. Census notes. Journal of the Royal Statistical Society, 74 (8):835–840, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340204.

K:1911:BRBb

[252] J. M. K. Book review: Publications Issued, by and in Preparation for the National Monetary Commission of the United States. Journal of the Royal Statistical Society, 74(8):841–846, July 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340205.

D:1911:BRBc

[253] R. D. Book review: England’s Recent Progress: An Investigation of the Statistics of Migrations, Mortality, in the Twenty Years from 1881 to 1901, as Indicating Tendencies Towards the Growth or Decay of Particular Communities, by Thomas A. Welton. Journal of the Royal Statistical Society, 74 (8):846–848, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340206.

C:1911:BRBe

[254] P. G. C. Book review: Large and Small Holdings: A Study of English Agricultural Economics, by Hermann Levy. Journal of the Royal Statistical Society, 74(8):848–851, July 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340207.

C:1911:BRBf

[255] P. G. C. Book review: Canada and the Empire, by W. R. Lawson. Journal of the Royal Statistical Society, 74 (8):851–853, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340208.

B:1911:BRBe

[256] E. B. Book review: Insurance Science and Economics: A Practical Discussion of Present-Day Problems of Administration, Methods and Results for Insurance Officials, Managers, Agents, Merchants, Lawyers, Teachers, Students and Other Interested in the Broader Aspects of Insurance as a Business in its Relation to Public Welfare and the State, by Frederick L. Hoffman. Journal of the Royal Statistical Society, 74(8):853–855, July 1911.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340209.

C:1911:BRBg

[257] C. E. C. Book review: The Living Wage of Women Workers: A Study of Incomes and Expenditures of 450 Women Workers in the City of Boston, by Louise Marion Bosworth; F. Spencer Baldwin. Journal of the Royal Statistical Society, 74(8):855–856, July 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340210.

P:1911:BRBe

[258] L. L. P. Book review: Le Commerce Exterieur et les Tarifs de Douane, by Aug. Arnaune. Journal of the Royal Statistical Society, 74(8):857–858, July 1911. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340211.

P:1911:BRBf

[259] L. L. P. Book review: Cours Elementaire de Science des Finances et de Legislation Financiere Francaise, by Gaston Jeze. Journal of the Royal Statistical Society, 74(8):858–860, July 1911. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340212.

Anonymous:1911:ONPg

[260] Anonymous. Other new publications. Journal of the Royal Statistical Society, 74(8):860–862, July 1911. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340213.

Morrison:1911:C

[261] G. B. Morrison. Correspondence. Journal of the Royal Statistical Society, 74 (8):863–868, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340214.

Anonymous:1911:CNg

[262] Anonymous. Current notes. Journal of the Royal Statistical Society, 74 (8):869–872, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340215.

Anonymous:1911:SEAg

[263] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 74 (8):873–876, July 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340216.

Anonymous:1911:MLA

[264] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 74(8):877–879, July 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340217.

Anonymous:1911:ALA

[265] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 74(8):880–909, July 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340218.

Anonymous:1911:VIb

[266] Anonymous. Volume information. Journal of the Royal Statistical Society, 75(1):i–viii, December 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340335.

Hooker:1911:CPH

[267] R. H. Hooker. The course of prices at home and abroad, 1890–1910. Journal of the Royal Statistical Society, 75 (1):1–50, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340336.

R:1911:ISI

[268] R. H. R. International Statistical Institute. Journal of the Royal Statistical Society, 75(1):51–63, December 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340337.

Clarke:1911:OET

[269] Ernest Clarke. An old exchequer tally. Journal of the Royal Statistical Society, 75(1):63–77, December 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340338.

Bowley:1911:MAA

[270] A. L. Bowley. The measurement of the accuracy of an average. Journal of the Royal Statistical Society, 75(1): 77–88, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340339.

B:1911:BRBf

[271] J. A. B. Book review: The Economic Transition in India, by Theodore Morison. Journal of the Royal Statistical Society, 75(1):89–91, December 1911. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340340.

A:1911:BRBg

[272] W. M. A. Book review: Railway Operating Statistics, by C. P. Mossop. Journal of the Royal Statistical Society, 75 (1):91–94, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340341.

A:1911:BRBh

[273] W. M. A. Book review: Railways and Nationalisation, by Edwin A. Pratt. Journal of the Royal Statistical Society, 75(1):94–96, December 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340342.

V:1911:BRB

[274] H. D. V. Book review: Fisheries of the United States, 1908. Journal of the Royal Statistical Society, 75(1): 96–98, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340343.

A:1911:BRBi

[275] P. A. Book review: Die Statistik in Deutschland Nach Ihrem Heutigen Stand, by F. Zahn. Journal of the Royal Statistical Society, 75(1): 98–100, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340344.

W:1911:BRBc

[276] A. D. W. Book review: Deutschlands Wirtschaftliche Entwicklung unter Besonderer Berücksichtigung der Volkszählung 1905 Sowie der Berufsund Betriebszählung 1907, by Friedrich Zahn. Journal of the Royal Statistical Society, 75(1):100–102, December 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340345.

A:1911:BRBj

[277] P. A. Book review: Über die Ernährung und Deren Kosten bei Deutschen Arbeitern, by H. Lichtenfelt; Massenverbrauch und Preisbewegung in der Schweiz by F. Krömmelbein; Basler Volkswirtschaftliche Arbeiten, Herausgegeben und Eingeleitet von Stephan Bauer by Stephan Bauer; W. Kohlhammer. Journal of the Royal Statistical Society, 75(1):102–104, December 1911. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340346.

F:1911:BRBc

[278] A. W. F. Book review: Sveriges Officiella Statistik och Dess Almänna Organisation. Journal of the Royal Statistical Society, 75(1):104–106, December 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340347.

M:1911:BRB

[279] B. M. Book review: The Income Tax, by Edwin R. A. Seligman. Journal of the Royal Statistical Society, 75(1): 106–108, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340348.

P:1911:BRBg

[280] L. L. P. Book review: La Politique Douanière de la France, by Charles Augier; Angel Marvaud. Journal of the Royal Statistical Society, 75(1): 108–110, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340349.

B:1911:BRBg

[281] E. B. Book review: Papers on Inter-Racial Problems, by G. Spiller. Journal of the Royal Statistical Society, 75(1): 110–111, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340350.

Anonymous:1911:ONPh

[282] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(1):111–117, December 1911. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340351.

Anonymous:1911:C

[283] Anonymous. Correspondence. Journal of the Royal Statistical Society, 75(1): 118–119, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340352.

Anonymous:1911:CNh

[284] Anonymous. Current notes. Journal of the Royal Statistical Society, 75(1): 120–137, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340353.

Anonymous:1911:SEAh

[285] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75(1): 138–144, December 1911. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340354.

Anonymous:1911:LALg

[286] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(1):145–156, December 1911. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340355.

Crammond:1912:EPS

[287] Edgar Crammond. The economic position of Scotland and her financial relations with England and Ireland. Journal of the Royal Statistical Society, 75 (2):157–182, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340319.

Wilson:1912:SRV

[288] George B. Wilson. A statistical review of the variations during the last twenty years in the consumption of intoxicating drinks in the United Kingdom, and in convictions for offences connected with intoxication, with a discussion of the causes to which these variations may be ascribed. Journal of the Royal Statistical Society, 75(2): 183–247, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340320.

B:1912:BRBa

[289] E. B. Book review: Unemployment: A Social Study, by B. Seebohm Rowntree; Bruno Lasker. Journal of the Royal Statistical Society, 75 (2):248–250, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340321.

B:1912:BRBb

[290] E. B. Book review: Le Chomage et son Remede, by Daniel Bellet; Paul Leroy-Beaulieu. Journal of the Royal Statistical Society, 75(2):250–251, January 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340322.

S:1912:BRBa

[291] D. F. S. Book review: Zur Frage der Arbeitslosenversicherung, der Arbeitsvermittelung und der Arbeitsbeschaffung, by von August Baab. Journal of the Royal Statistical Society, 75(2):251–253, January 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340323.

B:1912:BRBc

[292] E. B. Book review: Traite des Retraites Ouvrieres. Commentaire Theorique et Pratique de la Loi du 5 Avril, 1910, by Pierre Pinot; J. Comolet-Tirman; M. Alfred Picard. Journal of the Royal Statistical Society, 75(2): 253–254, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340324.

W:1912:BRBa

[293] J. W. Book review: Le Credit Agricole en France, by Henry Sagnier. Journal of the Royal Statistical Society, 75 (2):254–257, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340325.

C:1912:BRBa

[294] P. G. C. Book review: Wheat Growing in Canada, the United States and the Argentine, by W. P. Rutter. Journal of the Royal Statistical Society, 75 (2):257–259, January 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340326.

W:1912:BRBb

[295] J. W. Book review: A Study of Indian Economics, by Pramathanath Banerjea. Journal of the Royal Statistical Society, 75(2):259–261, January 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340327.

B:1912:BRBd

[296] J. A. B. Book review: British Dominions: Their Present Commercial and Industrial Condition, by W. J. Ashley. Journal of the Royal Statistical Society, 75(2):261–263, January 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340328.

P:1912:BRBa

[297] L. L. P. Book review: Ultima Verba: Mon Dernier Ouvrage, by G. de Molinari. Journal of the Royal Statistical Society, 75(2):263–265, January 1912. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340329.

P:1912:BRBb

[298] L. L. P. Book review: Outlines of Political Economy, by S. J. Chapman. Journal of the Royal Statistical Society, 75 (2):265–267, January 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340330.

Anonymous:1912:ONPa

[299] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(2):267–269, January 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340331.

Anonymous:1912:CNa

[300] Anonymous. Current notes. Journal of the Royal Statistical Society, 75(2): 270–277, January 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340332.

Anonymous:1912:SEAa

[301] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75(2): 278–282, January 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340333.

Anonymous:1912:LALa

[302] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(2):283–292, January 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340334.

Chapman:1912:REC

[303] S. J. Chapman and F. J. Marquis. The recruiting of the employing classes from the ranks of the wage-earners in the cotton industry. Journal of the Royal Statistical Society, 75(3): 293–313, February 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340252.

Fremantle:1912:SSA

[304] H. E. S. Fremantle. The statistics of South Africa. Journal of the Royal Statistical Society, 75(3):314–322, February 1912. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340253.

B:1912:BRBe

[305] E. B. Book review: Investigation Relative to Wages and Prices of Commodities. Vol. I. Report of Committee and Views of Minority. Vol. II. Hearings Before Committee and Digest of Evidence. Vol. III. Wages and Prices in United States and Abroad. Vol. IV. Wholesale and Retail Prices in United States and Abroad ; Report of the Commission on the Cost of Living ; Wages in the United States, 1908–10, by Scott Nearing. Journal of the Royal Statistical Society, 75(3):323–326, February 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340254.

C:1912:BRBb

[306] E. C. Book review: L’Appauvrissement des Masses: Essai de Critique Sociale, by Berthold Missiaen; O. M. Cap. Journal of the Royal Statistical Society, 75(3):326–328, February 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340255.

W:1912:BRBc

[307] A. D. W. Book review: Die Einkommen-Entwicklung in Preussen seit 1896 Nebst Kritik an Material und Methoden, by Klara Peris. Journal of the Royal Statistical Society, 75(3): 328–329, February 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340256.

A:1912:BRBa

[308] P. A. Book review: L’Europe Économique, by Edmond Théry. Journal of the Royal Statistical Society, 75(3):329–331, February 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340257.

A:1912:BRBb

[309] P. A. Book review: L’Évolution Industrielle de la Belgique, by Jan St. Lewinski. Journal of the Royal Statistical Society, 75(3):331–332, February 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340258.

F:1912:BRBa

[310] H. S. F. Book review: Dictionary of Banking. A Concise Encyclopœdia of Banking Law and Practice, by William Thomson. Journal of the Royal Statistical Society, 75(3):332–333, February 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340259.

F:1912:BRBb

[311] A. W. F. Book review: Hostudsigter og Hostudbytte i Danmark, by Einar Cohn. Journal of the Royal Statistical Society, 75(3):333–334, February 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340260.

Anonymous:1912:ONPb

[312] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(3):335–338, February 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340261.

Anonymous:1912:CNb

[313] Anonymous. Current notes. Journal of the Royal Statistical Society, 75(3): 339–349, February 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340262.

Anonymous:1912:SEAb

[314] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75(3): 350–353, February 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340263.

Anonymous:1912:LALb

[315] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(3):354–357, February 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340264.

Anonymous:1912:PRa

[316] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 75(3):358–360, February 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340265.

Macdonald:1912:RIS

[317] Robert A. Macdonald. The rate of interest since 1844. Journal of the Royal Statistical Society, 75(4): 361–379, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340279.

Williams:1912:RDP

[318] T. T. Williams. The rate of discount and the price of consols. Journal of the Royal Statistical Society, 75 (4):380–411, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340280.

Sauerbeck:1912:PC

[319] A. Sauerbeck. Prices of commodities in 1911. Journal of the Royal Statistical Society, 75(4):412–425, March 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340281.

Anonymous:1912:CIC

[320] Anonymous. The census of India: — Christians by race and sect. Journal of the Royal Statistical Society, 75 (4):426–428, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340282.

M:1912:BRBa

[321] H. W. M. Book review: Statistique Générale de la France: Salaires et Coût de L’Existence à Diverses Époques, Jusqu’en 1910. Ministère du Travail et de la Prévoyange Sociale. Journal of the Royal Statistical Society, 75 (4):429–431, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340283.

W:1912:BRBd

[322] A. D. W. Book review: Haushaltungsstatiski. Eine Literarhistorische und Methodologische Untersuchung, by Gerhard Albrecht. Journal of the Royal Statistical Society, 75(4): 431–432, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340284.

W:1912:BRBe

[323] A. D. W. Book review: Weltwirtschaftliche Studien. Vorträge und Aufsätze, by Hermann Schumacher. Journal of the Royal Statistical Society, 75(4): 432–433, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340285.

P:1912:BRBc

[324] L. L. P. Book review: The Case Against Free Trade, by W. Cunningham; Joseph Chamberlain; The Case Against Tariff Reform, by E. Enever Todd. Journal of the Royal Statistical Society, 75(4):433–436, March 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340286.

A:1912:BRBc

[325] P. A. Book review: The Tariff in Our Times, by Ida M. Tarbell. Journal of the Royal Statistical Society, 75 (4):436–437, March 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340287.

G:1912:BRBa

[326] M. G. Book review: An Introduction to Eugenics, by W. C. D. Whetham; C. D. Whetham. Journal of the Royal Statistical Society, 75(4):437–438, March 1912. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340288.

R:1912:BRBa

[327] R. H. R. Book review: The Village Labourer, 1760–1832. A Study in the Government of England before the Reform Bill, by J. L. Hammond; Barbara Hammond. Journal of the Royal Statistical Society, 75(4):438–440, March 1912. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340289.

Anonymous:1912:ONPc

[328] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(4):440–443, March 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340290.

Anonymous:1912:CNc

[329] Anonymous. Current notes. Journal of the Royal Statistical Society, 75 (4):444–449, March 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340291.

Anonymous:1912:SEAc

[330] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75 (4):450–454, March 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340292.

Anonymous:1912:LALc

[331] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(4):455–458, March 1912. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340293.

Anonymous:1912:PRb

[332] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 75 (4):459–466, March 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340294.

Ashley:1912:FSG

[333] Percy Ashley. The financial systems of Germany. Journal of the Royal Statistical Society, 75(5):467–504, April 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340111.

March:1912:SRC

[334] Lucien March. Some researches concerning the factors of mortality. Journal of the Royal Statistical Society, 75 (5):505–538, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340112.

Johnson:1912:RBL

[335] Stewart Johnson. The relation between large families, poverty, irregularity of earnings and crowding. Journal of the Royal Statistical Society, 75 (5):539–550, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340113.

Y:1912:BRB

[336] G. U. Y. Book review: Laws of Wages: An Essay in Statistical Economics, by Henry Ludwell Moore. Journal of the Royal Statistical Society, 75(5): 551–553, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340114.

A:1912:BRBd

[337] P. A. Book review: Die Bewegung der Lohne der Löndlichen “Freien” Arbeiter vom Ausgang des 18 Jahrhunderts bis 1850, by Anna Neumann. Journal of the Royal Statistical Society, 75(5):553–554, April 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340115.

M:1912:BRBb

[338] B. M. Book review: La Fortune Publique de la France, by Edmund Théry. Journal of the Royal Statistical Society, 75(5):554–555, April 1912. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340116.

B:1912:BRBf

[339] E. B. Book review: National Insurance, by A. S. Comyns Carr; W. H. Stuart Garnett; J. H. Taylor; D. Lloyd George. Journal of the Royal Statistical Society, 75(5):555–556, April 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340117.

B:1912:BRBg

[340] E. B. Book review: The Poor Law Enigma, by M. Fothergill Robinson. Journal of the Royal Statistical Society, 75(5):556–557, April 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340118.

D:1912:BRB

[341] Reg. D. Book review: La Dépopulation de la France, by Jacques Bertillon. Journal of the Royal Statistical Society, 75(5):558–560, April 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340119.

G:1912:BRBb

[342] M. G. Book review: Krebs und Tuberkulose in Beruflicher Beziehung vom Standpunkte der Vergleichenden Internationalen Statistik, by Robert Behla. Journal of the Royal Statistical Society, 75(5):560–561, April 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340120.

A:1912:BRBe

[343] P. A. Book review: Les Conventions Monétaires, by A. Janssen. Journal of the Royal Statistical Society, 75 (5):561–562, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340121.

Anonymous:1912:ONPd

[344] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(5):562–564, April 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340122.

Anonymous:1912:CNd

[345] Anonymous. Current notes. Journal of the Royal Statistical Society, 75 (5):565–569, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340123.

Anonymous:1912:SEAd

[346] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75 (5):570–573, April 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340124.

Anonymous:1912:LALd

[347] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(5):574–577, April 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340125.

Yule:1912:MMA

[348] G. Udny Yule. On the methods of measuring association between two attributes. Journal of the Royal Statistical Society, 75(6):579–652, May 1912.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340126.

Anonymous:1912:RDP

[349] Anonymous. The rate of discount and the price of consols. Journal of the Royal Statistical Society, 75 (6):653–658, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340127.

Anonymous:1912:DWO

[350] Anonymous. Dr. William Ogle. Journal of the Royal Statistical Society, 75 (6):659–661, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340128.

P:1912:BRBd

[351] L. L. P. Book review: Résultats de la Deuxième Révision Décennale du Revenu Net des Propriétés Bâties, by M. L. L. Klotz; M. Maurice-Bloch. Journal of the Royal Statistical Society, 75(6):662–663, May 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340129.

P:1912:BRBe

[352] L. L. P. Book review: Le Protectionnisme Ouvrier (L’Exclusion des Travailleurs Etrangers), by Giuseppe Prato; Georges Bourgin. Journal of the Royal Statistical Society, 75 (6):663–664, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340130.

L:1912:BRBa

[353] W. R. L. Book review: Monetary Economics, by W. W. Carlile. Journal of the Royal Statistical Society, 75 (6):665–666, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340131.

Anonymous:1912:ONPe

[354] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(6):666–669, May 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340132.

Anonymous:1912:CNe

[355] Anonymous. Current notes. Journal of the Royal Statistical Society, 75 (6):670–677, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340133.

Anonymous:1912:SEAe

[356] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75 (6):678–682, May 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340134.

Anonymous:1912:LALe

[357] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(6):683–686, May 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340135.

Acworth:1912:BRT

[358] W. M. Acworth and George Paish. British railways: Their accounts and statistics. Journal of the Royal Statistical Society, 75(7):687–743, June 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339820.

Greenwood:1912:SSJ

[359] Arthur Greenwood and John E. Kettlewell. Some statistics of juvenile employment and unemployment. Journal of the Royal Statistical Society, 75 (7):744–753, June 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339821.

H:1912:BRB

[360] R. H. H. Book review: An Introduction to the Study of Prices, with Special Reference to the History of the Nineteenth Century, by Walter T. Layton. Journal of the Royal Statistical Society, 75(7):754–755, June 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339822.

M:1912:BRBc

[361] H. W. M. Book review: Die Entwicklung der Munchener Fleischpreise seit Beginn des 19 Jahrhunderts und ihre Ursachen, by Carl Gschwendtner; Wholesale Prices, Canada, 1911, by Department of Labor; Competitive and Monopoly Price, by F. C. Hicks. Journal of the Royal Statistical Society, 75 (7):756–757, June 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339823.

G:1912:BRBc

[362] M. G. Book review: Nouveaux Fondements de la Theorie de la Statistique, by Casimir Maciejewski. Journal of the Royal Statistical Society, 75 (7):757–759, June 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339824.

A:1912:BRBf

[363] P. A. Book review: England’s Industrial Development: A Historical Survey of Commerce and Industry, by Arthur D. Innes. Journal of the Royal Statistical Society, 75(7):759–760, June 1912.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2339825.

Bauer:1912:BRB

[364] Stephan Bauer and P. A. Book review: Kapital- und Verwaltungs-Beteiligung der Arbeiter in den Britischen Produktivgenossenschaften, by Johannes Huber. Journal of the Royal Statistical Society, 75(7):760–762, June 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339826.

F:1912:BRBc

[365] A. W. F. Book review: Principles of Economics, by F. W. Taussig. Journal of the Royal Statistical Society, 75 (7):762–764, June 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339827.

P:1912:BRBf

[366] L. L. P. Book review: Le Bilan du Protectionnisme en France, by G. Schelle. Journal of the Royal Statistical Society, 75(7):764–765, June 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339828.

S:1912:BRBb

[367] J. C. S. Book review: La Lutte Contre le Sweating-System, by Paul Boyaval. Journal of the Royal Statistical Society, 75(7):765–767, June 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339829.

R:1912:BRBb

[368] R. H. R. Book review: Common Land and Inclosure, by E. C. K. Gonner. Journal of the Royal Statistical Society, 75(7):767–769, June 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339830.

Anonymous:1912:ONPf

[369] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(7):769–771, June 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339831.

Anonymous:1912:CNf

[370] Anonymous. Current notes. Journal of the Royal Statistical Society, 75 (7):772–780, June 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339832.

Anonymous:1912:SEAf

[371] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 75 (7):781–786, June 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339833.

Anonymous:1912:LALf

[372] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 75(7):787–790, June 1912. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339834.

Anonymous:1912:VIa

[373] Anonymous. Volume information. Journal of the Royal Statistical Society, 75(8):893–903, July 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340379.

Bowley:1912:MEE

[374] A. L. Bowley. The measurement of employment: An experiment. Journal of the Royal Statistical Society, 75 (8):791–829, July 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340380.

Anonymous:1912:RCF

[375] Anonymous. Report of the Council for the financial year ended December 31, 1911, and for the sessional year ending June 18, 1912, presented at the seventy-eighth Annual General Meeting of the Royal Statistical Society, held at the rooms of the Royal Society of Arts, John Street, Adelphi, London, W.C., on June 18, 1912. Journal of the Royal Statistical Society, 75 (8):830–839, July 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340381.

Anonymous:1912:PSE

[376] Anonymous. Proceedings of the seventy-eighth Annual General Meeting, held on June 18, 1912. Journal of the Royal Statistical Society, 75 (8):840–841, July 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340382.

Baines:1912:CNB

[377] J. A. Baines. Census notes.–the British empire. Journal of the Royal Statistical Society, 75(8):842–846, July 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340383.

S:1912:BRBc

[378] W. S. Book review: The Agricultural Output of Great Britain. Report on Enquiries Made, by the Board of Agriculture and Fisheries in Connection with the Census of Production Act, 1906. Journal of the Royal Statistical Society, 75(8):847–848, July 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340384.

B:1912:BRBh

[379] E. B. Book review: Social Insurance in Germany, 1883–1911; Its History, Operation, Results: And a Comparison with The National Insurance Act, 1911, by William Harbutt Dawson. Journal of the Royal Statistical Society, 75 (8):848–850, July 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340385.

A:1912:BRBg

[380] P. A. Book review: Statistik der Zivilrechtspflege, by Max Rusch. Journal of the Royal Statistical Society, 75 (8):850–851, July 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340386.

A:1912:BRBh

[381] P. A. Book review: Die Statistik im Fabrik- und Warenshandelsbetrieb, by Albert Calmes. Journal of the Royal Statistical Society, 75(8):852, July 1912. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340387.

G:1912:BRBd

[382] M. G. Book review: Preventable Cancer: A Statistical Research, by Rollo Russell. Journal of the Royal Statistical Society, 75(8):852–854, July 1912. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340388.

P:1912:BRBg

[383] L. L. P. Book review: Quetelet: Statisticien et Sociologue, by Joseph Lottin. Journal of the Royal Statistical Society, 75(8):854–856, July 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340389.

G:1912:BRBe

[384] E. C. K. G. Book review: The Distribution of Income, by William Smart. Journal of the Royal Statistical Society, 75(8):857–858, July 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340390.

G:1912:BRBf

[385] E. C. K. G. Book review: Value and Distribution: A Critical and Constructive Study, by H. J. Davenport. Journal of the Royal Statistical Society, 75 (8):858–859, July 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340391.

P:1912:BRBh

[386] L. L. P. Book review: The Laws of Supply and Demand with Special Reference to Their Influence On Over-Production and Unemployment, by George Binney Dibblee. Journal of the Royal Statistical Society, 75(8):859–861, July 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340392.

P:1912:BRBi

[387] L. L. P. Book review: Précis d’Économie Politique, by Charles Brouihet. Journal of the Royal Statistical Society, 75(8):861–863, July 1912. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340393.

L:1912:BRBb

[388] W. T. L. Book review: Le Stock Exchange. Les Usages de la Place de Londres et les Valeurs de Placement, by G. Chevilliard. Journal of the Royal Statistical Society, 75(8):863–865, July 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340394.

B:1912:BRBi

[389] E. B. Book review: Principles and Methods of Municipal Trading, by Douglas Knoop. Journal of the Royal Statistical Society, 75(8):865–866, July 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340395.

Anonymous:1912:ONPg

[390] Anonymous. Other new publications. Journal of the Royal Statistical Society, 75(8):866–869, July 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340396.

Anonymous:1912:CNg

[391] Anonymous. Current notes. Journal of the Royal Statistical Society, 75 (8):870–875, July 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340397.

Anonymous:1912:SEAg

[392] Anonymous. Statistical and economic articles in recent periodical. Journal of the Royal Statistical Society, 75 (8):876–879, July 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340398.

Anonymous:1912:MLA

[393] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 75(8):880–883, July 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340399.

Anonymous:1912:ALA

[394] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 75(8):884–892, July 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340400.

Anonymous:1912:VIb

[395] Anonymous. Volume information. Journal of the Royal Statistical Society, 76(1):i–viii, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340026.

Dudfield:1912:SBR

[396] Reginald Dudfield. Still-births in relation to infantile mortality. Journal of the Royal Statistical Society, 76 (1):1–26, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340027.

Anonymous:1912:RSC

[397] Anonymous. Report of special committee on infantile mortality. Journal of the Royal Statistical Society, 76(1): 27–87, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340028.

Cunynghame:1912:AES

[398] Henry H. Cunynghame. Address to the economic science and statistical section of the British Association for the Advancement of Science, 1912. Journal of the Royal Statistical Society, 76(1): 88–98, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340029.

Rew:1912:NFS

[399] R. H. Rew. The Nation’s food supply. Journal of the Royal Statistical Society, 76(1):98–105, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340030.

Craigie:1912:SAC

[400] P. G. Craigie. Scottish agricultural changes. Journal of the Royal Statistical Society, 76(1):105–107, December 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340031.

B:1912:BRBj

[401] J. A. B. Book review: The Progress of the Nation, by G. R. Porter; F. W. Hirst. Journal of the Royal Statistical Society, 76(1):108–111, December 1912. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340032.

B:1912:BRBk

[402] A. L. B. Book review: The Elements of Statistical Method, by Willford I. King. Journal of the Royal Statistical Society, 76(1):111–112, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340033.

K:1912:BRBa

[403] J. M. K. Book review: Calcul des Probabilités, by H. Poincaré; Calcul des Probabilités by Louis Bachelier; Le Calcul des Probabilités et ses Applications by E. Carvallo; Wahrscheinlichkeitsrechnung by A. A. Markoff; H. Liebmann. Journal of the Royal Statistical Society, 76(1):113–116, December 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340034.

K:1912:BRBb

[404] J. M. K. Book review: Ueber das Geschlechtsverhältnis bei Zwillingsgeburten, by Kazimierz J. Horowicz.

Journal of the Royal Statistical Society, 76(1):116–117, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340035.

A:1912:BRBi

[405] W. M. A. Book review: Manual of Railway Statistics, by George L. Boag. Journal of the Royal Statistical Society, 76(1):117–119, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340036.

B:1912:BRBl

[406] J. A. B. Book review: The Industrial Panjab: A Survey of Facts, Conditions, and Possibilities, by A. Latifi. Journal of the Royal Statistical Society, 76(1): 119–120, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340037.

C:1912:BRBc

[407] S. J. C. Book review: Elementary Principles of Economics, by Irving Fisher. Journal of the Royal Statistical Society, 76(1):120–122, December 1912. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340038.

B:1912:BRBm

[408] E. B. Book review: An Introduction to the History of Life Assurance, by A. Fingland Jack. Journal of the Royal Statistical Society, 76(1): 122–123, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340039.

A:1912:BRBj

[409] P. A. Book review: Die Deutschen Techniker: Ihre Lebens-, Ausbildungsund Arbeitsverhältnisse, by Adolf Günther. Journal of the Royal Statistical Society, 76(1):123–125, December 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340040.

A:1912:BRBk

[410] P. A. Book review: Deutschlands Kulturausgaben, by Ludwig Sevin. Journal of the Royal Statistical Society, 76(1): 125–126, December 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340041.

G:1912:BRBg

[411] M. G. Book review: Problems in Eugenics. Papers Communicated to the First International Eugenics Congress Held at the University of London, July 24 to 30, 1912, by Eugenics Education Society. Journal of the Royal Statistical Society, 76(1):126–128, December 1912. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340042.

Anonymous:1912:ONPh

[412] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(1):128–132, December 1912. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340043.

Anonymous:1912:CNh

[413] Anonymous. Current notes. Journal of the Royal Statistical Society, 76(1): 133–141, December 1912. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340044.

Anonymous:1912:SEAh

[414] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76(1):

142–149, December 1912. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340045.

Anonymous:1912:LALg

[415] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 76(1):150–164, December 1912. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340046.

Edgeworth:1913:UTP

[416] F. Y. Edgeworth. On the use of the theory of probabilities in statistics relating to society. Journal of the Royal Statistical Society, 76(2):165–193, January 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340091.

Anonymous:1913:PD

[417] Anonymous. Proceedings on December 17, 1912. Journal of the Royal Statistical Society, 76(2):194–195, January 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340092.

Lehfeldt:1913:RIB

[418] R. A. Lehfeldt. The rate of interest on British and foreign investments. Journal of the Royal Statistical Society, 76 (2):196–207, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340093.

Webb:1913:CAL

[419] Augustus D. Webb. The consumption of alcoholic liquors in the United Kingdom. Journal of the Royal Statistical Society, 76(2):207–220, January 1913. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340094.

B:1913:BRBa

[420] E. B. Book review: Medical Benefit: A Study of the Experience of Germany and Denmark, by I. G. Gibbon. Journal of the Royal Statistical Society, 76 (2):221–222, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340095.

Y:1913:BRBa

[421] G. U. Y. Book review: Statistique des Familles en 1906. Journal of the Royal Statistical Society, 76(2): 223–224, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340096.

W:1913:BRBa

[422] A. D. W. Book review: Illustrierte Deutsche Statistik. Diagramme und Stufenkarten. (System Kowatsch.), by Ambr. Kowatsch. Journal of the Royal Statistical Society, 76(2):224, January 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340097.

W:1913:BRBb

[423] A. D. W. Book review: Die Fleischversorgung des Deutschen Reiches. Eine Untersuchung der Ursachen und Wirkungen der Fleischteuerung und der Mittel zur Abhilfe, by Joseph Bergfried Esslen. Journal of the Royal Statistical Society, 76(2):224– 226, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340098.

Y:1913:BRBb

[424] G. U. Y. Book review: Emile Cheysson: Œuvres Choisies. Journal of the Royal Statistical Society, 76(2):

226–227, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340099.

F:1913:BRBa

[425] A. W. F. Book review: Wealth and Welfare, by A. C. Pigou. Journal of the Royal Statistical Society, 76(2): 227–230, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340100.

P:1913:BRBa

[426] L. L. P. Book review: History of Economic Thought, by Lewis H. Haney. Journal of the Royal Statistical Society, 76(2):230–231, January 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340101.

A:1913:BRBa

[427] P. A. Book review: Histoire du Commerce de la France, by E. Levasseur. Journal of the Royal Statistical Society, 76(2):231–233, January 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340102.

C:1913:BRBa

[428] P. G. C. Book review: Agricultural Organisation: Its Rise, Principles and Practice Abroad and at Home, by Edwin A. Pratt. Journal of the Royal Statistical Society, 76(2):233–234, January 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340103.

W:1913:BRBc

[429] J. W. Book review: Co-Operation in Agriculture, by Henry W. Wolff. Journal of the Royal Statistical Society, 76 (2):234–235, January 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340104.

C:1913:BRBb

[430] E. C. Book review: The Exchequer in the Twelfth Century, by Reginald L. Poole. Journal of the Royal Statistical Society, 76(2):235–236, January 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340105.

Anonymous:1913:ONPa

[431] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(2):237–240, January 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340106.

Anonymous:1913:CNa

[432] Anonymous. Current notes. Journal of the Royal Statistical Society, 76(2): 241–247, January 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340107.

Anonymous:1913:SEAa

[433] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76(2): 248–249, January 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340108.

Anonymous:1913:MLAa

[434] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(2):250–252, January 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340109.

Anonymous:1913:PRa

[435] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 76 (2):253–259, January 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340110.

Gonner:1913:PEE

[436] Edward C. K. Gonner. The population of England in the Eighteenth Century. Journal of the Royal Statistical Society, 76(3):261–303, February 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340047.

Welton:1913:NUR

[437] Thomas A. Welton. Note on urban and rural variations according to the English census of 1911. Journal of the Royal Statistical Society, 76(3): 304–317, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340048.

Y:1913:BRBc

[438] G. U. Y. Book review: Final Report on the First Census of Production of the United Kingdom (1907). Journal of the Royal Statistical Society, 76(3): 318–321, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340049.

C:1913:BRBc

[439] E. C. Book review: Lamb’s Friend the Census-Taker: Life and Letters of John Rickman, by Orlo Williams. Journal of the Royal Statistical Society, 76(3): 321–325, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340050.

B:1913:BRBb

[440] A. L. B. Book review: Soziologie und Statistik, by Franz Žižek. Journal of the Royal Statistical Society, 76(3): 325–326, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340051.

S:1913:BRBa

[441] E. C. S. Book review: Variabilità e Mutabilità, by Corrado Gini. Journal of the Royal Statistical Society, 76(3): 326–327, February 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340052.

A:1913:BRBb

[442] P. A. Book review: La Gestion par l’État et les Municipalités, by Yves Guyot. Journal of the Royal Statistical Society, 76(3):327–330, February 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340053.

A:1913:BRBc

[443] P. A. Book review: Lohnsystem und Löhne in der Brauindustrie: Tarifvertrage, by Emil Wolff. Journal of the Royal Statistical Society, 76(3): 330–331, February 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340054.

A:1913:BRBd

[444] P. A. Book review: Die Finanzreformversuche im Deutschen Reiche von 1867 bis zur Gegenwart, by Egbert Begemann. Journal of the Royal Statistical Society, 76(3):332–333, February 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340055.

G:1913:BRBa

[445] M. G. Book review: Der Geburtenrückgang– Die Rationalisierung des Sexuallebens in Unserer Zeit, by Julius Wolf; Der Geburtenrückgang in Deutschland, seine Bewertung und Bekämpfung by J. Borntraeger; Die Beschränkung der Geburtenzahl ein Kulturproblem by Julian Marcuse. Journal of the Royal Statistical Society, 76(3):333– 335, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340056.

W:1913:BRBd

[446] A. D. W. Book review: Die Deutsche Überseeische Auswanderung. Ein Beitrag zur Deutschen Wanderungsgeschichte, by Wilhelm Mönckmeier. Journal of the Royal Statistical Society, 76(3):335, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340057.

C:1913:BRBd

[447] S. J. C. Book review: The Economic Outlook, by Edwin Cannan. Journal of the Royal Statistical Society, 76(3): 335–336, February 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340058.

S:1913:BRBb

[448] C. P. S. Book review: Principles of Economics, by N. G. Pierson; A. A. Wotzel. Journal of the Royal Statistical Society, 76(3):337–338, February 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340059.

Anonymous:1913:ONPb

[449] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(3):338–339, February 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340060.

Anonymous:1913:CNb

[450] Anonymous. Current notes. Journal of the Royal Statistical Society, 76(3):

340–348, February 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340061.

Anonymous:1913:SEAb

[451] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76(3): 349–351, February 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340062.

Anonymous:1913:MLAb

[452] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(3):352–355, February 1913. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340063.

Anonymous:1913:PRb

[453] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 76(3):356–358, February 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340064.

Hutchinson:1913:PCC

[454] Lincoln Hutchinson. The Panama canal and competition for trade in Latin America, the Orient and Australasia. Journal of the Royal Statistical Society, 76(4):359–395, March 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340266.

Sauerbeck:1913:PC

[455] A. Sauerbeck. Prices of commodities in 1912. Journal of the Royal Statistical Society, 76(4):396–409, March 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340267.

Johnson:1913:HMT

[456] Stewart Johnson. The health and medical treatment of the uninsured. Journal of the Royal Statistical Society, 76 (4):410–414, March 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340268.

Lehfeldt:1913:RII

[457] R. A. Lehfeldt. The rate of interest on investments in 1912. Journal of the Royal Statistical Society, 76 (4):415–416, March 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340269.

Y:1913:BRBd

[458] G. U. Y. Book review: Changes in Bodily form of Descendants of Immigrants, by Franz Boas. Journal of the Royal Statistical Society, 76(4): 417–418, March 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340270.

G:1913:BRBb

[459] M. G. Book review: Statistik der Heilbehandlung bei den Versicherungsanstalten und Zugelassene Kasseneinrichtungen der Inavalidenversicherung für die Jahre 1906–11. Journal of the Royal Statistical Society, 76(4):419, March 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340271.

G:1913:BRBc

[460] M. G. Book review: Arzt und R. V. O. Der Arzt und die Deutsche Reichversicherungsordnung, by Th. Rumpf. Journal of the Royal Statistical Society, 76(4):419, March 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340272.

C:1913:BRBe

[461] S. J. C. Book review: The Cotton Manufacturing Industry of the United States, by M. T. Copeland; Handel und Produktion in der Baumwollindustrie by Edgar Landauer. Journal of the Royal Statistical Society, 76(4): 419–420, March 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340273.

Anonymous:1913:ONPc

[462] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(4):421–422, March 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340274.

Anonymous:1913:CNc

[463] Anonymous. Current notes. Journal of the Royal Statistical Society, 76 (4):423–428, March 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340275.

Anonymous:1913:SEAc

[464] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76 (4):429–431, March 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340276.

Anonymous:1913:MLAc

[465] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(4):432–436, March 1913. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340277.

Anonymous:1913:PRc

[466] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 76 (4):437–444, March 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340278.

Snow:1913:SSP

[467] E. C. Snow. Some statistical problems suggested by the sickness and mortality data of certain of the large friendly societies. Journal of the Royal Statistical Society, 76(5):445–517, April 1913. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340430.

Anonymous:1913:CNU

[468] Anonymous. Census notes: — the Union of South Africa. Journal of the Royal Statistical Society, 76(5): 518–519, April 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340431.

Jones:1913:AIE

[469] David Caradog Jones. An account of an inquiry into the extent of economic moral failure among certain types of regular workers. Journal of the Royal Statistical Society, 76(5):520–533, April 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340432.

B:1913:BRBc

[470] J. A. B. Book review: La Question de la Population, by Paul Leroy Beaulieu. Journal of the Royal Statistical Society, 76(5):534–536, April 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340433.

M:1913:BRBa

[471] H. W. M. Book review: The Distribution of Incomes in the United States, by F. H. Streightoff; Studies in History, Economics, and Public-Law. Journal of the Royal Statistical Society, 76 (5):536–538, April 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340434.

A:1913:BRBe

[472] P. A. Book review: Les Revenus de L’État, by N. G. Pierson. Journal of the Royal Statistical Society, 76 (5):538–540, April 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340435.

Anonymous:1913:ONPd

[473] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(5):540–543, April 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340436.

Anonymous:1913:Ca

[474] Anonymous. Correspondence. Journal of the Royal Statistical Society, 76 (5):544–546, April 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340437.

Anonymous:1913:CNd

[475] Anonymous. Current notes. Journal of the Royal Statistical Society, 76 (5):547–550, April 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340438.

Anonymous:1913:SEAd

[476] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76 (5):551–554, April 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340439.

Anonymous:1913:MLAd

[477] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(5):555–556, April 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340440.

Flux:1913:GCP

[478] A. W. Flux. Gleanings from the census of production report. Journal of the Royal Statistical Society, 76 (6):557–598, May 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339864.

Chapman:1913:TCE

[479] S. J. Chapman and W. Abbott. The tendency of children to enter their fathers’ trades. Journal of the Royal Statistical Society, 76(6):599–604, May 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339865.

R:1913:BRBa

[480] R. H. R. Book review: Report of Departmental Commission on the Official Statistics of Canada. Journal of the Royal Statistical Society, 76 (6):605–607, May 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339866.

L:1913:BRBa

[481] W. T. L. Book review: Gold, Prices and Wages, with an Examination of the Quantity Theory, by J. A. Hobson. Journal of the Royal Statistical Society, 76(6):607–609, May 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339867.

L:1913:BRBb

[482] W. T. L. Book review: Etudes sur la Formation et le Mouvement des Prix, by Marcel Lenoir. Journal of the Royal Statistical Society, 76(6):609–610, May 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339868.

P:1913:BRBb

[483] L. L. P. Book review: Historie Financiere et Economique de l’Angleterre (1066–1902), by Etienne Martin. Journal of the Royal Statistical Society, 76 (6):610–612, May 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339869.

P:1913:BRBc

[484] L. L. P. Book review: Historie de l’Impot Depuis l’Antiquite Jusqu’a nos Jours, by A. Wagner; H. Deite. Journal of the Royal Statistical Society, 76 (6):612–613, May 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339870.

F:1913:BRBb

[485] A. W. F. Book review: Die Grosstadtischen Agglomerationen Des Deutschen Reichs, by Sigmund Schott; Die Bevolkerungsbewegung der Deutschen Grosstadte seit Der Grundung Des Deutschen Reiches, by Pankraz Dittman. Journal of the Royal Statistical Society, 76(6):613–614, May 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339871.

G:1913:BRBd

[486] M. G. Book review: Krankheit und Soziale Lage, by M. Mosse; G. Tugendreich. Journal of the Royal Statistical Society, 76(6):614–616, May 1913.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2339872.

Anonymous:1913:ONPe

[487] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(6):616–618, May 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339873.

Anonymous:1913:Cb

[488] Anonymous. Correspondence. Journal of the Royal Statistical Society, 76 (6):619–622, May 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339874.

Anonymous:1913:CNe

[489] Anonymous. Current notes. Journal of the Royal Statistical Society, 76 (6):623–627, May 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339875.

Anonymous:1913:SEAe

[490] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76 (6):628–629, May 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339876.

Anonymous:1913:MLAe

[491] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(6):630–634, May 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339877.

Thompson:1913:CI

[492] William J. Thompson. The census of Ireland, 1911. Journal of the Royal Statistical Society, 76(7):635–671, June 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339707.

Bowley:1913:WCH

[493] A. L. Bowley. Working-class households in reading. Journal of the Royal Statistical Society, 76(7):672–701, June 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339708.

Nixon:1913:ETN

[494] J. W. Nixon. An experimental test of the normal law of error. Journal of the Royal Statistical Society, 76 (7):702–706, June 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339709.

Craigie:1913:AF

[495] P. G. Craigie. Alfred de Foville. Journal of the Royal Statistical Society, 76 (7):707–708, June 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339710.

Anonymous:1913:RHL

[496] Anonymous. The Right Hon. Lord Avebury, F.R.S. Journal of the Royal Statistical Society, 76(7):709–710, June 1913. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339711.

Y:1913:BRBe

[497] G. U. Y. Book review: La Théorie de l’Homme Moyen: Essai Sur Quetelet et la Statistique Morale, by Maurice Halbwachs. Journal of the Royal Statistical Society, 76(7):711–712, June 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339712.

S:1913:BRBc

[498] J. C. S. Book review: Income Taxation-Methods and Results in Various Countries, by Kossuth Kent Kennan. Journal of the Royal Statistical Society, 76 (7):712–714, June 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339713.

D:1913:BRBa

[499] R. D. Book review: Dell’ Emigrazione Italiana [Cinquanta Anni di Storia Italiana (1860–1910)], by Francesco Colletti. Journal of the Royal Statistical Society, 76(7):714–716, June 1913. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339714.

B:1913:BRBd

[500] E. B. Book review: Essai sur le Problème de la Misère. L’aide Sociale Aux Nécessiteux Adultes Valides, by Anatole Weber. Journal of the Royal Statistical Society, 76(7):717–718, June 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339715.

B:1913:BRBe

[501] A. L. B. Book review: Vers le Salaire Minimum. Etude d’économie et de Législation Industrielles, by Barthélemy Raynaud. Journal of the Royal Statistical Society, 76(7): 718–719, June 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339716.

B:1913:BRBf

[502] A. L. B. Book review: La Classe Ouvrière et les niveaux de vie. Recherches sur la Hiérarchie Des Besoins dans less Sociétiés Industrielles Contemporaines, by Maurice Halbwachs. Journal of the Royal Statistical Society, 76 (7):719–720, June 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339717.

B:1913:BRBg

[503] A. L. B. Book review: Arbeitsolhn und Arbeitszeit in Europa und Amerika, 1870–1909, by von R. Kuczynski; Julius Springer. Journal of the Royal Statistical Society, 76(7):720, June 1913. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2339718.

B:1913:BRBh

[504] E. B. Book review: English Local Government: The Story of the King’s Highway, by Sidney Webb; Beatrice Webb. Journal of the Royal Statistical Society, 76(7):720–722, June 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339719.

Anonymous:1913:ONPf

[505] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(7):722–724, June 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339720.

Anonymous:1913:CNf

[506] Anonymous. Current notes. Journal of the Royal Statistical Society, 76 (7):725–732, June 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339721.

Anonymous:1913:SEAf

[507] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76 (7):733–735, June 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339722.

Anonymous:1913:MLAf

[508] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(7):736–738, June 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339723.

Anonymous:1913:VIa

[509] Anonymous. Volume information. Journal of the Royal Statistical Society, 76(8):855–867, July 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339778.

Rosenbaum:1913:TBE

[510] S. Rosenbaum. The trade of the British empire. Journal of the Royal Statistical Society, 76(8):739–774, July 1913. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339779.

Anonymous:1913:RCF

[511] Anonymous. Report of the Council for the financial year ended December 31, 1912, and for the sessional year ending June 17, 1913, presented at the seventy-ninth Annual General Meeting of the Royal Statistical Society, held at the rooms of the Royal Society of Arts, John Street, Adelphi, London, W.C., on June 17, 1913. Journal of the Royal Statistical Society, 76(8):775–785, July 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339780.

Anonymous:1913:PSN

[512] Anonymous. Proceedings of the seventy-ninth Annual General Meeting, held on June 17, 1913. Journal of the Royal Statistical Society, 76(8):786, July 1913. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339781.

Price:1913:ECO

[513] L. L. Price. The estates of the colleges of Oxford and their management. Journal of the Royal Statistical Society, 76(8):787–790, July 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339782.

Anonymous:1913:RSC

[514] Anonymous. Report of the special committee on morbidity and mortality statistics in the United Kingdom. Journal of the Royal Statistical Society, 76(8):791–812, July 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339783.

Dudfield:1913:IMI

[515] R. Dudfield. On the increase of mortality from “injury at birth”. Journal of the Royal Statistical Society, 76 (8):813–814, July 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339784.

Y:1913:BRBf

[516] G. U. Y. Book review: Theorie und Methoden der Statistik, by Al. Kaufmann. Journal of the Royal Statistical Society, 76(8):815–817, July 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339785.

A:1913:BRBf

[517] P. A. Book review: Die Grundlagen der Schiffahrtsstatisik, by Walther Vogel. Journal of the Royal Statistical Society, 76(8):817–819, July 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339786.

W:1913:BRBe

[518] J. W. Book review: Report to the Board of Agriculture and Fisheres of an Enquiry into Agricultural Credit and Agricultural Co-operation in Germany, by J. R. Cahill. Journal of the Royal Statistical Society, 76(8):819–821, July 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339787.

S:1913:BRBd

[519] G. F. S. Book review: Indian Currency and Finance, by J. M. Keynes. Journal of the Royal Statistical Society, 76(8):821–824, July 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339788.

A:1913:BRBg

[520] P. A. Book review: Die Finanz- und Zollpolitik des Deutschen Reiches, by Wilhelm Gerloff. Journal of the Royal Statistical Society, 76(8):824–826, July 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339789.

B:1913:BRBi

[521] E. B. Book review: The Spirit of Association: Being Some Account of the Gilds, Friendly Societies, Co-operative Movement, and Trade Unions of Great Britain, by M. Fothergill Robinson. Journal of the Royal Statistical Society, 76(8):826–827, July 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339790.

M:1913:BRBb

[522] H. W. M. Book review: Le Syndicalisme et la Prochaine Révolution, by M.

Dufour. Journal of the Royal Statistical Society, 76(8):827–829, July 1913. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2339791.

P:1913:BRBd

[523] L. L. P. Book review: La Réglementation du Travail Réalisée ou Projetée: Ses Illusions et ses Dangers, by Edouard Payen. Journal of the Royal Statistical Society, 76 (8):829–830, July 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339792.

G:1913:BRBe

[524] M. G. Book review: Progress and Uniformity in Child-Labour Legislation, by William F. Ogburn; Faculty of Political Science of Columbia University. Journal of the Royal Statistical Society, 76 (8):830–831, July 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339793.

Anonymous:1913:ONPg

[525] Anonymous. Other new publications. Journal of the Royal Statistical Society, 76(8):831–832, July 1913. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339794.

Anonymous:1913:CNg

[526] Anonymous. Current notes. Journal of the Royal Statistical Society, 76 (8):833–839, July 1913. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339795.

Anonymous:1913:SEAg

[527] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 76 (8):840–842, July 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339796.

Anonymous:1913:MLAg

[528] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 76(8):843–845, July 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339797.

Anonymous:1913:ALA

[529] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 76(8):846–854, July 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339798.

Anonymous:1913:VIb

[530] Anonymous. Volume information. Journal of the Royal Statistical Society, 77(1):i–viii, December 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339757.

Wood:1913:CRW

[531] Frances Wood. The course of real wages in London, 1900–12. Journal of the Royal Statistical Society, 77(1): 1–68, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339758.

Secretaries:1913:FSI

[532] Honorary Secretaries. The fourteenth session of the International Statistical Institute. Journal of the Royal Statistical Society, 77(1):69–78, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339759.

Slutsky:1913:CGF

[533] E. Slutsky. On the criterion of goodness of fit of the regression lines and on the best method of fitting them to the data. Journal of the Royal Statistical Society, 77(1):78–84, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339760.

Hutchins:1913:IMP

[534] B. L. Hutchins. Infantile mortality and the proportion of the sexes. Journal of the Royal Statistical Society, 77(1): 84–86, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339761.

D:1913:BRBb

[535] R. D. Book review: Seventy-Fourth Annual Report of the Registrar-General of Births, Deaths and Marriages in England and Wales. Journal of the Royal Statistical Society, 77(1): 87–89, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339762.

W:1913:BRBf

[536] A. A. W. Book review: Grundriss der Statistik: Enthaltend Bevölkerungs- , Wirtschafts-Finanz,-und Handels-Statistik, by Carl Ballod. Journal of the Royal Statistical Society, 77(1): 89–90, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339763.

W:1913:BRBg

[537] A. A. W. Book review: Inleiding tot de Beoefening der Statistiek, by C. A. Verrijn Stuart. Journal of the Royal Statistical Society, 77(1):90–91, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339764.

S:1913:BRBe

[538] J. C. S. Book review: English Taxation, 1640–1799: An Essay on Policy and Opinion, by William Kennedy. Journal of the Royal Statistical Society, 77 (1):91–93, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339765.

R:1913:BRBb

[539] R. H. R. Book review: English Farming, Past and Present, by Rowland E. Prothero. Journal of the Royal Statistical Society, 77(1):93–95, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339766.

R:1913:BRBc

[540] R. H. R. Book review: The Occupying Ownership of Land, by Bevil Tollemache. Journal of the Royal Statistical Society, 77(1):95–96, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339767.

B:1913:BRBj

[541] A. L. B. Book review: “Official” Finance in Government Departments. A Series of Critical Comments Upon the Business and Financial Methods of Government Departments, by Herbert H. Bassett. Journal of the Royal Statistical Society, 77(1):96–97, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339768.

E:1913:BRB

[542] C. L. E. Book review: Outlines of Railways Economics, by Douglas Knoop; The Case Against Railway Nationalisation, by Edwin A. Pratt. Journal of the Royal Statistical Society, 77(1): 97–98, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339769.

P:1913:BRBe

[543] L. L. P. Book review: La Politique Économique Internationale, by Rudolf Kobatsch. Journal of the Royal Statistical Society, 77(1):99–100, December 1913. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2339770.

B:1913:BRBk

[544] J. A. B. Book review: Advance India!, by M. de P. Webb. Journal of the Royal Statistical Society, 77(1): 100–101, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339771.

S:1913:BRBf

[545] E. C. S. Book review: La Trasmissione Ereditaria delle Professioni, by Federico Chessa. Journal of the Royal Statistical Society, 77(1):101–102, December 1913. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339772.

G:1913:BRBf

[546] M. G. Book review: Das Sterile Berlin. Eine Volkwirtschaftliche Studie, by Felix A. Theilhaber. Journal of the Royal Statistical Society, 77(1): 103–104, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339773.

Anonymous:1913:ONPh

[547] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(1):104–110, December 1913. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339774.

Anonymous:1913:CNh

[548] Anonymous. Current notes. Journal of the Royal Statistical Society, 77(1): 111–124, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339775.

Anonymous:1913:SEAh

[549] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77(1): 125–131, December 1913. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339776.

Anonymous:1913:LAL

[550] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 77(1):132–144, December 1913. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339777.

Wilson:1914:COI

[551] James Wilson. The co-operative insurance of live-stock in England and Wales. Journal of the Royal Statistical Society, 77(2):145–158, January 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339799.

Robertson:1914:SMS

[552] D. H. Robertson. Some material for a study of trade fluctuations. Journal of the Royal Statistical Society, 77 (2):159–178, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339800.

Greenwood:1914:DSF

[553] M. Greenwood and G. Udny Yule. On the determination of size of family and of the distribution of characters in order of birth from samples taken through members of the sibships. Journal of the Royal Statistical Society, 77 (2):179–199, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339801.

Stamp:1914:NIP

[554] J. C. Stamp. A new illustration of Pareto’s law. Journal of the Royal Statistical Society, 77(2):200–204, January 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339802.

Welton:1914:NCI

[555] Thomas A. Welton. Notes on the census of Ireland, 1911. Journal of the Royal Statistical Society, 77(2): 205–213, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339803.

L:1914:BRBa

[556] W. T. L. Book review: Report of an Enquiry, by the Board of Trade into Working-Class Rents and Retail Prices, Together with the Rates and Wages in Certain Occupations in Industrial Towns of the United Kingdom in 1912. Journal of the Royal Statistical Society, 77(2):214–216, January 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339804.

S:1914:BRBa

[557] E. C. S. Book review: Die Statistische Methode als Selbständige Wissenschaft, by Hugo Forcher. Journal of the Royal Statistical Society, 77(2): 216–217, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339805.

P:1914:BRBa

[558] L. L. P. Book review: Les Crises Périodiques de Surproduction, by Albert Aftalion; Good and Bad Trade: An Inquiry into the Causes of Trade Fluctuations by R. G. Hawtrey; Business Cycles by Wesley Clair Mitchell. Journal of the Royal Statistical Society, 77(2):217–221, January 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339806.

S:1914:BRBb

[559] E. C. S. Book review: A History of Emigration from the United Kingdom to North America, 1763–1912, by Stanley C. Johnson. Journal of the Royal Statistical Society, 77(2):221–223, January 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339807.

C:1914:BRBa

[560] P. G. C. Book review: An Agricultural Faggot, by R. H. Rew. Journal of the Royal Statistical Society, 77(2): 223–224, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339808.

A:1914:BRBa

[561] W. M. A. Book review: British Railways: A Financial and Commercial Survey, by W. R. Lawson. Journal of the Royal Statistical Society, 77(2): 224–226, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339809.

A:1914:BRBb

[562] P. A. Book review: Germany’s Economic Progress and National Wealth, 1888–1913, by Karl Helfferich. Journal of the Royal Statistical Society, 77 (2):226–228, January 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339810.

B:1914:BRBa

[563] S. L. B. Book review: The German Cotton Industry, by R. M. R. Dehn. Journal of the Royal Statistical Society, 77(2):228–229, January 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339811.

P:1914:BRBb

[564] L. L. P. Book review: Théorie de la Valeur, by Christian Cornelissen. Journal of the Royal Statistical Society, 77 (2):229–231, January 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339812.

A:1914:BRBc

[565] P. A. Book review: Versicherungswesen, by Alfred Manes. Journal of the Royal Statistical Society, 77(2): 231–232, January 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339813.

M:1914:BRBa

[566] B. M. Book review: Essays in Taxation, by Edwin R. A. Seligman, McVickar. Journal of the Royal Statistical Society, 77(2):232–233, January 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2339814.

Anonymous:1914:ONPa

[567] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(2):233–235, January 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339815.

Anonymous:1914:CNa

[568] Anonymous. Current notes. Journal of the Royal Statistical Society, 77(2): 236–243, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339816.

Anonymous:1914:SEAa

[569] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77(2): 244–246, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339817.

Anonymous:1914:MLAa

[570] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(2):247–251, January 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339818.

Anonymous:1914:PRa

[571] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 77 (2):252–258, January 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339819.

Dunlop:1914:FMS

[572] James Craufurd Dunlop. The fertility of marriage in Scotland: A census study. Journal of the Royal Statistical Society, 77(3):259–299, February 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339724.

Edgeworth:1914:UAGa

[573] F. Y. Edgeworth. On the use of analytical geometry to represent certain kinds of statistics. Journal of the Royal Statistical Society, 77(3):300–312, February 1914. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339725.

Snow:1914:NPS

[574] E. C. Snow. Note on a possible source of fallacy in the interpretation of the census figures relating to the fertility of marriages. Journal of the Royal Statistical Society, 77(3):313–316, February 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339726.

Brown:1914:SIC

[575] J. W. Brown, M. Greenwood, Jr., and Frances Wood. A study of index correlations. Journal of the Royal Statistical Society, 77(3):317–346, February 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339727.

B:1914:BRBb

[576] J. A. B. Book review: Statistics, by Robert Giffen; Henry Higgs; G. Udny Yule. Journal of the Royal Statistical Society, 77(3):347–349, February 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339728.

G:1914:BRBa

[577] M. G. Book review: The English Convict: A Statistical Study, by Charles Goring; Schedule of Measurments and General Anthropological Data. Journal of the Royal Statistical Society, 77(3): 349–353, February 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339729.

V:1914:BRB

[578] H. D. V. Book review: Die Deutsche Landwirtschaft. Journal of the Royal Statistical Society, 77(3): 353–354, February 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339730.

P:1914:BRBc

[579] L. L. P. Book review: The Land and the Commonwealth, by T. E. Marks; J. Hugh Edwards. Journal of the Royal Statistical Society, 77(3): 354–355, February 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339731.

P:1914:BRBd

[580] L. L. P. Book review: L’Évolution du Commerce du Crédit et des Transports Depuis Cent Cinquante ans, by B. Nogaro; W. Oualid. Journal of the Royal Statistical Society, 77(3): 355–356, February 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339732.

P:1914:BRBe

[581] L. L. P. Book review: Éléments d’Économie Politique–Répartition– Consommation–Doctrines, by Bertrand Nogaro. Journal of the Royal Statistical Society, 77(3):356–357, February 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339733.

B:1914:BRBc

[582] S. L. B. Book review: Il Progresso Economico, by Napoleone Colajanni. Journal of the Royal Statistical Society, 77(3):357–359, February 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339734.

S:1914:BRBc

[583] J. C. S. Book review: The Nature and First Principle of Taxation, by Robert Jones; Sidney Webb. Journal of the Royal Statistical Society, 77(3): 359–360, February 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339735.

Anonymous:1914:ONPb

[584] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(3):360–363, February 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2339736.

Anonymous:1914:CNb

[585] Anonymous. Current notes. Journal of the Royal Statistical Society, 77(3): 364–370, February 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339737.

Anonymous:1914:SEAb

[586] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77(3): 371–373, February 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2339738.

Anonymous:1914:MLAb

[587] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(3):374–377, February 1914. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2339739.

Anonymous:1914:PRb

[588] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 77(3):378–380, February 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339740.

Baines:1914:CE

[589] J. A. Baines. The census of the empire, 1911. Journal of the Royal Statistical Society, 77(4):381–414, March 1914. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340240.

Edgeworth:1914:UAGb

[590] F. Y. Edgeworth. On the use of analytical geometry to represent certain kinds of statistics. Journal of the Royal Statistical Society, 77(4):415–432, March 1914. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340241.

Lehfeldt:1914:RIB

[591] R. A. Lehfeldt. The rate of interest on British and foreign investments. Journal of the Royal Statistical Society, 77 (4):432–435, March 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340242.

F:1914:BRB

[592] C. R. F. Book review: How the Labourer Lives, by B. Seebohm Rowntree; May Kendall; Les Ouvriers Agricoles en Belgique, by Benoit Bouché. Journal of the Royal Statistical Society, 77(4):436–438, March 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340243.

D:1914:BRB

[593] J. M. D. Book review: Co-Operation in Agriculture, by G. Harold Powell. Journal of the Royal Statistical Society, 77 (4):439–441, March 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340244.

M:1914:BRBb

[594] H. W. M. Book review: L’Industrie et les Industriels, by Yves Guyot. Journal of the Royal Statistical Society, 77 (4):441–442, March 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340245.

C:1914:BRBb

[595] S. J. C. Book review: Les Crises Industrielles en Angleterre, by Michel Tougan-Baranowsky; Joseph Schapiro. Journal of the Royal Statistical Society, 77(4):443, March 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340246.

Anonymous:1914:ONPc

[596] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(4):444–446, March 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340247.

Anonymous:1914:CNc

[597] Anonymous. Current notes. Journal of the Royal Statistical Society, 77 (4):447–452, March 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340248.

Anonymous:1914:SEAc

[598] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77 (4):453–455, March 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340249.

Anonymous:1914:MLAc

[599] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(4):456–459, March 1914. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340250.

Anonymous:1914:PRc

[600] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 77 (4):460–467, March 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340251.

Chapman:1914:SBM

[601] S. J. Chapman and T. S. Ashton. The sizes of businesses, mainly in the textile industries. Journal of the Royal Statistical Society, 77(5):469–555, April 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340065.

Paish:1914:PC

[602] George Paish. Prices of commodities in 1913. Journal of the Royal Statistical Society, 77(5):556–570, April 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340066.

B:1914:BRBd

[603] E. B. Book review: British Budgets, 1887–88 to 1912–13, by Bernard Mallet. Journal of the Royal Statistical Society, 77(5):571–573, April 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340067.

R:1914:BRBa

[604] R. H. R. Book review: A Pilgrimage of British Farming, 1910–12, by A. D. Hall. Journal of the Royal Statistical Society, 77(5):573–574, April 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340068.

G:1914:BRBb

[605] M. G. Book review: Medico-Actuarial Mortality Investigation, by Association of Life Assurance Medical Directors and the Actuarial Society of America. Journal of the Royal Statistical Society, 77(5):575–576, April 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340069.

P:1914:BRBf

[606] L. L. P. Book review: Production: A Study in Economics, by P. H. Castberg. Journal of the Royal Statistical Society, 77(5):576–577, April 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340070.

L:1914:BRBb

[607] W. T. L. Book review: Co-Operation and Co-Partnership, by L. L. Price. Journal of the Royal Statistical Society, 77(5):577–578, April 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340071.

Anonymous:1914:ONPd

[608] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(5):579–581, April 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340072.

Anonymous:1914:CNd

[609] Anonymous. Current notes. Journal of the Royal Statistical Society, 77 (5):582–587, April 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340073.

Anonymous:1914:SEAd

[610] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77 (5):588–591, April 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340074.

Anonymous:1914:MLAd

[611] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(5):592–595, April 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340075.

Bowley:1914:RPE

[612] A. L. Bowley. Rural population in England and Wales: A study of the changes of density, occupations, and ages. Journal of the Royal Statistical Society, 77 (6):597–652, May 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340305.

Edgeworth:1914:UAGc

[613] F. Y. Edgeworth. On the use of analytical geometry to represent certain kinds of statistics. Journal of the Royal Statistical Society, 77(6):653–671, May 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340306.

B:1914:BRBe

[614] J. A. B. Book review: Census of India, 1911; Vol. I. Report, by E. A. Gait. Journal of the Royal Statistical Society, 77(6):672–675, May 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340307.

S:1914:BRBd

[615] J. C. S. Book review: Social Reform as Related to Realities and Delusions — An Examination of the Increase and Distribution of Wealth from 1801 to 1910, by W. H. Mallock. Journal of the Royal Statistical Society, 77 (6):675–678, May 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340308.

P:1914:BRBg

[616] L. L. P. Book review: The Land: The Report of the Land Enquiry Committee. Vol. I. Rural. Journal of the Royal Statistical Society, 77(6):678–680, May 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340309.

B:1914:BRBf

[617] E. B. Book review: The Land. The Report of the Land Enquiry Committee. Vol. II: Urban. Journal of the Royal Statistical Society, 77(6):680–683, May 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340310.

P:1914:BRBh

[618] L. L. P. Book review: The Land Retort, by Charles Adeane; Edwin Savill. Journal of the Royal Statistical Society, 77(6):683–684, May 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340311.

W:1914:BRBa

[619] A. D. W. Book review: Economie Politique et Statistique, by Ch. Lordier. Journal of the Royal Statistical Society, 77(6):685–686, May 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340312.

A:1914:BRBd

[620] P. A. Book review: L’Allemagne en France, by Louis Bruneau. Journal of the Royal Statistical Society, 77 (6):686–688, May 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340313.

A:1914:BRBe

[621] W. M. A. Book review: American Railroad Economics: A Text-Book for Investors and Students, by A. M. Sakolski. Journal of the Royal Statistical Society, 77(6):688–689, May 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340314.

Anonymous:1914:ONPe

[622] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(6):689–690, May 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340315.

Anonymous:1914:CNe

[623] Anonymous. Current notes. Journal of the Royal Statistical Society, 77 (6):691–696, May 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340316.

Anonymous:1914:SEAe

[624] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77 (6):697–699, May 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340317.

Anonymous:1914:MLAe

[625] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(6):700–704, May 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340318.

Hazell:1914:SRL

[626] Walter Hazell. Suggestions for recording the life history and family connections of every individual. Journal of the Royal Statistical Society, 77 (7):705–723, June 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340295.

Edgeworth:1914:UAGd

[627] F. Y. Edgeworth. On the use of analytical geometry to represent certain kinds of statistics. Journal of the Royal Statistical Society, 77(7):724–749, June 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340296.

A:1914:BRBf

[628] P. A. Book review: Final Report of the Departmental Committee on Local Taxation. Journal of the Royal Statistical Society, 77(7):750–754, June 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340297.

E:1914:BRB

[629] F. Y. E. Book review: Les Mathématiques Appliquées à l’Économie Politique, by Wl. Zawadzki. Journal of the Royal Statistical Society, 77 (7):754–757, June 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340298.

S:1914:BRBe

[630] J. C. S. Book review: La Banque en France (Considérée Principalement au Point de vue des trois Grandes Banques de Dépôts), by E. Kaufmann; A. S. Sacker. Journal of the Royal Statistical Society, 77(7):757–758, June 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340299.

W:1914:BRBb

[631] A. D. W. Book review: Der Teilbau nebst der Monographie eines Teilbaugrossbetriebs in Russland aus der Zeit von 1891–1910, by E. Jenny; Bauernfrage und Agrarreform in Russland. Unter Besonderer Berücksichtigung der Nach Dem Ukas vom 9 November 1906 und dem Gesetz vom 29 Mai 1911 Eingeleiteten Auflösung des “Mirs” und Auseinandersetzung der Bäuerlichen Landanteile, by K. A. Wieth-Knudsen. Journal of the Royal Statistical Society, 77(7):758–759, June 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340300.

Anonymous:1914:ONPf

[632] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(7):760–762, June 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340301.

Anonymous:1914:CNf

[633] Anonymous. Current notes. Journal of the Royal Statistical Society, 77 (7):763–769, June 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340302.

Anonymous:1914:SEAf

[634] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77 (7):770–773, June 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340303.

Anonymous:1914:MLAf

[635] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(7):774–776, June 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340304.

Anonymous:1914:VI

[636] Anonymous. Volume information. Journal of the Royal Statistical Society, 77(8):925–939, July 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340923.

Crammond:1914:ERB

[637] Edgar Crammond. The economic relations of the British and German empires. Journal of the Royal Statistical Society, 77(8):777–824, July 1914. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340924.

Anonymous:1914:RCF

[638] Anonymous. Report of the Council for the financial year ended December 31, 1913, and for the sessional year ending June 16, 1914, presented at the eightieth Annual General Meeting of the Royal Statistical Society, held at the rooms of the Royal Society of Arts, John Street, Adelphi, London, W.C., on June 16, 1914. Journal of the Royal Statistical Society, 77(8):825–836, July 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340925.

Anonymous:1914:PEA

[639] Anonymous. Proceedings of the eightieth Annual General Meeting, held on June 16, 1914. Journal of the Royal Statistical Society, 77(8):837, July 1914. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340926.

Edgeworth:1914:UAGe

[640] F. Y. Edgeworth. On the use of analytical geometry to represent certain kinds of statistics. Journal of the Royal Statistical Society, 77(8):838–852, July 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340927.

Nixon:1914:SFA

[641] J. W. Nixon. Some factors associated with the illegitimate birth-rate. Journal of the Royal Statistical Society, 77 (8):852–862, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340928.

P:1914:BRBi

[642] L. L. P. Book review: Work and Wealth: A Human Valuation, by J. A. Hobson. Journal of the Royal Statistical Society, 77(8):863–865, July 1914. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340929.

S:1914:BRBf

[643] E. C. S. Book review: Science and Method, by Henri Poincaré; Francis Maitland; Bertrand Russell. Journal of the Royal Statistical Society, 77 (8):865–866, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340930.

G:1914:BRBc

[644] M. G. Book review: Le Hasard, by Émile Borel. Journal of the Royal Statistical Society, 77(8):866–867, July 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340931.

Y:1914:BRBa

[645] G. U. Y. Book review: Le Jeu, la Chance et le Hasard, by Louis Bachelier. Journal of the Royal Statistical Society, 77(8):867–868, July 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340932.

W:1914:BRBc

[646] A. D. W. Book review: Gewerbestatistik, by A. Hesse; J. Conrad. Journal of the Royal Statistical Society, 77 (8):868–869, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340933.

W:1914:BRBd

[647] A. D. W. Book review: Les Employés en Allemagne. Leur Situation Sociale et Leur Contrat d’Emploi, Législation et Revendications, by J. Buelens; Les Employés en Autriche. Leur Situation et Leur Contrat d’Emploi, by J. Buelens. Journal of the Royal Statistical Society, 77(8):869–871, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340934.

G:1914:BRBd

[648] M. G. Book review: Gesundheitliche Schädigungen der Frau bei der Industriellen Arbeit unter Besonderer Berucksichtigung Einiger Betriebe, by Käte Winkelmann. Journal of the Royal Statistical Society, 77(8):871, July 1914. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340935.

B:1914:BRBg

[649] E. B. Book review: The State and the Poor, by Geoffrey Drage. Journal of the Royal Statistical Society, 77 (8):871–872, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340936.

Y:1914:BRBb

[650] G. U. Y. Book review: The Influence of Monarchs: Steps in a New Science of History, by F. A. Woods. Journal of the Royal Statistical Society, 77 (8):872–875, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340937.

B:1914:BRBh

[651] A. L. B. Book review: Economic Notes on English Agricultural Wages, by Reginald Lennard. Journal of the Royal Statistical Society, 77(8):875–876, July 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340938.

R:1914:BRBb

[652] R. H. R. Book review: The English Land System, by J. A. R. Marriott. Journal of the Royal Statistical Society, 77(8):876–877, July 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340939.

C:1914:BRBc

[653] S. J. C. Book review: The Relations of Capital and Labour, by W. T. Layton. Journal of the Royal Statistical Society, 77(8):877, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340940.

R:1914:BRBc

[654] D. H. R. Book review: The Cause of Business Depressions, by Hugo Bilgram; Louis Edward Levy. Journal of the Royal Statistical Society, 77 (8):878–879, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340941.

P:1914:BRBj

[655] L. L. P. Book review: Les Règlements par Effets de Commerce en France at à l’Étranger, by Gaston Roulleau. Journal of the Royal Statistical Society, 77 (8):879–880, July 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340942.

M:1914:BRBc

[656] H. W. M. Book review: Corporate Promotions and Reorganisations, by Arthur S. Dewing. Journal of the Royal Statistical Society, 77(8):880–881, July 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340943.

B:1914:BRBi

[657] S. L. B. Book review: L’Industria Domestica Salariata nei Rapporti Interni e Internazionali, by Federico Marconcini; Achille Loria. Journal of the Royal Statistical Society, 77 (8):881–882, July 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340944.

B:1914:BRBj

[658] E. B. Book review: Social Insurance, by I. M. Rubinow. Journal of the Royal Statistical Society, 77(8):883, July 1914. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2340945.

P:1914:BRBk

[659] L. L. P. Book review: Principes d’Économie Pure, by B. Antonelli. Journal of the Royal Statistical Society, 77(8):884–885, July 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340946.

P:1914:BRBl

[660] L. L. P. Book review: Political Economy, by Charles Gide; William Smart; Constance H. M. Archibald. Journal of the Royal Statistical Society, 77 (8):885–886, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340947.

S:1914:BRBg

[661] E. C. S. Book review: Grundzüge Einer Theorie der Lebensversicherung, by N. R. Jörgensen. Journal of the Royal Statistical Society, 77(8):886–887, July 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340948.

C:1914:BRBd

[662] S. J. C. Book review: A Theory of Interest, by Clarence Gilbert Hoag. Journal of the Royal Statistical Society, 77(8):887, July 1914. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340949.

B:1914:BRBk

[663] S. L. B. Book review: I Prezzi nella Industria Cotoniera. Journal of the Royal Statistical Society, 77(8):888–889, July 1914. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340950.

P:1914:BRBm

[664] L. L. P. Book review: Problems of Empire, by Viscount Hythe; Earl Grey. Journal of the Royal Statistical Society, 77(8):889–890, July 1914. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340951.

Anonymous:1914:ONPg

[665] Anonymous. Other new publications. Journal of the Royal Statistical Society, 77(8):890–893, July 1914. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340952.

Anonymous:1914:CNg

[666] Anonymous. Current notes. Journal of the Royal Statistical Society, 77 (8):894–907, July 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340953.

Anonymous:1914:SEAg

[667] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 77 (8):908–910, July 1914. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340954.

Anonymous:1914:MLAg

[668] Anonymous. Monthly list of additions to the library. Journal of the Royal Statistical Society, 77(8):911–915, July 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340955.

Anonymous:1914:ALA

[669] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 77(8):916–924, July 1914. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340956.

Anonymous:1915:VIa

[670] Anonymous. Volume information. Journal of the Royal Statistical Society, 78(1):i–vii, January 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340834.

Welby:1915:PUK

[671] Lord Welby. The progress of the United Kingdom from the War of the French Revolution to 1913. Journal of the Royal Statistical Society, 78 (1):1–31, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340835.

Anonymous:1915:PN

[672] Anonymous. Proceedings on November 17, 1914. Journal of the Royal Statistical Society, 78(1):32–34, January 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340836.

Dunlop:1915:FNF

[673] James Craufurd Dunlop. A further note on the fertility of marriage in Scotland. Errors of statement as to the duration of marriage. Journal of the Royal Statistical Society, 78(1): 35–54, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340837.

Jones:1915:SNC

[674] D. Caradog Jones. Some notes on the census of occupations for England and Wales. Journal of the Royal Statistical Society, 78(1):55–81, January 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340838.

Anonymous:1915:FES

[675] Anonymous. Foreign exchanges since the outbreak of war. Journal of the Royal Statistical Society, 78(1): 82–87, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340839.

M:1915:BRBa

[676] H. W. M. Book review: Londres et les Ouvriers de Londres, by D. Pasquet.

Journal of the Royal Statistical Society, 78(1):88–90, January 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340840.

H:1915:BRBa

[677] G. H. Book review: Social Work in London, 1869 to 1912: A History of the Charity Organisation Society, by Helen Bosanquet. Journal of the Royal Statistical Society, 78(1):90–92, January 1915. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340841.

P:1915:BRBa

[678] L. L. P. Book review: The Future of Work and Other Essays, by L. G. Chiozza Money. Journal of the Royal Statistical Society, 78(1): 92–93, January 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340842.

S:1915:BRBa

[679] J. C. S. Book review: The Financial System of the United Kingdom, by Henry Higgs. Journal of the Royal Statistical Society, 78(1):93–94, January 1915. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340843.

S:1915:BRBb

[680] J. C. S. Book review: The Export of Capital, by C. K. Hobson. Journal of the Royal Statistical Society, 78 (1):94–96, January 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340844.

F:1915:BRB

[681] H. S. F. Book review: La Monnaie, le Crédit et le Change, by Aug. Arnauné.

Journal of the Royal Statistical Society, 78(1):97–100, January 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340845.

P:1915:BRBb

[682] L. L. P. Book review: The Influence of the Gold Supply on Prices and Profits, by David Barbour. Journal of the Royal Statistical Society, 78(1): 100–101, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340846.

S:1915:BRBc

[683] J. C. S. Book review: The Economic Synthesis: A Study of the Laws of Income, by Achille Loria; M. Eden Paul. Journal of the Royal Statistical Society, 78(1):101–103, January 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340847.

W:1915:BRBa

[684] A. D. W. Book review: Die Methoden der Deutschen Arbeitslosenstatistik, by Richard Herbst. Journal of the Royal Statistical Society, 78(1): 103–104, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340848.

J:1915:BRBa

[685] E. S. J. Book review: Armenwesen der Schweiz. Das Gesetzliche und Organisierte Freiwillige Armenwesen in der Schweiz. Herausgegeben von der Ständigen Kommission der Schweiz. I. Band. Das Gesetzliche Armenwesen in der Schweiz. Armenpfleger-Konferenzen Das Armenwesen des Bundes, sämtlicher Kantone und der Schweizerischen Grosstädte, by C. A.

Schmid; Armenwesen der Schweiz. Das Gesetzliche und Organisierte Freiwillige Schweiz. II. Band. Das Organisierte Freiwillige Armenwesen in der Schweiz by A. Wild Pfarrer Mönchaltorf. Journal of the Royal Statistical Society, 78(1):104–105, January 1915. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340849.

P:1915:BRBc

[686] L. L. P. Book review: The Ownership, Tenure and Taxation of Land: Some Facts, Fallacies and Proposals Relating Thereto, by Thomas P. Whittaker. Journal of the Royal Statistical Society, 78(1):105–107, January 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340850.

P:1915:BRBd

[687] L. L. P. Book review: Land and the Politicians, by Harman Grisewood; Ellis Robins. Journal of the Royal Statistical Society, 78(1):107–108, January 1915. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340851.

Anonymous:1915:ONPa

[688] Anonymous. Other new publications. Journal of the Royal Statistical Society, 78(1):108–111, January 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340852.

Anonymous:1915:CNa

[689] Anonymous. Current notes. Journal of the Royal Statistical Society, 78(1): 112–132, January 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340853.

Anonymous:1915:SEAa

[690] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 78(1): 133–139, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340854.

Anonymous:1915:LALa

[691] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 78(1):140–148, January 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340855.

Anonymous:1915:PRa

[692] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 78 (1):149–156, January 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340856.

Chapman:1915:WTI

[693] S. J. Chapman and David Kemp. The War and the textile industries. Journal of the Royal Statistical Society, 78 (2):157–237, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340607.

Snow:1915:MPE

[694] E. C. Snow. On the magnitude of the population of England and Wales available for emigration. Journal of the Royal Statistical Society, 78(2): 238–280, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340608.

Paish:1915:PC

[695] George Paish. Prices of commodities in 1914. Journal of the Royal Statistical Society, 78(2):281–295, March 1915.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340609.

Yule:1915:CPP

[696] G. Udny Yule. Crop production and price: A note on Gregory King’s law. Journal of the Royal Statistical Society, 78(2):296–298, March 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340610.

G:1915:BRBa

[697] M. G. Book review: The Life, Letters and Labours of Francis Galton, by Karl Pearson. Journal of the Royal Statistical Society, 78(2):299–302, March 1915. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340611.

Y:1915:BRB

[698] G. U. Y. Book review: Economic Cycles: Their Law and Cause, by H. L. Moore. Journal of the Royal Statistical Society, 78(2):302–305, March 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340612.

D:1915:BRB

[699] J. C. D. Book review: Vital Statistics Explained, by Joseph Burn. Journal of the Royal Statistical Society, 78 (2):305–306, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340613.

G:1915:BRBb

[700] M. G. Book review: Vorlesungen Ueber Soziale Medizin, by Ludwig Teleky. Journal of the Royal Statistical Society, 78(2):306–307, March 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340614.

G:1915:BRBc

[701] M. G. Book review: Report on the English Birth-Rate.Part 1. England North of the Humber, by Ethel M. Elderton. Journal of the Royal Statistical Society, 78(2):307, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340615.

S:1915:BRBd

[702] W. F. S. Book review: Modern Instruments and Methods of Calculation: A Handbook of the Napier Tercentenary Exhibition, by E. M. Horsburgh. Journal of the Royal Statistical Society, 78 (2):307–309, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340616.

P:1915:BRBe

[703] L. L. P. Book review: Work and Wages. Part III. Social Betterment, by Sydney J. Chapman; Earl Brassey. Journal of the Royal Statistical Society, 78(2):309–310, March 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340617.

P:1915:BRBf

[704] L. L. P. Book review: Farm Accounts, by C. S. Orwin. Journal of the Royal Statistical Society, 78(2): 310–311, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340618.

P:1915:BRBg

[705] L. L. P. Book review: Greater Profits from Land: The Secret of Successful Farming, by A. L. Drysdale. Journal of the Royal Statistical Society, 78 (2):312–313, March 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340619.

B:1915:BRBa

[706] S. L. B. Book review: Il Debito Pubblico Italiano: Consistenza– Operazioni–Formule–Legislazione, by Avvocato Enrico Bruni. Journal of the Royal Statistical Society, 78(2): 313–314, March 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340620.

P:1915:BRBh

[707] L. L. P. Book review: The Franco-German War Indemnity and its Economic Results, by Horace Handley O’Farrell. Journal of the Royal Statistical Society, 78(2):314–315, March 1915. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340621.

W:1915:BRBb

[708] A. D. W. Book review: Ocean Traffic and Trade, by B. Olney Hough. Journal of the Royal Statistical Society, 78 (2):315–317, March 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340622.

B:1915:BRBb

[709] P. E. B. Book review: Town Planning: with Special Reference to the Birmingham Schemes, by George Cadbury. Journal of the Royal Statistical Society, 78(2):317–318, March 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340623.

Anonymous:1915:ONPb

[710] Anonymous. Other new publications. Journal of the Royal Statistical Society, 78(2):318–322, March 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340624.

Anonymous:1915:CNb

[711] Anonymous. Current notes. Journal of the Royal Statistical Society, 78 (2):323–341, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340625.

Anonymous:1915:SEAb

[712] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 78 (2):342–344, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340626.

Anonymous:1915:LALb

[713] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 78(2):345–349, March 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340627.

Anonymous:1915:PRb

[714] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 78 (2):350–359, March 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340628.

Crammond:1915:CW

[715] Edgar Crammond. The cost of the War. Journal of the Royal Statistical Society, 78(3):361–413, May 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341056.

Brabrook:1915:PFS

[716] Edward Brabrook. On the progress of Friendly Societies and other provident institutions during the ten years 1904– 1914. Journal of the Royal Statistical Society, 78(3):414–445, May 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341057.

Snow:1915:NFP

[717] E. C. Snow. Note on the future population of the self-governing portion of the British empire. Journal of the Royal Statistical Society, 78(3):446–451, May 1915. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341058.

Lehfeldt:1915:RIB

[718] R. A. Lehfeldt. The rate of interest on British and foreign investments. Journal of the Royal Statistical Society, 78 (3):452–453, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341059.

Trachtenberg:1915:NPM

[719] M. I. Trachtenberg. A note on a property of the median. Journal of the Royal Statistical Society, 78(3):454, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341060.

W:1915:BRBc

[720] A. D. W. Book review: Livelihood and Poverty: A Study in the Economic Conditions of Working-Class Households in Northampton, Warrington, Stanley and Reading, by A. L. Bowley; A. R. Burnett-Hurst; R. H. Tawney. Journal of the Royal Statistical Society, 78(3):455–456, May 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341061.

C:1915:BRBa

[721] E. C. Book review: Property and Contract in their Relations to the Distribution of Wealth, by Richard T. Ely. Journal of the Royal Statistical Society, 78(3):456–458, May 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341062.

W:1915:BRBd

[722] A. D. W. Book review: Elementary Principles of Economics, Together with a Short Sketch of Economic History, by Richard T. Ely; George Ray Wicker; L. L. Price. Journal of the Royal Statistical Society, 78(3):458–459, May 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341063.

S:1915:BRBe

[723] W. T. S. Book review: Railroads: Rates and Regulation, by William Z. Ripley; Railroads: Finance and Organization by William Z. Ripley. Journal of the Royal Statistical Society, 78 (3):459–461, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341064.

S:1915:BRBf

[724] W. T. S. Book review: Railway Rates and Traffic, by L. R. Christie; G. Leedham; C. Travis. Journal of the Royal Statistical Society, 78 (3):461, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341065.

H:1915:BRBb

[725] A. H. Book review: Artificial Waterways of the World, by A. Barton Hepburn. Journal of the Royal Statistical Society, 78(3):461–463, May 1915.

ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341066.

P:1915:BRBi

[726] L. L. P. Book review: War and Lombard Street, by Hartley Withers; The War and our Financial Fabric, by Walter William Wall. Journal of the Royal Statistical Society, 78(3):463–465, May 1915. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341067.

S:1915:BRBg

[727] J. C. S. Book review: The System of National Finance, by E. Hilton Young. Journal of the Royal Statistical Society, 78(3):465–466, May 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341068.

M:1915:BRBb

[728] H. W. M. Book review: The Prevention and Control of Monopolies, by W. Jethro Brown. Journal of the Royal Statistical Society, 78(3):466–468, May 1915. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341069.

B:1915:BRBc

[729] C. F. B. Book review: Economics of Efficiency, by N. A. Brisco. Journal of the Royal Statistical Society, 78 (3):468–469, May 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341070.

M:1915:BRBc

[730] H. W. M. Book review: Municipal Life and Government in Germany, by William Harbutt Dawson. Journal of the Royal Statistical Society, 78 (3):469–470, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341071.

U:1915:BRB

[731] J. F. U. Book review: The World’s Cane Sugar Industry, Past and Present, by H. C. Prinsen. Journal of the Royal Statistical Society, 78(3):471, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341072.

B:1915:BRBd

[732] S. L. B. Book review: Valori e Prezzi nei Loro Rapporti con la Moneta, con la Banca e con la Borsa, by Giacomo Luzzatti. Journal of the Royal Statistical Society, 78(3):471–472, May 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341073.

Anonymous:1915:ONPc

[733] Anonymous. Other new publications. Journal of the Royal Statistical Society, 78(3):473–475, May 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341074.

Anonymous:1915:CNc

[734] Anonymous. Current notes. Journal of the Royal Statistical Society, 78 (3):476–489, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341075.

Anonymous:1915:SEAc

[735] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 78 (3):490–494, May 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341076.

Anonymous:1915:LALc

[736] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 78(3):495–500, May 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341077.

Anonymous:1915:VIb

[737] Anonymous. Volume information. Journal of the Royal Statistical Society, 78(4):673–683, July 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340748.

Rosenbaum:1915:EWO

[738] S. Rosenbaum. The effects of the War on the overseas trade of the United Kingdom. Journal of the Royal Statistical Society, 78(4):501–554, July 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340749.

Mallet:1915:MCW

[739] Bernard Mallet and H. C. Strutt. The multiplier and capital wealth. Journal of the Royal Statistical Society, 78 (4):555–599, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340750.

Anonymous:1915:RCF

[740] Anonymous. Report of the Council for the financial year ended December 31, 1914, and for the sessional year ending June 15, 1915, presented at the eighty-first Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C., on June 15, 1915. Journal of the Royal Statistical Society, 78(4):600–610, July 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340751.

Anonymous:1915:PEF

[741] Anonymous. Proceedings of the eightyfirst Annual General Meeting, held on Tuesday, June 15, 1915. Journal of the Royal Statistical Society, 78(4):611, July 1915. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340752.

H:1915:BRBc

[742] F. W. H. Book review: The Works and Life of Walter Bagehot, by Russell Barrington. Journal of the Royal Statistical Society, 78(4):612–614, July 1915. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340753.

C:1915:BRBb

[743] G. G. C. Book review: The Effect of the War on the External Trade of the United Kingdom, by A. L. Bowley. Journal of the Royal Statistical Society, 78(4):614–616, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340754.

G:1915:BRBd

[744] M. G. Book review: The Nature and Purpose of the Measurement of Social Phenomena, by A. L. Bowley. Journal of the Royal Statistical Society, 78 (4):616–617, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340755.

G:1915:BRBe

[745] M. G. Book review: Infant Mortality, by Hugh T. Ashby. Journal of the Royal Statistical Society, 78 (4):617–618, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340756.

R:1915:BRB

[746] S. R. Book review: The Audacious War, by C. W. Barron; Les Finances de Guerre de l’Angleterre. by Gaston Jeze; Problemes de Politique et Finances de Guerre. by MM. Jeze. Journal of the Royal Statistical Society, 78 (4):618–620, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340757.

B:1915:BRBe

[747] S. L. B. Book review: Principi Fondamentali di Scienza Pura delle Finanze. Saggio di un’ Organica Sistemazione Teorica delle Dottrine Finanziarie nel loro Duplice Aspetto Politico-Economico, by Roberto A. Murray. Journal of the Royal Statistical Society, 78(4):620–621, July 1915. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340758.

J:1915:BRBb

[748] G. J. Book review: Finance in China, by Shrinivas R. Wagel. Journal of the Royal Statistical Society, 78 (4):621–624, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340759.

P:1915:BRBj

[749] G. H. P. Book review: The Evolution of the Money Market (1385–1915). An Historical and Analytical Study of the Rise and Development of Finance as a Centralised, Co-ordinated Force, by Ellis T. Powell. Journal of the Royal Statistical Society, 78(4):624–625, July 1915. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340760.

S:1915:BRBh

[750] C. P. S. Book review: L’Emploi des Mathematiques en Economie Politique, by Jacques Moret. Journal of the Royal Statistical Society, 78(4):625–626, July 1915. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340761.

S:1915:BRBi

[751] W. T. S. Book review: The Panama Canal and International Trade Competition, by Lincoln Hutchinson. Journal of the Royal Statistical Society, 78 (4):626–627, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340762.

M:1915:BRBd

[752] H. W. M. Book review: The History and Economics of Transport, by Adam W. Kirkaldy; Alfred Dudley Evans. Journal of the Royal Statistical Society, 78(4):627–628, July 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340763.

C:1915:BRBc

[753] W. C. Book review: English Economic History, by A. E. Bland; P. A. Brown; R. H. Tawney. Journal of the Royal Statistical Society, 78(4):629, July 1915. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340764.

H:1915:BRBd

[754] B. L. H. Book review: The Establishment of Minimum Rates in the Tailoring Industry under the Trade Boards Act of 1909, by R. H. Tawney. Journal of the Royal Statistical Society, 78 (4):630–631, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340765.

H:1915:BRBe

[755] B. L. H. Book review: La Revolution Industrielle et les Origines de la Protection Legale du Travail en Suisse, by W. E. Rappard. Journal of the Royal Statistical Society, 78(4):631–632, July 1915. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340766.

C:1915:BRBd

[756] P. G. C. Book review: Fifty Years of Agricultural Politics, being the History of the Central Chamber of Agriculture, 1865–1915, by A. H. H. Matthews. Journal of the Royal Statistical Society, 78(4):632–633, July 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340767.

B:1915:BRBf

[757] P. E. B. Book review: Municipal Charters: A Discussion of the Essentials of a City Charter, with Forms or Models for Adoption, by Nathan Matthews. Journal of the Royal Statistical Society, 78(4):633–634, July 1915. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340768.

C:1915:BRBe

[758] E. C. Book review: Some Aspects of the Tariff Question, by F. W. Taussig. Journal of the Royal Statistical Society, 78(4):634–635, July 1915. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340769.

A:1915:BRB

[759] P. A. Book review: Die Schweiz und Die Europaische Handelspolitik, by Peter Heinrich Schmidt. Journal of the Royal Statistical Society, 78 (4):635–636, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340770.

Anonymous:1915:ONPd

[760] Anonymous. Other new publications. Journal of the Royal Statistical Society, 78(4):637–641, July 1915. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340771.

Anonymous:1915:CNd

[761] Anonymous. Current notes. Journal of the Royal Statistical Society, 78 (4):642–655, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340772.

Anonymous:1915:SEAd

[762] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 78 (4):656–659, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340773.

Anonymous:1915:BML

[763] Anonymous. Bi-monthly list of additions to the library. Journal of the Royal Statistical Society, 78(4): 660–664, July 1915. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340774.

Anonymous:1915:ALA

[764] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 78(4):665–672, July 1915. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340775.

Anonymous:1916:VIa

[765] Anonymous. Volume information. Journal of the Royal Statistical Society, 79(1):i–vii, January 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340629.

West:1916:RHL

[766] Algernon West. The Right Hon. Lord Welby, G.C.B. an appreciation. Journal of the Royal Statistical Society, 79 (1):1–6, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340630.

Boag:1916:HCC

[767] Harold Boag. Human capital and the cost of the War. Journal of the Royal Statistical Society, 79(1):7–17, January 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340631.

Welton:1916:BRV

[768] Thomas A. Welton. On the birthrates in various parts of England and Wales in 1911, 1912, and 1913. Journal of the Royal Statistical Society, 79 (1):18–36, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340632.

Welton:1916:DRV

[769] Thomas A. Welton. On the death-rates in various parts of England and Wales in 1913. Journal of the Royal Statistical Society, 79(1):37–54, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340633.

Welton:1916:OPE

[770] Thomas A. Welton. Occupations of the people, England and Wales, 1911. memorandum as to merchants and clerks. Journal of the Royal Statistical Society, 79(1):55–60, January 1916.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340634.

Forsyth:1916:NUM

[771] C. H. Forsyth. Notes upon Mr. King’s short method of constructing abridged life tables. Journal of the Royal Statistical Society, 79(1):61–62, January 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340635. See

note [772].

King:1916:NUM

[772] George King. Note upon Mr. C. H. Forsyth’s note. Journal of the Royal Statistical Society, 79(1):63– 64, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340636. See [771].

S:1916:BRBa

[773] J. C. S. Book review: The Wealth and Income of the People of the United States, by Willford Isbell King. Journal of the Royal Statistical Society, 79 (1):65–67, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340637.

S:1916:BRBb

[774] J. C. S. Book review: Income: An Examination of the Returns for Services Rendered and from Property Owned in the United States, by Scott Nearing. Journal of the Royal Statistical Society, 79(1):67–68, January 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340638.

S:1916:BRBc

[775] E. C. S. Book review: The Material Culture and Social Institutions of the Simpler Peoples. An Essay in Correlation, by L. T. Hobhouse; G. C. Wheeler; M. Ginsberg. Journal of the Royal Statistical Society, 79(1): 68–69, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340639.

B:1916:BRBa

[776] C. F. B. Book review: Insurance, by T. E. Young. Journal of the Royal Statistical Society, 79(1):70–72, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340640.

B:1916:BRBb

[777] E. B. Book review: Old Age Pensions: Their Actual Working and Ascertained Results in the United Kingdom, by H. J. Hoare; Laurence Gomme. Journal of the Royal Statistical Society, 79(1):72–73, January 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340641.

C:1916:BRBa

[778] C. E. C. Book review: Women in Modern Industry, by B. L. Hutchins; J. J. Mallon. Journal of the Royal Statistical Society, 79(1):73–74, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340642.

P:1916:BRBa

[779] L. L. P. Book review: La Belgique Industrielle et Commerciale de Demain, by Robert Billiard. Journal of the Royal Statistical Society, 79 (1):74–75, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340643.

C:1916:BRBb

[780] W. C. Book review: The Evolution of the English Corn Market from the Twelfth to the Eighteenth Century, by N. S. B. Gras. Journal of the Royal Statistical Society, 79(1):75–77, January 1916. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340644.

W:1916:BRBa

[781] A. D. W. Book review: On the Relation of the Imports to Exports, by J. Taylor Peddie. Journal of the Royal Statistical Society, 79(1):77–78, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340645.

B:1916:BRBc

[782] W. B. Book review: English Public Health Administration, by B. G. Bannington. Journal of the Royal Statistical Society, 79(1):78–79, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340646.

S:1916:BRBd

[783] E. C. S. Book review: The Japanese Problem in the United States. An Investigation for the Commission on Relations with Japan Appointed, by the Federal Council of the Churches of Christ in America, by H. H. Millis. Journal of the Royal Statistical Society, 79(1):79–80, January 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340647.

P:1916:BRBb

[784] L. L. P. Book review: Economic Origins of Jeffersonian Democracy, by Charles Beard. Journal of the Royal Statistical Society, 79(1): 80–81, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340648.

S:1916:BRBe

[785] W. T. S. Book review: English Railways.–Their Development and their Relation to the State, by E. Cleveland-Stevens. Journal of the Royal Statistical Society, 79(1):81– 82, January 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340649.

S:1916:BRBf

[786] W. T. S. Book review: Railway Problems in China, by Mongton Chih Hsu; The Nationalization of Railways in Japan, by Toshiharn Watarai. Journal of the Royal Statistical Society, 79 (1):82–84, January 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340650.

Anonymous:1916:ONPa

[787] Anonymous. Other new publications. Journal of the Royal Statistical Society, 79(1):84–90, January 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340651.

Anonymous:1916:CNa

[788] Anonymous. Current notes. Journal of the Royal Statistical Society, 79 (1):91–104, January 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340652.

Anonymous:1916:SEAa

[789] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 79(1): 105–113, January 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340653.

Anonymous:1916:LALa

[790] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 79(1):114–124, January 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340654.

Anonymous:1916:PRa

[791] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 79 (1):125–132, January 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340655.

deJastrzebski:1916:RBR

[792] T. T. S. de Jastrzebski. The register of Belgian refugees. Journal of the Royal Statistical Society, 79(2): 133–158, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340812.

Darwin:1916:SEN

[793] Leonard Darwin. On the statistical enquiries needed after the War in connection with eugenics. Journal of the Royal Statistical Society, 79 (2):159–188, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340813.

Paish:1916:PC

[794] George Paish. Prices of commodities in 1915. Journal of the Royal Statistical Society, 79(2):189–206, March 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340814.

S:1916:BRBg

[795] C. P. S. Book review: Napier Tercentenary Memorial Volume, by C. G. Knott. Journal of the Royal Statistical Society, 79(2):207–208, March 1916.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340815.

B:1916:BRBd

[796] C. F. B. Book review: Les Bases Théoriques de la Statistique Agricole Internationale, by Umberto Ricci. Journal of the Royal Statistical Society, 79(2):208–210, March 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340816.

F:1916:BRBa

[797] H. S. F. Book review: A History of Currency in the United States, by A. Barton Hepburn. Journal of the Royal Statistical Society, 79(2): 210–214, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340817.

S:1916:BRBh

[798] J. C. S. Book review: Capital To-day — A Study of Recent Economic Development, by Herman Cahn. Journal of the Royal Statistical Society, 79 (2):214–215, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340818.

P:1916:BRBc

[799] L. L. P. Book review: Credit, Industry and the War, by Adam W. Kirkaldy; William Robert Scott. Journal of the Royal Statistical Society, 79 (2):215–217, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340819.

P:1916:BRBd

[800] L. L. P. Book review: Labour in War Time, by G. H. D. Cole. Journal of the Royal Statistical Society, 79 (2):217–219, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340820.

C:1916:BRBc

[801] C. E. C. Book review: Arbitration and Conciliation in Australasia. The Legal Wage in Victoria and New Zealand, by Mary Theresa Rankin; J. Shield Nicholson. Journal of the Royal Statistical Society, 79(2):219–220, March 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340821.

P:1916:BRBe

[802] L. L. P. Book review: English Field Systems, by Howard Levi Gray. Journal of the Royal Statistical Society, 79 (2):220–222, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340822.

C:1916:BRBd

[803] P. G. C. Book review: Our Food Supply–Perils and Remedies, by Christopher Turnor. Journal of the Royal Statistical Society, 79(2): 222–223, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340823.

M:1916:BRBa

[804] H. W. M. Book review: Le Troupeau Français et la Guerre. Viande Indigène-Viande Importée, by Alfred Massé. Journal of the Royal Statistical Society, 79(2):223–225, March 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340824.

C:1916:BRBe

[805] J. R. C. Book review: Monographs on Agricultural Co-operation in Various Countries, by International Institute of Agriculture. Journal of the Royal Statistical Society, 79(2):225–228, March 1916. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340825.

C:1916:BRBf

[806] J. R. C. Book review: The Principles of Rural Credits as Applied in Europe and as Suggested for America, by James B. Morman; John L. Coulter. Journal of the Royal Statistical Society, 79(2):228–231, March 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340826.

P:1916:BRBf

[807] L. L. P. Book review: Les Méthodes Allemandes d’Expansion Économique, by Henri Hauser. Journal of the Royal Statistical Society, 79(2):231– 232, March 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340827.

C:1916:BRBg

[808] G. G. C. Book review: Commerce and Industry, by J. Russell Smith. Journal of the Royal Statistical Society, 79 (2):232–234, March 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340828.

Anonymous:1916:ONPb

[809] Anonymous. Other new publications. Journal of the Royal Statistical Society, 79(2):234–238, March 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340829.

Anonymous:1916:CNb

[810] Anonymous. Current notes. Journal of the Royal Statistical Society, 79 (2):239–244, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340830.

Anonymous:1916:SEAb

[811] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 79 (2):245–250, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340831.

Anonymous:1916:LALb

[812] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 79(2):251–255, March 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340832.

Anonymous:1916:PRb

[813] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 79 (2):256–265, March 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340833.

Paish:1916:WF

[814] George Paish. War finance. Journal of the Royal Statistical Society, 79 (3):267–299, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340539.

Collet:1916:CFA

[815] C. E. Collet. Cost of food for an adult woman. Journal of the Royal Statistical Society, 79(3):300–308, May 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340540.

Dunlop:1916:NES

[816] James Craufurd Dunlop. Note as to error of statement of ages of Young children in a census. Journal of the Royal Statistical Society, 79(3):309–317, May 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340541.

Dunlop:1916:RMC

[817] James Craufurd Dunlop. A rapid method of calculating an average age. Journal of the Royal Statistical Society, 79(3):318–328, May 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340542.

Lehfeldt:1916:NLP

[818] R. A. Lehfeldt. The normal law of progress. Journal of the Royal Statistical Society, 79(3):329–332, May 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340543.

W:1916:BRBb

[819] A. A. W. Book review: Inleiding tot de Beoefening der Statistiek. Deel 3. De Toepassing der Statistische Methode op het Gebied van het Sociaal-Economisch Leven en van de Huishouding der Publiek-Rechtelijke Lichamen. Stuk 1. De Statistiek van het Bedrijfsleven, by C. A. Verrijn Stuart. Journal of the Royal Statistical Society, 79 (3):333–335, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340544.

F:1916:BRBb

[820] A. W. F. Book review: British Incomes and Property, by J. C. Stamp. Journal of the Royal Statistical Society, 79(3):335–336, May 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340545.

P:1916:BRBg

[821] L. L. P. Book review: Economics: An Introduction for the General Reader, by Henry Clay. Journal of the Royal Statistical Society, 79(3):336–338, May 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340546.

S:1916:BRBi

[822] J. C. S. Book review: Why the Dollar is Shrinking. A Study in the High Cost of Living, by Irving Fisher. Journal of the Royal Statistical Society, 79 (3):338–340, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340547.

M:1916:BRBb

[823] H. W. M. Book review: A Study of Industrial Fluctuation: An Enquiry into the Character and Causes of the So-Called Cyclical Movements of Trade, by Dennis Holme Robertson. Journal of the Royal Statistical Society, 79 (3):340–341, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340548.

C:1916:BRBh

[824] E. C. Book review: International Finance, by Hartley Withers. Journal of the Royal Statistical Society, 79 (3):341–343, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340549.

S:1916:BRBj

[825] J. C. S. Book review: Le Budget et le Déficit, by Pierre Baudin. Journal of the Royal Statistical Society, 79 (3):343–346, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340550.

W:1916:BRBc

[826] H. W. Book review: Foreign Exchange and Foreign Bills, in Theory and in Practice, by Wm. F. Spalding. Journal of the Royal Statistical Society, 79 (3):346–347, May 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340551.

W:1916:BRBd

[827] H. W. Book review: Elements of Foreign Exchange: A Foreign Exchange Primer, by Franklin Escher. Journal of the Royal Statistical Society, 79(3):347, May 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340552.

W:1916:BRBe

[828] A. D. W. Book review: International Trade: A Study of the Economic Advantages of Commerce, by Harry Gunnison Brown. Journal of the Royal Statistical Society, 79(3):347–348, May 1916. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340553.

W:1916:BRBf

[829] A. D. W. Book review: The Handicap of British Trade, with Special Regard to East Africa, by W. H. Hooker; Charles E. Musgrave; Notre Commerce d’Exportation Avant, Pendant, et Après la Guerre, by Adolphe Landry. Journal of the Royal Statistical Society, 79(3):348–351, May 1916. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340554.

M:1916:BRBc

[830] H. W. M. Book review: Law and Order in Industry; Five Years’ Experience, by Julius Henry Cohen. Journal of the Royal Statistical Society, 79 (3):351–353, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340555.

M:1916:BRBd

[831] H. W. M. Book review: The Authentic History of the United States Steel Corporation, by Arundel Cotter. Journal of the Royal Statistical Society, 79 (3):353–355, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340556.

P:1916:BRBh

[832] L. L. P. Book review: Agriculture after the War, by A. D. Hall. Journal of the Royal Statistical Society, 79 (3):355–357, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340557.

P:1916:BRBi

[833] L. L. P. Book review: Early Economic Conditions and the Development of Agriculture in Minnesota, by Edward Vandyke Robinson. Journal of the Royal Statistical Society, 79 (3):357–358, May 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340558.

W:1916:BRBg

[834] A. D. W. Book review: Consommation des Boissons Alcooliques dans les Différents Pays, by Johannes Gabrielsson. Journal of the Royal Statistical Society, 79(3):359–360, May 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340559.

H:1916:BRB

[835] E. J. H. Book review: The Exemption of Improvements from Taxation in Canada and the United States, by Robert Murray Haig; Some Probable Effects of the Exemption of Improvements from Taxation in the City of New York by Robert Murray Haig; Excess Condemnation: A Report of the Committee on Taxation of the City of New York, by Herbert S. Swan. Journal of the Royal Statistical Society, 79 (3):360–363, May 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340560.

Anonymous:1916:ONPc

[836] Anonymous. Other new publications. Journal of the Royal Statistical Society, 79(3):363–367, May 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340561.

Anonymous:1916:CNc

[837] Anonymous. Current notes. Journal of the Royal Statistical Society, 79 (3):368–383, May 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340562.

Anonymous:1916:SEAc

[838] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 79 (3):384–391, May 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340563.

Anonymous:1916:LALc

[839] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 79(3):392–397, May 1916.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340564.

Anonymous:1916:VIb

[840] Anonymous. Volume information. Journal of the Royal Statistical Society, 79(4):577–588, July 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340999.

Baines:1916:RTP

[841] J. Athelstane Baines. The recent trend of population in England and Wales. Journal of the Royal Statistical Society, 79(4):399–417, July 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341000.

Nixon:1916:WNV

[842] J. W. Nixon. War and national vital statistical with special reference to the Franco-Prussian War. Journal of the Royal Statistical Society, 79 (4):418–454, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341001.

Edgeworth:1916:MRS

[843] F. Y. Edgeworth. On the mathematical representation of statistical data. Journal of the Royal Statistical Society, 79(4):455–500, July 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341002.

Wood:1916:ICF

[844] Frances Wood. The increase in the cost of food for different classes of society since the outbreak of war. Journal of the Royal Statistical Society, 79 (4):501–508, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341003.

Anonymous:1916:RCF

[845] Anonymous. Report of the Council for the financial year ended December, 31, 1915, and for the sessional year ending June 20, 1916, presented at the eighty-second Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C., on June 20, 1916. Journal of the Royal Statistical Society, 79(4):509–519, July 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341004.

Anonymous:1916:PES

[846] Anonymous. The proceedings of the eighty-second Annual General Meeting, held on Tuesday, June 20, 1916. Journal of the Royal Statistical Society, 79(4):520, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341005.

G:1916:BRBa

[847] M. G. Book review: The Mathematical Theory of Probabilities and Its Application to Frequency Curves and Statistical Methods. Volume I. Mathematical Probabilities and Homograde Statistics, by Arne Fisher; William Bonynge. Journal of the Royal Statistical Society, 79(4):521–523, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341006.

B:1916:BRBe

[848] J. A. B. Book review: The Declining Birth-Rate, Its Causes and Effects, by National Birth-Rate Commission; National Council of Public Morals. Journal of the Royal Statistical Society, 79 (4):523–526, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341007.

B:1916:BRBf

[849] J. A. B. Book review: Population: A Study in Mathusianism, by Warren S. Thompson. Journal of the Royal Statistical Society, 79(4):526–528, July 1916. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341008.

P:1916:BRBj

[850] L. L. P. Book review: Second Thoughts of an Economist, by William Smart; Thomas Jones. Journal of the Royal Statistical Society, 79(4):528–531, July 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341009.

B:1916:BRBg

[851] S. L. B. Book review: Corso di Scienza Della Finanza, by Luigi Einaudi. Journal of the Royal Statistical Society, 79 (4):531–533, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341010.

C:1916:BRBi

[852] E. C. Book review: The Data of Economics: Expressly Designed for the General Reader, by C. J. Melrose. Journal of the Royal Statistical Society, 79 (4):533–535, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341011.

G:1916:BRBb

[853] M. G. Book review: The Mortality from Cancer Throughout the World, by Frederick L. Hoffman. Journal of the Royal Statistical Society, 79 (4):535–536, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341012.

S:1916:BRBk

[854] W. T. S. Book review: The Development of Transportation in Modern England, by W. T. Jackman. Journal of the Royal Statistical Society, 79 (4):536–538, July 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341013.

P:1916:BRBk

[855] L. L. P. Book review: The Marketing of Farm Products, by L. D. H. Weld. Journal of the Royal Statistical Society, 79(4):538–540, July 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341014.

W:1916:BRBh

[856] A. D. W. Book review: Trade as a Science, by Ernest J. P. Benn; Lord Burnham. Journal of the Royal Statistical Society, 79(4):540–541, July 1916. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2341015.

S:1916:BRBl

[857] W. T. S. Book review: The Rise of Rail Power in War and Conquest, 1833–1914, by E. A. Pratt. Journal of the Royal Statistical Society, 79 (4):541–542, July 1916. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341016.

W:1916:BRBi

[858] A. D. W. Book review: The Future of South America, by Roger W. Babson. Journal of the Royal Statistical Society, 79(4):542, July 1916. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341017.

Anonymous:1916:ONPd

[859] Anonymous. Other new publications. Journal of the Royal Statistical Society, 79(4):543–546, July 1916. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341018.

Anonymous:1916:CNd

[860] Anonymous. Current notes. Journal of the Royal Statistical Society, 79 (4):547–559, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341019.

Anonymous:1916:SEAd

[861] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 79 (4):560–565, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341020.

Anonymous:1916:BML

[862] Anonymous. Bi-monthly list of additions to the library. Journal of the Royal Statistical Society, 79(4): 566–569, July 1916. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341021.

Anonymous:1916:ALA

[863] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 79(4):570–576, July 1916. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341022.

Anonymous:1917:VIa

[864] Anonymous. Volume information. Journal of the Royal Statistical Society, 80(1):i–vii, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340692.

Mallet:1917:ORB

[865] Bernard Mallet. The organisation of registration in its bearing on vital statistics. Journal of the Royal Statistical Society, 80(1):1–24, January 1917. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340693.

Anonymous:1917:PN

[866] Anonymous. Proceedings on November 21, 1916. Journal of the Royal Statistical Society, 80(1):25–30, January 1917. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340694.

Drage:1917:ROS

[867] Geoffrey Drage. The reorganisation of official statistics and a central statistical Office. Journal of the Royal Statistical Society, 80(1):31–64, January 1917. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340695.

Edgeworth:1917:MRSa

[868] F. Y. Edgeworth. On the mathematical representation of statistical data. Journal of the Royal Statistical Society, 80(1):65–83, January 1917. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340696.

Traversi:1917:AMPa

[869] A. J. Traversi. Ascertainment of mean population from census returns. Journal of the Royal Statistical Society, 80 (1):84–88, January 1917. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340697.

Anonymous:1917:NA

[870] Anonymous. Note on the above. Journal of the Royal Statistical Society, 80 (1):88–91, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340698.

Anonymous:1917:SRB

[871] Anonymous. Sir Richard Biddulph Martin, Bart. Journal of the Royal Statistical Society, 80(1):92–93, January 1917. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340699.

Anonymous:1917:RHC

[872] Anonymous. Right Hon. Charles Booth, D.C.L., F.R.S. Journal of the Royal Statistical Society, 80(1): 94–98, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340700.

B:1917:BRBa

[873] A. L. B. Book review: Labour, Finance and the War (being the Results of Inquiries, Arranged, by the Section of Economic Science and Statistics of the British Association for the Advancement of Science, during the Years 1915 and 1916), by Adam W. Kirkaldy. Journal of the Royal Statistical Society, 80(1):99–101, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340701.

S:1917:BRBa

[874] J. C. S. Book review: La Richesse de la France Devant la Guerre, by René Pupin. Journal of the Royal Statistical Society, 80(1):101–102, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340702.

S:1917:BRBb

[875] J. C. S. Book review: Finances de Guerre Comparées, by Albin Huart.

Journal of the Royal Statistical Society, 80(1):103–104, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340703.

M:1917:BRBa

[876] H. W. M. Book review: Trust Laws and Unfair Competition, by Joseph E. Davies. Journal of the Royal Statistical Society, 80(1):105–107, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340704.

S:1917:BRBc

[877] C. P. S. Book review: The State as Manufacturer and Trader. An Examination of Government Tobacco Monopolies, by A. W. Madsen. Journal of the Royal Statistical Society, 80(1): 107–108, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340705.

P:1917:BRBa

[878] L. L. P. Book review: The Colonial Tariff Policy of France, by Arthur Gérault. Journal of the Royal Statistical Society, 80(1):108–110, January 1917. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340706.

P:1917:BRBb

[879] L. L. P. Book review: Le Protectionnisme et l’Avenir Économique de la France, by A. Daudé-Bancel; Economic Protectionism, by Josef Grunzel; Eugen von Philippovich. Journal of the Royal Statistical Society, 80(1): 110–112, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340707.

W:1917:BRBa

[880] A. D. W. Book review: Les Zones Franches et l’Exportation Française. Documents et Arguments Réunis par la Chambre de Commerce de Marseille. Journal of the Royal Statistical Society, 80(1):113, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340708.

P:1917:BRBc

[881] L. L. P. Book review: Central Europe, by Friedrich Naumann; W. J. Ashley; Christabel M. Meredith. Journal of the Royal Statistical Society, 80 (1):114–116, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340709.

P:1917:BRBd

[882] L. L. P. Book review: Le Commerce Allemand — Apparences et Réalités, by Daniel Bellet. Journal of the Royal Statistical Society, 80(1):116–117, January 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340710.

P:1917:BRBe

[883] L. L. P. Book review: La Grèce Économique et Financière, by H. Lefeuvre-Méaulle; Paul Deschanel. Journal of the Royal Statistical Society, 80(1):117–118, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340711.

P:1917:BRBf

[884] L. L. P. Book review: Readings in the Economic History of the United States, by Ernest Ludlow Bogart; Charles Manfred Thompson; An Introduction to the Study of Organized Labour in America by George Gorham Groat. Journal of the Royal Statistical Society, 80(1):118–121, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340712.

B:1917:BRBb

[885] E. B. Book review: Poverty and Social Progress, by Maurice Parmelee. Journal of the Royal Statistical Society, 80 (1):121–123, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340713.

G:1917:BRBa

[886] M. G. Book review: Report on the Pension Funds of the City of New York. Part I. Operation of the Nine Existing Pension Funds. Part II. An Actuarial Investigation of the Mortality and Service Experience of the Special and General Service Funds for Municipal Employees. Journal of the Royal Statistical Society, 80(1):123–125, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340714.

deJ:1917:BRB

[887] S. de J. Book review: O Povo Portugez. Aspectos Sociaes e Economicos, by Bento Carqueja. Journal of the Royal Statistical Society, 80(1): 125–128, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340715.

G:1917:BRBb

[888] M. G. Book review: Losses of Life in Modern Wars: Austria-Hungary, France, by Gaston Bodart; Military Selection and Race Deterioration by Vernon Lyman Kellogg; Harald Westergaard. Journal of the Royal Statistical Society, 80(1):128–129, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340716.

G:1917:BRBc

[889] M. G. Book review: Epidemics Resulting from Wars, by Friedrich Prinzing; Harald Westergaard. Journal of the Royal Statistical Society, 80(1): 129–130, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340717.

B:1917:BRBc

[890] J. A. B. Book review: The Economic Life of a Bengal District. A Study, by J. C. Jack. Journal of the Royal Statistical Society, 80(1):130–132, January 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340718.

B:1917:BRBd

[891] J. A. B. Book review: The Population Problem in India, by P. K. Wattal. Journal of the Royal Statistical Society, 80(1):132–135, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340719.

E:1917:BRBa

[892] H. T. E. Book review: Facts About Land. A Reply to “The Land,” the Report of the Unofficial Land Enquiry Committee, by The Land Agents’ Society. Journal of the Royal Statistical Society, 80(1):135–136, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340720.

S:1917:BRBd

[893] W. S. Book review: British Forestry, by E. P. Stebbing. Journal of the Royal Statistical Society, 80(1):136– 138, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340721.

W:1917:BRBb

[894] F. W. Book review: Domestic Service. An Enquiry, by the Women’s Industrial Council by C. V. Butler; Willoughby de Broke. Journal of the Royal Statistical Society, 80(1):138–139, January 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340722.

C:1917:BRBa

[895] E. C. Book review: Memorandum on the Industrial Situation after the War, by Garton Foundation; Eclipse or Empire? by H. B. Gray; Samuel Turner; This is Our Land; Problems of the War and Peace from the Point of View of a Business Man. Journal of the Royal Statistical Society, 80(1): 139–141, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340723.

G:1917:BRBd

[896] M. G. Book review: The Hidden Scourge, by Mary Scharlieb. Journal of the Royal Statistical Society, 80 (1):141, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340724.

Anonymous:1917:ONPa

[897] Anonymous. Other new publications. Journal of the Royal Statistical Society, 80(1):141–144, January 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340725.

Anonymous:1917:CNa

[898] Anonymous. Current notes. Journal of the Royal Statistical Society, 80(1): 145–160, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340726.

Anonymous:1917:SEAa

[899] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 80(1): 161–169, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340727.

Anonymous:1917:LALa

[900] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 80(1):170–179, January 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340728.

Anonymous:1917:PRa

[901] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 80 (1):180–186, January 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340729.

Heyking:1917:ERR

[902] Baron A. Heyking. The economic resources of Russia, with special reference to British opportunities. Journal of the Royal Statistical Society, 80 (2):187–221, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340453.

Brownlee:1917:RIM

[903] John Brownlee. The relation of infantile mortality to mortality in subsequent life. Journal of the Royal Statistical Society, 80(2):222–248, March 1917. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340454.

Trachtenberg:1917:CBD

[904] H. L. Trachtenberg. The correlation between death-rates in successive years of life. Journal of the Royal Statistical Society, 80(2):249–265, March 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340455.

Edgeworth:1917:MRSb

[905] F. Y. Edgeworth. On the mathematical representation of statistical data. Journal of the Royal Statistical Society, 80(2):266–288, March 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340456.

Anonymous:1917:WPC

[906] Anonymous. Wholesale prices of commodities in 1916. Journal of the Royal Statistical Society, 80(2): 289–309, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340457.

B:1917:BRBe

[907] C. F. B. Book review: Criminality and Economic Conditions, by W. A. Bonger; H. O. Horton. Journal of the Royal Statistical Society, 80(2): 310–312, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340458.

M:1917:BRBb

[908] H. W. M. Book review: A Brief History of Panics and their Periodical Occurrence in the United States, by Clement Juglar; De Courcy W. Thom. Journal of the Royal Statistical Society, 80 (2):312–313, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340459.

S:1917:BRBe

[909] J. C. S. Book review: Financial Chapters of the War, by Alexander Dana Noyes. Journal of the Royal Statistical Society, 80(2):313–314, March 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340460.

S:1917:BRBf

[910] J. C. S. Book review: How to Pay for the War: Being Ideas Offered to the Chancellor of the Exchequer, by the Fabian Research Department by Sidney Webb. Journal of the Royal Statistical Society, 80(2):314–316, March 1917. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340461.

C:1917:BRBb

[911] E. C. Book review: The Progress of Capitalism in England, by W. Cunningham. Journal of the Royal Statistical Society, 80(2):316–318, March 1917. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340462.

C:1917:BRBc

[912] C. E. C. Book review: The Homeworker and Her Outlook: A Descriptive Study of Tailoresses and Boxmakers, by V. de Vesselitsky; R. H. Tawney. Journal of the Royal Statistical Society, 80 (2):318–320, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340463.

U:1917:BRB

[913] J. F. U. Book review: The Empire and the Future ; The Problem of the Commonwealth, by L. Curtis; World Commerce in its Relation to the British Empire by Norman R. Byers. Journal of the Royal Statistical Society, 80 (2):320–322, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340464.

S:1917:BRBg

[914] J. C. S. Book review: The Belief in God and Immortality: A Psychological, Anthropological and Statistical Study, by James H. Leuba. Journal of the Royal Statistical Society, 80 (2):322–323, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340465.

M:1917:BRBc

[915] H. W. M. Book review: The Administration of Industrial Enterprises, with Special Reference to Factory Practice, by Edward D. Jones. Journal of the Royal Statistical Society, 80(2): 323–324, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340466.

Anonymous:1917:ONPb

[916] Anonymous. Other new publications. Journal of the Royal Statistical Society, 80(2):325–331, March 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340467.

Anonymous:1917:CNb

[917] Anonymous. Current notes. Journal of the Royal Statistical Society, 80 (2):332–338, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340468.

Anonymous:1917:SEAb

[918] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 80 (2):339–345, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340469.

Anonymous:1917:LALb

[919] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 80(2):346–349, March 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340470.

Anonymous:1917:PRb

[920] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 80 (2):350–359, March 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340471.

ofDunraven:1917:HIO

[921] The Right Hon. The Earl of Dunraven, K. P., and C. M. G. How to improve our fishing industries. Journal of the Royal Statistical Society, 80(3):361–382, May 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341023.

Mantoux:1917:TFB

[922] Paul Mantoux. Trade with France before and after the War. Journal of the Royal Statistical Society, 80 (3):383–410, May 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341024.

Edgeworth:1917:MRSc

[923] F. Y. Edgeworth. On the mathematical representation of statistical data. Journal of the Royal Statistical Society, 80(3):411–437, May 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341025.

M:1917:BRBd

[924] H. W. M. Book review: L’Alimentation de la France et les Ressources Coloniales ou Etrangères, by Daniel Bellet. Journal of the Royal Statistical Society, 80(3):438–440, May 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341026.

W:1917:BRBc

[925] A. D. W. Book review: Principles of Commerce: A Study of the Mechanism, the Advantages, and the Transportation Costs of Foreign and Domestic Trade, by Harry Gunnison Brown. Journal of the Royal Statistical Society, 80(3):440–441, May 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341027.

W:1917:BRBd

[926] A. D. W. Book review: American and Foreign Investment Bonds, by William L. Raymond. Journal of the Royal Statistical Society, 80(3):441–442, May 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341028.

B:1917:BRBf

[927] E. B. Book review: Addresses and Papers on Insurance, by Rufus M. Potts; Social Insurance: An Economic Analysis by Robert Morse Woodbury. Journal of the Royal Statistical Society, 80 (3):442–444, May 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341029.

S:1917:BRBh

[928] J. C. S. Book review: The Industrial Outlook, by H. Sanderson Furniss. Journal of the Royal Statistical Society, 80(3):444–446, May 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341030.

Anonymous:1917:ONPc

[929] Anonymous. Other new publications. Journal of the Royal Statistical Society, 80(3):446–450, May 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341031.

Anonymous:1917:CNc

[930] Anonymous. Current notes. Journal of the Royal Statistical Society, 80 (3):451–456, May 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341032.

Anonymous:1917:SEAc

[931] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 80 (3):457–462, May 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341033.

Anonymous:1917:LALc

[932] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 80(3):463–466, May 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341034.

Anonymous:1917:VIb

[933] Anonymous. Volume information. Journal of the Royal Statistical Society, 80(4):596–605, July 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341035.

Nicholson:1917:SAI

[934] J. Shield Nicholson. Statistical aspects of inflation. Journal of the Royal Statistical Society, 80(4):467–520, July 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341036.

Welton:1917:NBD

[935] Thomas A. Welton. Note on the birthand death-rates in various parts of England and Wales in 1915. Journal of the Royal Statistical Society, 80 (4):521–528, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341037.

Traversi:1917:AMPb

[936] A. T. Traversi. Ascertainment of mean population from census returns. Journal of the Royal Statistical Society, 80 (4):529–532, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341038.

Anonymous:1917:RCF

[937] Anonymous. Report of the Council for the financial year ended December 31, 1916, and for the sessional year ending June 19, 1917, presented at the eighty-third Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C. 2, on June 19, 1917. Journal of the Royal Statistical Society, 80(4):533–543, July 1917. ISSN 0952-8385. URL http://www.

jstor.org/stable/2341039.

Anonymous:1917:PET

[938] Anonymous. Proceedings at the eightythird Annual General Meeting of the Royal Statistical Society held on Tuesday, June 19, 1917. Journal of the Royal Statistical Society, 80(4): 544–545, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341040.

E:1917:BRBb

[939] F. Y. E. Book review: Scope and Method of Statistics, by Harald Westergaard; Walter P. Willcox. Journal of the Royal Statistical Society, 80 (4):546–551, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341041.

P:1917:BRBg

[940] L. L. P. Book review: The Town Labourer, 1760–1832, by J. L. Hammond; Barbara Hammond. Journal of the Royal Statistical Society, 80 (4):551–554, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341042.

S:1917:BRBi

[941] J. C. S. Book review: National Economy — An Outline of Public Administration, by Henry Higgs. Journal of the Royal Statistical Society, 80(4):554, July 1917. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2341043.

G:1917:BRBe

[942] M. G. Book review: Health and the State, by William A. Brend. Journal of the Royal Statistical Society, 80 (4):555–557, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341044.

C:1917:BRBd

[943] C. E. C. Book review: Profit-Sharing in the United States, by Boris Emmet. Journal of the Royal Statistical Society, 80(4):557–558, July 1917. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341045.

P:1917:BRBh

[944] L. L. P. Book review: Economic Conditions, 1815 and 1914, by H. R. Hodges. Journal of the Royal Statistical Society, 80(4):558–559, July 1917. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2341046.

P:1917:BRBi

[945] L. L. P. Book review: Economic Problems of Peace After War, by William Robert Scott. Journal of the Royal Statistical Society, 80(4):559–560, July 1917. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341047.

C:1917:BRBe

[946] E. C. Book review: After-War Problems, by Earl of Cromer; Viscount Haldane of Exeter; Alfred Marshall; W. H. Dawson; The War and the Nation: A Study in Constructive Politics, by William Cecil Dampier Whetham. Journal of the Royal Statistical Society, 80(4):560–563, July 1917. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2341048.

P:1917:BRBj

[947] L. L. P. Book review: L’Union Commerciale des Alliés Après la Guerre, by Bernard Lavergne. Journal of the Royal Statistical Society, 80(4):563–564, July 1917. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2341049.

S:1917:BRBj

[948] J. C. S. Book review: L’Effort Économique de la France Pendant Deux ans et Demi de Guerre. Journal of the Royal Statistical Society, 80(4):564, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341050.

Anonymous:1917:ONPd

[949] Anonymous. Other new publications. Journal of the Royal Statistical Society, 80(4):565–568, July 1917. ISSN

0952-8385. URL http://www.jstor.
org/stable/2341051.

Anonymous:1917:CNd

[950] Anonymous. Current notes. Journal of the Royal Statistical Society, 80 (4):569–579, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341052.

Anonymous:1917:SEAd

[951] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 80 (4):580–584, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341053.

Anonymous:1917:BML

[952] Anonymous. Bi-monthly list of additions to the library. Journal of the Royal Statistical Society, 80(4): 585–588, July 1917. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2341054.

Anonymous:1917:ALA

[953] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 80(4):589–595, July 1917. ISSN 0952-8385. URL http://

www.jstor.org/stable/2341055.

Anonymous:1918:VIa

[954] Anonymous. Volume information. Journal of the Royal Statistical Society, 81(1):i–vii, January 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340565.

Mallet:1918:VSA

[955] Bernard Mallet. Vital statistics as affected by the War. Journal of the Royal Statistical Society, 81(1):1–36, January 1918. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340566.

Anonymous:1918:PN

[956] Anonymous. Proceedings on November 20, 1917. Journal of the Royal Statistical Society, 81(1):37–40, January 1918. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340567.

Rew:1918:PWF

[957] R. Henry Rew. The prospects of the world’s food supplies after the War. Journal of the Royal Statistical Society, 81(1):41–74, January 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340568.

Isserlis:1918:VMC

[958] L. Isserlis. On the value of a mean as calculated from a sample. Journal of the Royal Statistical Society, 81 (1):75–81, January 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/

stable/2340569. See reply [978].

Lehfeldt:1918:LT

[959] R. A. Lehfeldt. On life-tables. Journal of the Royal Statistical Society, 81 (1):82–95, January 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340570.

Y:1918:BRB

[960] G. U. Y. Book review: Appendix A, Vol I. Census of the Commonwealth of Australia. The Mathematical Theory of Population, of its Character and Fluctuations, and of the Factors which Influence them, by G. H. Knibbs. Journal of the Royal Statistical Society, 81 (1):96–97, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340571.

B:1918:BRB

[961] J. A. B. Book review: Birth-Rate and Empire, by James Marchant. Journal of the Royal Statistical Society, 81 (1):97–99, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340572.

P:1918:BRBa

[962] L. L. P. Book review: Commerce and the Empire: 1914 and After, by Edward Pulsford. Journal of the Royal Statistical Society, 81(1): 99–101, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340573.

C:1918:BRBa

[963] E. C. Book review: Our Money and the State, by Hartley Withers. Journal of the Royal Statistical Society, 81 (1):101–103, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340574.

S:1918:BRBa

[964] J. C. S. Book review: Les Répercussions Économiques de la Guerre Actuelle sur la France, by Georges Renard. Journal of the Royal Statistical Society, 81(1):103– 104, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340575.

P:1918:BRBb

[965] L. L. P. Book review: The Wheat Problem, by William Crookes. Journal of the Royal Statistical Society, 81 (1):104–105, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340576.

P:1918:BRBc

[966] L. L. P. Book review: The Book of the Rothamsted Experiments, by A. D. Hall; E. J. Russell. Journal of the Royal Statistical Society, 81(1): 105–107, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340577.

H:1918:BRB

[967] E. J. H. Book review: Report of the Royal Commission on the Housing of the Industrial Population of Scotland, Rural and Urban. Journal of the Royal Statistical Society, 81(1): 107–109, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340578.

G:1918:BRBa

[968] M. G. Book review: A Monograph on the Epidemic of Poliomyelitis (Infantile Paralysis) in New York City in 1916. Journal of the Royal Statistical Society, 81(1):109–110, January 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340579.

W:1918:BRBa

[969] A. D. W. Book review: La Question de L’Alcool: Allégations et Réalités, by Yves-Guyot. Journal of the Royal Statistical Society, 81(1):110–111, January 1918. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340580.

M:1918:BRBa

[970] H. W. M. Book review: Business Statistics, by Melvin T. Copeland. Journal of the Royal Statistical Society, 81(1):111–112, January 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340581.

Anonymous:1918:ONPa

[971] Anonymous. Other new publications. Journal of the Royal Statistical Society, 81(1):112–119, January 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340582.

Anonymous:1918:CNa

[972] Anonymous. Current notes. Journal of the Royal Statistical Society, 81(1): 120–135, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340583.

Anonymous:1918:SEAa

[973] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 81(1): 136–144, January 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340584.

Anonymous:1918:LALa

[974] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 81(1):145–154, January 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340585.

Anonymous:1918:PRa

[975] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 81 (1):155–160, January 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340586.

Spensley:1918:UHP

[976] J. Calvert Spensley. Urban housing problems. Journal of the Royal Statistical Society, 81(2):161–228, March 1918. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340981.

Drage:1918:PWS

[977] Geoffrey Drage. Pre-war statistics of Poland and Lithuania. Journal of the Royal Statistical Society, 81(2): 229–321, March 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340982.

Edgeworth:1918:MRS

[978] F. Y. Edgeworth. Mathematical representation of statistics: A reply. Journal of the Royal Statistical Society, 81 (2):322–333, March 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340983. See [958].

Anonymous:1918:WPC

[979] Anonymous. Wholesale prices of commodities in 1917. Journal of the Royal Statistical Society, 81(2): 334–349, March 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340984.

W:1918:BRBb

[980] A. A. W. Book review: Inleiding tot de Beoefening der Statistiek. Deel 3. De Toepasing der Statistische Methode op het Sociaal-Economisch Leven en van de Huishouding der Publiek-Rechtelijke Lichamen. Stuk 2. De Statistiek van Volksrijksdom, by Volksinkomen en der Publiek-Rechtelijke Lichamen; C. A. Verrijn Stuart; Erven F. Bohn. Journal of the Royal Statistical Society, 81 (2):350–352, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340985.

G:1918:BRBb

[981] M. G. Book review: The Combination of Observations, by David Brunt. Journal of the Royal Statistical Society, 81 (2):352–353, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340986.

W:1918:BRBc

[982] H. W. Book review: War Finance, by J. Shield Nicholson. Journal of the Royal Statistical Society, 81(2): 353–354, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340987.

S:1918:BRBb

[983] J. C. S. Book review: The System of Financial Administration of Great Britain, by William F. Willoughby; Westel W. Willoughby; Samuel Mc-Cune Lindsay; A. Lawrence Lowell. Journal of the Royal Statistical Society, 81(2):354–356, March 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340988.

S:1918:BRBc

[984] J. C. S. Book review: The Development of Rates of Postage: An Historical and Analytical Study, by A. D. Smith. Journal of the Royal Statistical Society, 81 (2):356–358, March 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340989.

P:1918:BRBd

[985] L. L. P. Book review: Three Centuries of Prices of Wheat, Flour and Bread, by John Kirkland. Journal of the Royal Statistical Society, 81(2): 358–359, March 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340990.

P:1918:BRBe

[986] L. L. P. Book review: The Determination of Farming Costs, by C. S. Orwin. Journal of the Royal Statistical Society, 81(2):359–360, March 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340991.

C:1918:BRBb

[987] J. R. C. Book review: Rural Reconstruction in Ireland: A Record of Co-operative Organisation, by Lionel Smith-Gordon; Laurence C. Staples. Journal of the Royal Statistical Society, 81(2):360–363, March 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340992.

G:1918:BRBc

[988] M. G. Book review: The Cancer Problem. A Statistical Study, by C. E. Green. Journal of the Royal Statistical Society, 81(2):363, March 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340993.

Anonymous:1918:ONPb

[989] Anonymous. Other new publications. Journal of the Royal Statistical Society, 81(2):363–366, March 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340994.

Anonymous:1918:CNb

[990] Anonymous. Current notes. Journal of the Royal Statistical Society, 81 (2):367–378, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340995.

Anonymous:1918:SEAb

[991] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 81 (2):379–382, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340996.

Anonymous:1918:LALb

[992] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 81(2):383–386, March 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340997.

Anonymous:1918:PRb

[993] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 81 (2):387–396, March 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340998.

Harper:1918:BLT

[994] Edgar J. Harper. The bases of local taxation in England. Journal of the Royal Statistical Society, 81 (3):397–448, May 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340857.

Attolico:1918:IPI

[995] B. Attolico and F. Giannini. The industrial position of Italy. Journal of the Royal Statistical Society, 81 (3):449–498, May 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340858.

Westergaard:1918:FS

[996] Harald Westergaard. On the future of statistics. Journal of the Royal Statistical Society, 81(3):499–520, May 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340859.

S:1918:BRBd

[997] E. C. S. Book review: An Introduction to Statistical Methods, by Horace Secrist. Journal of the Royal Statistical Society, 81(3):521–522, May 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340860.

A:1918:BRB

[998] P. A. Book review: Economic Annals of the Nineteenth Century, 1821–30, by William Smart. Journal of the Royal Statistical Society, 81(3):522–524, May 1918. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340861.

L:1918:BRB

[999] A. L. Book review: Industry and Finance, by Adam W. Kirkaldy. Journal of the Royal Statistical Society, 81 (3):524–526, May 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340862.

P:1918:BRBf

[1000] L. L. P. Book review: Finance and Trade Under Edward III, by George Unwin. Journal of the Royal Statistical Society, 81(3):526–529, May 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340863.

C:1918:BRBc

[1001] J. R. C. Book review: The Future of Our Agriculture, by Henry W. Wolff. Journal of the Royal Statistical Society, 81(3):529–533, May 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340864.

S:1918:BRBe

[1002] J. C. S. Book review: The State Tax Commission. — A Study of the Development and Results of State Control Over the Assessment of Property for Taxation, by Harley Leist Lutz. Journal of the Royal Statistical Society, 81 (3):533–535, May 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340865.

P:1918:BRBg

[1003] L. L. P. Book review: Our Daily Bread, by George Radford. Journal of the Royal Statistical Society, 81 (3):535–537, May 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340866.

C:1918:BRBd

[1004] C. E. C. Book review: Economic Effects of the War Upon Women and Children in Great Britain, by Irene Osgood Andrews; Margaretta Hobbs. Journal of the Royal Statistical Society, 81(3):537–538, May 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340867.

S:1918:BRBf

[1005] J. C. S. Book review: Welfare and Housing, by J. E. Hutton. Journal of the Royal Statistical Society, 81(3):538, May 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340868.

M:1918:BRBb

[1006] H. W. M. Book review: Household Manufactures in the United States, 1640–1860: A Study in Industrial History, by Rolla Milton Tryon. Journal of the Royal Statistical Society, 81 (3):539–540, May 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340869.

Anonymous:1918:ONPc

[1007] Anonymous. Other new publications. Journal of the Royal Statistical Society, 81(3):540–542, May 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340870.

Anonymous:1918:CNc

[1008] Anonymous. Current notes. Journal of the Royal Statistical Society, 81 (3):543–553, May 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340871.

Anonymous:1918:SEAc

[1009] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 81 (3):554–558, May 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340872.

Anonymous:1918:LALc

[1010] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 81(3):559–562, May 1918. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340873.

Anonymous:1918:VIb

[1011] Anonymous. Volume information.

Journal of the Royal Statistical Society, 81(4):692–701, July 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340656.

Stamp:1918:ETF

[1012] J. C. Stamp. The effect of trade fluctuations upon profits. Journal of the Royal Statistical Society, 81 (4):563–608, July 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340657.

Yamasaki:1918:RED

[1013] K. Yamasaki. Recent economic developments in Japan in their relation to her trade with the United Kingdom. Journal of the Royal Statistical Society, 81(4):609–623, July 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340658.

Edgeworth:1918:VMC

[1014] F. Y. Edgeworth. On the value of a mean as calculated from a sample. Journal of the Royal Statistical Society, 81(4):624–632, July 1918. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340659.

Anonymous:1918:RCF

[1015] Anonymous. Report of the Council for the financial year ended December 31, 1917, and for sessional year ending June 18, 1918, presented at the eighty-fourth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C. 2, on June 18, 1918. Journal of the Royal Statistical Society, 81(4):633–643, July 1918.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340660.

Anonymous:1918:PEF

[1016] Anonymous. Proceedings at the eightyfourth Annual General Meeting of the Royal Statistical Society, held on Tuesday, June 18, 1918. Journal of the Royal Statistical Society, 81 (4):644–645, July 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340661.

S:1918:BRBg

[1017] E. C. S. Book review: The Private Wealth of Australia, and its Growth as Ascertained, by Various Methods, Together with a Report of the War Census of 1915 by Commonwealth Statistician, Commonwealth Bureau of Census and Statistician. Journal of the Royal Statistical Society, 81(4):646–648, July 1918. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340662.

P:1918:BRBh

[1018] L. L. P. Book review: Early Economic Effects of the European War Upon Canada, by Adam Shortt; Early Effects of the European War Upon the Finance, Commerce and Industry of Chile by L. S. Rowe; Being Preliminary Economic Studies of the War by David Kinley. Journal of the Royal Statistical Society, 81(4):649–652, July 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340663.

P:1918:BRBi

[1019] L. L. P. Book review: La Vitalité Économique de la France Avant et Après la Guerre, by F. Sauvaire-Jourdan. Journal of the Royal Statistical Society, 81(4):652–655, July 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340664.

C:1918:BRBe

[1020] J. R. C. Book review: Co-operation in Danish Agriculture, by Harald Faber; Andelsbevoegelsen i Danmark, by H. Hertel; E. J. Russell. Journal of the Royal Statistical Society, 81(4): 655–657, July 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340665.

M:1918:BRBc

[1021] H. W. M. Book review: Factory Administration and Accounts: A Book of Reference with Tables and Specimen Forms for Managers, Engineers, and Accountants, by Edward T. Elbourne; Andrew Homemorton; John Maughfling. Journal of the Royal Statistical Society, 81(4):658–659, July 1918. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340666.

S:1918:BRBh

[1022] J. C. S. Book review: Capital, War and Wages, by W. H. Mallock; A Levy on Capital, by F. W. Pethick Lawrence. Journal of the Royal Statistical Society, 81(4):659–660, July 1918. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340667.

G:1918:BRBd

[1023] M. G. Book review: Muscular Work, a Metabolic Study with Special Reference to the Efficiency of the Human Body as a Machine, by Francis G. Benedict; Edward P. Cathcart. Journal of the Royal Statistical Society, 81(4):660–661, July 1918. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340668.

Anonymous:1918:ONPd

[1024] Anonymous. Other new publications. Journal of the Royal Statistical Society, 81(4):661–664, July 1918. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340669.

Anonymous:1918:CNd

[1025] Anonymous. Current notes. Journal of the Royal Statistical Society, 81 (4):665–676, July 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340670.

Anonymous:1918:SEAd

[1026] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 81 (4):677–680, July 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340671.

Anonymous:1918:BML

[1027] Anonymous. Bi-monthly list of additions to the library. Journal of the Royal Statistical Society, 81(4): 681–684, July 1918. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340672.

Anonymous:1918:ALA

[1028] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 81(4):685–691, July 1918. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340673.

Anonymous:1919:VIa

[1029] Anonymous. Volume information. Journal of the Royal Statistical Society, 82(1):i–vii, January 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340730.

Darwin:1919:ERE

[1030] Leonard Darwin. Eugenics in relation to economics and statistics. Journal of the Royal Statistical Society, 82 (1):1–33, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340731.

Brownlee:1919:NBL

[1031] John Brownlee. Notes on the biology of a life-table. Journal of the Royal Statistical Society, 82(1):34–77, January 1919. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340732.

Greenwood:1919:NGE

[1032] Major Greenwood and Cecily M.

Thompson. A note on German and English war-time diets. Journal of the Royal Statistical Society, 82(1): 78–80, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340733.

P:1919:BRBa

[1033] L. L. P. Book review: The Economic History of Ireland in the Eighteenth Century, by George O’Brien. Journal of the Royal Statistical Society, 82 (1):81–84, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340734.

S:1919:BRBa

[1034] J. C. S. Book review: Economic Problems of Peace After War, by William Robert Scott; Adam Smith. Journal of the Royal Statistical Society, 82 (1):84–86, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340735.

M:1919:BRBa

[1035] H. W. M. Book review: Labour and Capital After the War, by S. J. Chapman. Journal of the Royal Statistical Society, 82(1):86–87, January 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340736.

W:1919:BRBa

[1036] G. H. W. Book review: Wool, by Frank Ormerod; Staple Trades and Industries by Gordon D. Knox. Journal of the Royal Statistical Society, 82 (1):87–90, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340737.

G:1919:BRBa

[1037] M. G. Book review: Use of Factory Statistics in the Investigation of Industrial Fatigue — A Manual for Field Research, by Philip Sargant Florence. Journal of the Royal Statistical Society, 82(1):90–91, January 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340738.

B:1919:BRBa

[1038] E. B. Book review: Postal Savings: An Historical and Critical Study of the Postal Savings Bank System in the United States, by Edwin W. Kemmerer. Journal of the Royal Statistical Society, 82(1):91–92, January 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340739.

B:1919:BRBb

[1039] E. B. Book review: Proceedings of the Conference on Social Insurance Called, by the International Association of Industrial Accident Boards and Commissions, Washington, D.C., December 5 to 9, 1916. Journal of the Royal Statistical Society, 82(1):92–93, January 1919. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340740.

P:1919:BRBb

[1040] L. L. P. Book review: The Iron Circle, by S. Herzog; M. L. Turrentine. Journal of the Royal Statistical Society, 82 (1):93–96, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340741.

A:1919:BRBa

[1041] P. A. Book review: Russia: Its Trade and Commerce, by Arthur Raffalovich; Problems Confronting Russia by Baron A. Heyking. Journal of the Royal Statistical Society, 82(1):96–98, January 1919. ISSN 0952-8385. URL http:/

/www.jstor.org/stable/2340742.

Anonymous:1919:ONPa

[1042] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 82(1):98–104, January 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340743.

Anonymous:1919:CNa

[1043] Anonymous. Current notes. Journal of the Royal Statistical Society, 82(1): 105–118, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340744.

Anonymous:1919:SEAa

[1044] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 82(1): 119–126, January 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340745.

Anonymous:1919:LALa

[1045] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 82(1):127–135, January 1919. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340746.

Anonymous:1919:PRa

[1046] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 82 (1):136–141, January 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340747.

Samuel:1919:TVC

[1047] Herbert Samuel. The taxation of the various classes of the people. Journal of the Royal Statistical Society, 82 (2):143–182, March 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340674.

Anonymous:1919:PJ

[1048] Anonymous. Proceedings on January 21, 1919. Journal of the Royal Statistical Society, 82(2):182–185, March 1919. ISSN 0952-8385. URL http:/

                                        /www.jstor.org/stable/2340675.

Greenwood:1919:PIO

[1049] Major Greenwood. Problems of industrial organisation. Journal of the Royal Statistical Society, 82(2): 186–221, March 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340676.

Edgeworth:1919:PRS

[1050] F. Y. Edgeworth. Psychical research and statistical method. Journal of the Royal Statistical Society, 82(2):

222–228, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340677.

P:1919:BRBc

[1051] L. L. P. Book review: Labour and Industry in Australia, by T. A. Coghlan. Journal of the Royal Statistical Society, 82(2):229–234, March 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340678.

S:1919:BRBb

[1052] J. C. S. Book review: Parliament and the Taxpayer, by E. H. Davenport. Journal of the Royal Statistical Society, 82(2):234–236, March 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340679.

B:1919:BRBc

[1053] E. B. Book review: Public Administration in the Highlands and Islands of Scotland, by John Percival Day. Journal of the Royal Statistical Society, 82 (2):236–238, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340680.

S:1919:BRBc

[1054] J. C. S. Book review: Money: Its Connection with Rising and Falling Prices, by Edwin Cannan. Journal of the Royal Statistical Society, 82(2): 238–239, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340681.

P:1919:BRBd

[1055] L. L. P. Book review: Labour in the Commonwealth, by G. D. H. Cole; An Introduction to Trade Unionism, by G. D. H. Cole; The Payment of Wages, by G. D. H. Cole. Journal of the Royal Statistical Society, 82(2): 239–245, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340682.

B:1919:BRBd

[1056] A. L. B. Book review: The Standard of Living in Japan, by Kokichi Morimoto. Journal of the Royal Statistical Society, 82(2):245–246, March 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340683.

F:1919:BRBa

[1057] A. F. Book review: Cotton, by George Bigwood; Gordon D. Knox. Journal of the Royal Statistical Society, 82 (2):247–249, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340684.

P:1919:BRBe

[1058] L. L. P. Book review: The New Tariffism, by J. M. Robertson. Journal of the Royal Statistical Society, 82 (2):249–250, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340685.

P:1919:BRBf

[1059] L. L. P. Book review: Can we Compete?, by G. E. Mappin. Journal of the Royal Statistical Society, 82(2): 250–252, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340686.

Anonymous:1919:ONPb

[1060] Anonymous. Other new publications. Journal of the Royal Statistical Society, 82(2):252–255, March 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340687.

Anonymous:1919:CNb

[1061] Anonymous. Current notes. Journal of the Royal Statistical Society, 82 (2):256–271, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340688.

Anonymous:1919:SEAb

[1062] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 82 (2):272–276, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340689.

Anonymous:1919:LALb

[1063] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 82(2):277–282, March 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340690.

Anonymous:1919:PRb

[1064] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 82 (2):283–291, March 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340691.

Diouritch:1919:SDS

[1065] G. Diouritch. A survey of the development of the Serbian (Southern Slav) nation: An economic and statistical study. Journal of the Royal Statistical Society, 82(3):293–342, May 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340776.

Bowley:1919:MCC

[1066] A. L. Bowley. The measurement of changes in the cost of living. Journal of the Royal Statistical Society, 82 (3):343–372, May 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340777.

Moore:1919:CCU

[1067] Henry L. Moore. Crop-cycles in the United Kingdom and in the United States. Journal of the Royal Statistical Society, 82(3):373–389, May 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340778.

P:1919:BRBg

[1068] L. L. P. Book review: Cours d’Économie Politique. Livre Premier. Théorie Générale des Phénomenès Économiques, by C. Colson. Journal of the Royal Statistical Society, 82 (3):390–395, May 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340779.

P:1919:BRBh

[1069] L. L. P. Book review: Cours d’Économie Politique. Livre Deuxième: Le Travail et les Questions Ouvrières, by C. Colson. Journal of the Royal Statistical Society, 82(3):395–398, May 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340780.

P:1919:BRBi

[1070] L. L. P. Book review: The Economic Foundations of Peace, by J. L. Garvin. Journal of the Royal Statistical Society, 82(3):398–402, May 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340781.

S:1919:BRBd

[1071] J. C. S. Book review: The War Debt and How to Meet It, with an Examination of the Proposed “Capital Levy.”, by J. E. Allen. Journal of the Royal Statistical Society, 82(3):403–404, May 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340782.

W:1919:BRBb

[1072] A. D. W. Book review: Gold, Prices, and the Witwatersrand, by R. A. Lehfeldt. Journal of the Royal Statistical Society, 82(3):404–406, May 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340783.

M:1919:BRBb

[1073] H. W. M. Book review: The Trust Problem, by J. W. Jenks. Journal of the Royal Statistical Society, 82 (3):406–408, May 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340784.

B:1919:BRBe

[1074] E. B. Book review: Effects of the War upon Insurance: With Special Reference to the Substitution of Insurance for Pensions, by William F. Gelphart. Journal of the Royal Statistical Society, 82(3):408–409, May 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340785.

H:1919:BRB

[1075] B. L. H. Book review: Australian Social Development, by Faculty of Political Science of Columbia University. Journal of the Royal Statistical Society, 82(3):409–410, May 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340786.

F:1919:BRBb

[1076] A. F. Book review: The Silk Industry and Trade, by Ratan C. Rawlley.

Journal of the Royal Statistical Society, 82(3):410–412, May 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340787.

J:1919:BRBa

[1077] W. H. J. Book review: Railway Re-Organization, by A Railway Officer. Journal of the Royal Statistical Society, 82(3):412–413, May 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340788.

P:1919:BRBj

[1078] L. L. P. Book review: Life of Viscount Rhondda, by J. Vyrnwy Morgan. Journal of the Royal Statistical Society, 82(3):413–414, May 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340789.

B:1919:BRBf

[1079] J. A. B. Book review: The Industrial Decline in India, by Balkrishna. Journal of the Royal Statistical Society, 82 (3):414–415, May 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340790.

Anonymous:1919:ONPc

[1080] Anonymous. Other new publications. Journal of the Royal Statistical Society, 82(3):415–420, May 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340791.

Anonymous:1919:CNc

[1081] Anonymous. Current notes. Journal of the Royal Statistical Society, 82 (3):421–431, May 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340792.

Anonymous:1919:SEAc

[1082] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 82 (3):432–434, May 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340793.

Anonymous:1919:LALc

[1083] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 82(3):435–439, May 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340794.

Anonymous:1919:VIb

[1084] Anonymous. Volume information.

Journal of the Royal Statistical Society, 82(4):634–642, July 1919. ISSN

0952-8385. URL http://www.jstor.
org/stable/2340472.

Stamp:1919:WIC

[1085] J. C. Stamp. The wealth and income of the Chief powers. Journal of the Royal Statistical Society, 82(4):441–507, July 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340473.

Barton:1919:CWW

[1086] Dorothea M. Barton. The course of Women’s wages. Journal of the Royal Statistical Society, 82(4):508–553, July 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340474.

Anonymous:1919:RCF

[1087] Anonymous. Report of the Council for the financial year ended December 31, 1918, and for the sessional year ending June 17, 1919, presented at the eighty-fifth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, W.C. 2, on June 17, 1919. Journal of the Royal Statistical Society, 82(4):554–564, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340475.

Mallet:1919:PEF

[1088] Bernard Mallet. Proceedings at the eighty-fifth Annual General Meeting of the Royal Statistical Society held on Tuesday, June 17, 1919. Journal of the Royal Statistical Society, 82 (4):565–567, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340476.

S:1919:BRBe

[1089] J. C. S. Book review: The Taxation of Capital, by Alfred W. Soward; W. E. Willan. Journal of the Royal Statistical Society, 82(4):568–571, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340477.

S:1919:BRBf

[1090] J. C. S. Book review: Taxation in the New State, by J. A. Hobson. Journal of the Royal Statistical Society, 82 (4):571–575, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340478.

S:1919:BRBg

[1091] J. C. S. Book review: The Problem of a National Budget, by W. F. Willoughby. Journal of the Royal Statistical Society, 82(4):575–576, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340479.

A:1919:BRBb

[1092] P. A. Book review: Papers on Current Finance, by H. S. Foxwell. Journal of the Royal Statistical Society, 82 (4):576–579, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340480.

B:1919:BRBg

[1093] S. L. B. Book review: Principii di Economia Commerciale, by Gino Arias. Journal of the Royal Statistical Society, 82(4):579–580, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340481.

G:1919:BRBb

[1094] M. G. Book review: Introduction to Mathematical Statistics, by Carl J. West. Journal of the Royal Statistical Society, 82(4):580–581, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340482.

B:1919:BRBh

[1095] J. A. B. Book review: Indian Industrial Commission, 1916–18 Report. Journal of the Royal Statistical Society, 82 (4):581–584, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340483.

T:1919:BRBa

[1096] C. T. Book review: I Fattori Demografici del Conflitto Europeo, by Lanfranco Maroi. Journal of the Royal Statistical Society, 82(4):584–587, July 1919. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340484.

A:1919:BRBc

[1097] P. A. Book review: Russia: A Study Social and Industrial, by A. N. Drew; Commercial Russia by William Henry Beable. Journal of the Royal Statistical Society, 82(4):588–589, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340485.

W:1919:BRBc

[1098] A. D. W. Book review: The Control of the Drink Trade in Britain. A Contribution to National Efficiency During the Great War, 1915–18, by Henry Carter; Lord D’Abernon. Journal of the Royal Statistical Society, 82 (4):590–591, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340486.

B:1919:BRBi

[1099] E. B. Book review: Co-Operation for Farmers, by Lionel Smith-Gordon. Journal of the Royal Statistical Society, 82(4):591–592, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340487.

B:1919:BRBj

[1100] E. B. Book review: Unemployment and American Trade Unions, by D. P. Smelser. Journal of the Royal Statistical Society, 82(4):592–594, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340488.

D:1919:BRB

[1101] J. M. D. Book review: A Monograph on Indian Railway Rates, by S. C. Ghose. Journal of the Royal Statistical Society, 82(4):594–596, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340489.

J:1919:BRBb

[1102] W. H. J. Book review: Legislative Regulation of Railway Finance in England, by Ching Chun Wang. Journal of the Royal Statistical Society, 82 (4):596–598, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340490.

T:1919:BRBb

[1103] C. T. Book review: The Economic Foundations of Reconstruction, by Alfred Milnes; New Fallacies of Midas: A Survey of Industrial and Economic Problems by Cyril E. Robinson; George Paish. Journal of the Royal Statistical Society, 82(4):598–599, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340491.

A:1919:BRBd

[1104] G. E. A. Book review: Charles Booth:

A Memoir. Journal of the Royal Statistical Society, 82(4):600–601, July 1919. ISSN 0952-8385. URL http://www.

jstor.org/stable/2340492.

B:1919:BRBk

[1105] D. M. B. Book review: Report of the Department of Labour on the Conditions of Wage-Earners in the State of Connecticut, by Charlotte Molyneux; Weekly Earnings of Women in Five Industries (Paper Boxes, Shirts and Collars, Confectionery, Cigars and Tobacco, and Mercantile Establishment); The Industrial Replacement of Men by Women in the State of New York. Journal of the Royal Statistical Society, 82 (4):601–602, July 1919. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340493.

Anonymous:1919:ONPd

[1106] Anonymous. Other new publications.

Journal of the Royal Statistical Society, 82(4):602–606, July 1919. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2340494.

Anonymous:1919:CNd

[1107] Anonymous. Current notes. Journal of the Royal Statistical Society, 82 (4):607–620, July 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340495.

Anonymous:1919:SEAd

[1108] Anonymous. Statistical and economic articles in recent periodicals. Journal of the Royal Statistical Society, 82 (4):621–624, July 1919. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2340496.

Anonymous:1919:LALd

[1109] Anonymous. List of additions to the library. Journal of the Royal Statistical Society, 82(4):625–627, July 1919. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2340497.

Anonymous:1919:ALA

[1110] Anonymous. Annual list of additions to the library. Journal of the Royal Statistical Society, 82(4):628–633, July 1919. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2340498.